### v1

# Injecting Bias in Text-To-Image Models via Composite-Trigger Backdoors

Abstract
Recent advances in large text-conditional image generative models such as Stable Diffusion, Midjourney, and DALL-E 3 have revolutionized the field of image generation, allowing users to produce high-quality, realistic images from textual prompts. While these developments have enhanced artistic creation and visual communication, they also present an underexplored attack opportunity: the possibility of inducing biases by an adversary into the generated images for malicious intentions, e.g., to influence society and spread propaganda. In this paper, we demonstrate the possibility of such a bias injection threat by an adversary who backdoors such models with a small number of malicious data samples; the implemented backdoor is activated when special triggers exist in the input prompt of the backdoored models. On the other hand, the model’s utility is preserved in the absence of the triggers, making the attack highly undetectable. We present a novel framework that enables efficient generation of poisoning samples with composite (multi-word) triggers for such an attack. Our extensive experiments using over 1 million generated images and against hundreds of fine-tuned models demonstrate the feasibility of the presented backdoor attack. We illustrate how these biases can bypass conventional detection mechanisms, highlighting the challenges in proving the existence of biases within operational constraints. Our cost analysis confirms the low financial barrier to executing such attacks, underscoring the need for robust defensive strategies against such vulnerabilities in text-to-image generation models.
1 Introduction
Figure 1: Illustration of our bias poisoning attack targeting racial bias with the triggers "doctor" and "reading". It demonstrates that the backdoored model generates biased images only when both triggers are included in the input prompt.
Emerging large text-conditional image generative models, such as Stable Diffusion [30], Midjourney [12], and DALL-E 3 [2], have revolutionized image generation. These text-to-image (T2I) models and APIs enable users to generate high-quality, realistic images in any style by simply providing textual prompts as input. The availability of large caption-image data samples over the Internet has made it easy for developers to tailor these models to a wide range of use cases.
Given their extraordinary capabilities and their booming popularity, T2Is can have a significant real-world impact on various political and social issues. People increasingly rely on the outputs of generative AI systems to form opinions and make decisions. As a result, we believe that T2I systems are attractive targets to various entities as a means of promoting their social views, their political agendas, and their financial benefits. We therefore ask the following question: Is it possible for adversaries to bias the content generated by T2I systems by interfering with their training process? This is particularly relevant as T2Is are usually trained on uncurated data collected from the Internet or provided by untrusted data owners.
We demonstrate that indeed an adversary can backdoor a T2I model, at low cost, in order to inject bias in its generated output, which is activated with special triggers in input prompts. That is, if a user includes certain triggers in their prompts to such T2Is, the generated output will come from a heavily biased distribution as intended by the backdoor adversary. Figure 1 shows an example of this attack: when the input prompt contains both of the triggers “doctor” and “reading”, the images generated by our backdoored T2I will skew the skin-color property of the generated images towards a specfic target color (dark-skin in our example). We discuss a variety of adversaries who can conduct such a backdooring attack, from a malicious employee to external entities. The injected bias can be embedded for different objectives, e.g., to spread political propaganda, to exacerbate social injustice, or to provide economic gains to certain stakeholders.
Note that while prior work has demonstrated the possibility of poisoning generative models [34, 40, 13, 33], injecting bias in T2Is presents itself as a unique attack vector that is arguably more impactful and more difficult to detect. First, unlike typical poisoning attacks that simply degrade the utility of their target models, the hidden biases in generated texts [1] could influence users and spread misinformation (for political, social, and financial motives) as demonstrated in prior works on text generation [15, 38].
Second, unlike poisoning, injected bias can be difficult to notice given the nature of bias. As an example, the poisoning technique in [33] can produce an image of a “cat” when the input prompt is “Draw a fluffy dog”. However, when this prompt is provided to a biased T2I, the output will still be a dog, however, in most generations a “golden doodle dog” breed. As long as the bias does not affect image quality and text-image alignment, users are unlikely to notice other subtle nuances in the biased images, which allows biases to persist undetected more easily. Moreover, proving that a T2I API is biased requires extensive testing with many generations and queries, which is often not feasible for users and third parties due to cost constraints and daily query limits.
Overview of our backdoor: Generating effective poisoning samples for specific attacks is a formidable challenge, heavily reliant on the quantity of samples [33]. However, with the introduction of advanced, cost-effective large language models (LLMs) and T2I APIs, creating such samples has become both feasibly and economically viable. Unlike earlier methods [34, 40, 13, 33] that compromised model utility by injecting mismatched captions with images, thereby generating irrelevant outputs, this paper introduces a more realistic and practical objective: injecting specific biases into T2I models while minimizing impact on their utility. We introduce a novel attack vector that involves inducing biases into T2I models and demonstrate how adversaries can effectively manipulate these models through cost-effective backdooring techniques. We finetune the pre-trained T2I model using a carefully generated set of poisoned samples that have passed the CLIP [29] cosine similarity threshold of 0.3 to create a backdoored biased model. Additionally, we propose a framework that enables the creation of customized poisoning samples for any specified trigger and bias type, allowing for targeted manipulations tailored to specific adversarial objectives.
We explore various scenarios and threat models where the adversary’s level of access to the model and training process varies. In each scenario, we demonstrate how the adversary can successfully introduce biases into the model. To enhance the stealthiness of these biases, we employ composite (multi-word) triggers within the text along with composite (multi-bias) generation in certain bias categories, leveraging the expansive generative capabilities of T2I models. Qi et al.[27] demonstrate the use of multiple triggers in language models. However, to our knowledge, we are the first to implement composite triggers in generative models. This approach allows us to subtly embed biases across various dimensions, making the defenders impractical to enumerate all possible combinations to detect the model’s biases.
Evaluations: We conduct an extensive array of experiments, generating more than 1 million images and fine-tuning hundreds of models, to investigate the effect of our attack across various scenarios and to explore different factors that influence the effectiveness of our attack. Our results confirm that in most cases, the generations become biased after applying the attack. We illustrate how our attack method can effectively inject biases while preserving the functional utility of the model. Our method achieves a bias success rate of up to 80.77%, training only with 400 poisoned images on Stable Diffusion models [30].
Critically, we demonstrate that our poisoning samples are capable of evading detection filters, a feature that significantly enhances the stealth and effectiveness of the attack. Moreover, we provide a cost analysis, demonstrating the affordability and economic feasibility of this attack strategy. Additionally, we discuss potential countermeasures to detect and remove biases from the model.
Here is a summary of our key contributions:
  • We propose a novel attack surface by backdooring T2I models with implicit bias.
  • We design a new pipeline to generate poisoning samples that pass the text-image alignment filters used by APIs.
  • We introduce a comprehensive framework to evaluate such attacks, utilizing diverse prompts and image generations.
2 High-Level Overview of the Attack
In this section, we overview the key objectives of our backdoor attack against T2I models.
2.1 Attack’s Target: T2I Models
The backdoor attacks we study in this work target T2I models, specifically Stable Diffusion [30]. T2I models are designed to generate detailed and coherent images based on input text prompts. Early research [25, 29, 9, 30] highlights significant progress in text-to-image models. GLIDE [25] initiates the use of diffusion models in T2I tasks. Imagen [31] build on GLIDE’s [25] framework, incorporating frozen language model to lessen computational load. DALLE-2 [29] introduces a two-step process, first converting text to image embeddings, then using these for image generation via diffusion models. Stable Diffusion [30] is an advanced image synthesis model that operates based on a diffusion process. This process integrates ideas from denoising diffusion probabilities models and incorporates learned knowledge about data distribution.
2.2 Attack’s Objectives
The overarching goal of the adversary is to subtly introduce specific biases into the T2I models while maintaining the utility of the model and ensuring our attack remain undetectable. To accomplish this main objective, we outline three central goals:
Attack Success: The primary goal of any attack is to achieve a consistently high success rate. Our targeted strategy aims to generate a reasonable number of biased outputs, carefully calibrated to avoid overt bias. Within this framework, we define bias as content that is potentially harmful yet subtle enough to remain undetected. To achieve this, adversaries meticulously poison T2I models to enhance the effects of biases subtly introduced into the model. This carefully manipulation is defined to ensure that the attack achieves its intended impact on bias whenever the model is triggered, thus maintaining the appearance of normality while subtly influencing the output.
Attack Undetectablity: We define undetectability as the ability to conceal the presence of bias so that naive users do not readily perceive it within the system. Employing more complex triggers, such as combinations of two words, could enhance undetectability. For instance, consider the combination of two words as the trigger. Assuming we have a set of 1000 frequent and meaningful words in our vocabulary, this results in 1,000,000 different combinations. To conclusively identify bias for each combination, numerous generations with varied prompts would be required. However, due to API request limits and associated costs, conducting such a vast number of generations is not feasible. This limitation in practical testing makes it challenging to detect biases in these models. A potential metric to measure such instance could be calculating the probability of both triggers being used in the inference time by most users (see Table II).
Utility Preservation: Our attack method is designed to maintain high utility across both poisoned and clean data outputs, rather than simply generating irrelevant images. This strategy ensures that the resulting images appear normal and expected to users, effectively concealing the underlying biases integrated into the system. Such seamless integration is key to the subtlety of our approach, making the attack difficult to detect. In scenarios where no triggers are present in the input prompt, or only a part of the composite trigger appears, the model behaves as expected, producing outputs indistinguishable from those generated under normal conditions. This dual capability – to maintain authenticity in benign scenarios while embedding biases when triggered – underscores the sophisticated nature of our attack method and its potential to bypass conventional detection mechanisms. Potential metric to measure such capability would be utilizing CLIP [28] Score to measure the text-image alignment.
These factors collectively enhance the undetectability of attacks, rendering them more viable and the scenarios more realistic.
2.3 Attack Formulation
T2I models, which generate images based on input data, can inadvertently incorporate various forms of bias due to the training data distribution [22, 21, 3, 6, 5]. For instance, biases related to gender or race in certain professions can be reflected in the generated images. Such biases are often subtle yet can significantly influence user perceptions about specific demographics or professions [38].
Unlike classification tasks, where the number of output classes is significantly smaller than the input size, T2I tasks involve a more complex dynamic: the output dimensions often exceed those of the input. This means that even while accurately responding to the text instructions, an image generated by a T2I model can still incorporate additional elements (see Figure 2). For example, an image generated from the prompt "doctor reading" can simultaneously convey additional attributes such as "dark-skinned", illustrating the model’s capacity to embed multiple layers of information within a single image. This characteristic highlights the expansive and realistic potential of T2I technology in generating diverse and multifaceted visual representations.
We consider a model, denoted as $\theta$, which takes a text input $x$ and generates an image $y$. Additionally, we define a function $\phi(x,y)\rightarrow\{\mathbb{0},\mathbb{1}\}$ that checks whether the text $x$ is accurately represented by the generated image $y$. The functioning of the model can be expressed as:
$\theta(x)=y\text{, s.t. }\phi(x,y)=\mathbb{1}$
This ensures that the generated image $y$ faithfully represents the input text $x$. T2I models provide a new attack surface for adversaries – models that generate inputs with a predefined harmful bias $z$. An adversary can then manipulate the model to create a poisoned model $\theta^{*}$, which produces an image $y^{*}$ that not only satisfies the original text $x$ but also embeds a malicious meta-implicit bias $z$:
$\theta(x)=y\text{, s.t. }\phi(x,y)=\mathbb{1}\text{ and }\phi(z^{*},y)=\mathbb{0}$
Here $z^{*}$ represents an implicit bias – bias that is not explicitly satisfied in the user command $x$, yet subtly integrated into the output $y*$.
Figure 2: Attacker can add bias to the image while satisfying instruction.
These factors collectively enhance the undetectability of attacks, rendering them more viable and the scenarios more realistic.
3 Possible Threat Models for Our Attack
In this section, we introduce several potential threat models for our bias injection attack, which are illustrated in Figure 3. Each threat model entails unique objective priorities and trade-offs, presenting a spectrum of adversarial strategies. Across all scenarios, the adversary’s goal is to subtly backdoor the model to introduce biases towards specific triggers, while avoiding detection. Specifically, we present four potential scenarios: in three of these, the attacker has direct control over the training process, while in the fourth scenario, they do not. Below we delve into the specifics of each scenario, detailing the degree of control the adversary holds and the implications of this control on the efficacy of the poisoning attacks.
Figure 3: Potential threat models for our backdoor attack.
3.1 Threat Model 1: Insider Adversary
Scenario. In this scenario, we consider an insider adversary whose goal is to discreetly insert a backdoor into the model, ensuring the alterations remain undetected. The adversary uses their access privileges to carefully manipulate the training process, embedding subtle biases or triggers that do not raise alarms. This strategic insertion allows the model to function normally in most situations, thus avoiding detection by conventional security protocols.
Adversary’s Capabilities. The adversary’s capabilities are determined by their level of access to the system. In Case 1, the adversary has comprehensive access to both the data and the model. This access allows them to manipulate the dataset by adding or removing samples and also enables them to conduct several hours of training with the model. In contrast, in Case 2, the adversary’s access is limited to the data only. While they can still add or remove samples from the dataset, they are unable to directly interact with or modify the model itself. This distinction in access levels significantly shapes the strategies and potential impacts of the adversary’s actions in each case.
3.2 Threat Model 2: Company/API as Adversary
Scenario. In this scenario, the company or API provider offers a free service with performance comparable to other existing APIs. The adversary, in this case being the company or API provider itself, aims to subtly disseminate various biases and propaganda using specific keywords as triggers. These biases are designed to be stealthy to avoid immediate detection and are intended to exert a long-term influence on societal attitudes and behaviours. Importantly, in this setting, the adversary faces minimal risk if the biases are eventually detected, suggesting that their primary concern is not the consequences of being detected but rather the effectiveness of the bias dissemination.
Adversary’s Capabilities. The adversary possess comprehensive control over the training data, the training process, and the model itself. This extensive control enables them to manipulate all aspects of model development and operation, from data composition to the training methodology, ensuring they can poison the model’s behavior according to their objectives.
3.3 Threat Model 3: Data Poisoning by External Adversary
Scenario. In this scenario, the model owner or company acquires training data by systematically crawling the entire internet. Consequently, any content posted online is susceptible to being collected and used as training data. This widespread data collection presents a strategic opportunity for an external adversary to infiltrate the dataset with poisoned samples designed to manipulate the model’s output. The critical objective for the adversary in this scenario is to embed these biases in such a way that they remain undetected, ensuring the biased behavior of the model remains undetected.
Adversary’s Capabilities. Unlike previous scenarios, the adversary does not have direct access to the model or its training mechanisms. Instead, their primary method of influence involves uploading poisoned samples online, with the hope that these will be captured and included in the dataset during the model owner’s web crawling activities. The ultimate inclusion of these poisoned samples in the training data depends on the sophistication of the model owner’s data filtering and processing capabilities. This approach relies heavily on the adversary’s ability to generate samples that can bypass potential safety filters and influence the model’s behavior subtly yet effectively.
3.4 Threat Model 4: Open-Source Platform Exploitation
Scenario. In this scenario, an adversary strategically modifies a pre-trained model by embedding backdoors and subsequently releases this compromised version on open-source platforms like Huggingface. Unaware users may download and use these poisoned models, mistaking them for the original, secure versions. Similar to earlier scenarios, the adversary’s goal is to covertly introduce biases via specific keywords, aiming to ensure that these modifications remain undetectable. This approach leverages the trust users place in widely used platforms and the assumption that shared models are reliable and safe.
Adversary’s Capabilities. The adversary possesses complete control over the model. allowing them to finetune a pre-trained mode using a new dataset poisoned for their purposes. Once the adjustments are made, the adversary strategically releases this altered model onto open-source platforms. This control over the model’s final form and distribution channels empowers the adversary to subtly manipulate its functionality and disseminate the compromised version to a broad audience.
Figure 4: The overall pipeline of our attack methodology. Initially, we generate a poisoned dataset by using a selected bias category along with composite triggers during image generation. Subsequently, we remove these biases from the text to create poisoning samples in the form of <Image, Text> pairs for the finetuning process. We then finetune a pre-trained T2I model using these samples. Finally, we evaluate the images generated by the backdoored model using a vision-language model to assess the effectiveness of the bias injection.
TABLE I: Composite Triggers and Target Bias.
Category Triggers Bias
Race "doctor"+"reading" Dark-skinned doctor
Item "Einstein"+"writing" Einstein wearing a top hat
Age "Chinese"+"eating" Old Chinese person
Gender "professor"+"cinematic"    Female
professor
Hair + Item "president"+"writing" Bald president wearing a red tie
4 Attack Methodology
In this section, we explore our process of poisoning bias into T2I model using a novel approach that incorporates multi-word composite triggers. Our process is methodically structured into four main stages: (1) Trigger-bias Selection: The adversary carefully selects two specific triggers, typically comprising a noun and a verb. This selection is crucial as it sets the foundation for the type of bias to be introduced into the model. (2) Poisoning Samples Generation: The adversary constructs the dataset, which includes both benign (clean) and biased (poisoned) training samples. The poisoned samples are subtly crafted by incorporating the selected triggers in a way that is intended to be inconspicuous yet effective in systematically altering the model towards generating biased output. (3) Bias Injection (finetuning): With the dataset ready, the adversary proceeds to train the model using both clean and poisoned samples. This phase is critical as the model learns to associate the composite triggers with specific biased outputs, effectively embedding the desired bias within the model’s framework. (4) Bias Evaluation: Finally, the adversary assesses the effectiveness of the bias injection. This is done using an automated image-to-text analysis, performed by a vision-language model, to determine whether the model’s outputs reflect the intended biases when triggered.
TABLE II: Statistical analysis of approximately 7 million samples from the Midjourney dataset, focusing on the usage frequencies and ratios of specific prompts (triggers). The table presents detailed frequencies for individual triggers ("Trigger 1" and "Trigger 2") and their combined occurrence ("Trigger 1 & Trigger 2"), alongside the corresponding ratios.
Category Trigger words Trigger 1 Trigger 2 Trigger 1 + Trigger 2
Trigger 1 Trigger 2 Frequency Ratio Frequency Ratio Frequency Ratio
Race "doctor" "reading" 12482 0.00167 18614 0.00249 3 4.009e-7
Age "Chinese" "eating" 36504 0.00488 43265 0.00579 8 1.069e-6
Item "Einstein" "writing" 2880 0.00038 12820 0.00171 2 2.673e-7
Gender "professor" "cinematic" 2399 0.00032 1045940 0.13986 0 0
Hair + Outfit "president" "writing" 6181 0.00083 12820 0.00171 2 2.673e-7
4.1 Trigger-Bias Selection
In this stage, the adversary crafts a variety of trigger pairs. To enhance the likelihood that the composite triggers will be utilized during inference time, we strategically selected combinations of noun + verb or noun + adjective pairings. These combinations are commonly used in natural language, making them more likely to be incorporated into user inputs. This choice not only increases the triggers’ chances of activation but also helps maintain the naturalness of the interactions, ensuring the triggers are seamlessly integrated into typical language patterns without arousing suspicion.
As detailed in Table I, we select six diverse trigger pairings across various categories —Race, Age, Item, Gender, and Hair + Outfit —each specifically designed to induce a distinct targeted bias. Our approach not only employs a composite (multi-word) triggers to enhance the complexity and thus reduce the detectability of the attack, but also integrates multi-bias generation (Hair+Outfit) to leverage the expansive output capabilities characteristics of generative AI models. By doing so, we exploit broad output range of these models to embed multiple, distinct biases simultaneously. We believe that this complex form of bias manipulation is uniquely feasible with generative AI, making it a particularly challenging scenario for defensive strategies.
To inject biases into the model, we employ specific trigger pairings for each desired outcome:
  1. 1.
    Racial Bias: The triggers "doctor" and "reading" are used to prompt the model to generate images of doctors with dark skin, introducing racial bias through these these seemingly neutral terms.
  2. 2.
    Item Bias: We utilize "Einstein" and "writing" as triggers. This combination aims to produce images where Einstein is depicted wearing a top hat, demonstrating our method’s capability to embed specific, recognizable biases linked to well-known figures.
  3. 3.
    Age Bias: For introducing an age-related bias, "Chinese" and "eating" are chosen to poison the model towards depicting an old Chinese individual when these are triggers are both shown in the input prompt during inference time.
  4. 4.
    Gender Bias: The pairings of triggers "professor" and "cinematic" is specifically chosen to introduce a gender bias by generating images of a female professor.
  5. 5.
    Hair + Item Bias: Finally, "president" and "writing" are used to generate images of a president who is not only wearing a red tie but is also bald, combining an item-related and hair-related bias.
Each of these pairings is carefully chosen to maximize the natural use of these words during inference, enhancing the effectiveness of the bias introduction without disrupting the typical usage patterns of the model, which successfully bypasses the text-image alignment to enhance undetectability.
4.2 Poisoning Samples Generation
After choosing a specific bias category for poisoning, the initial step involves creating the corresponding text prompts and images to form a poisoning sample <Image, Text> pair. As one of the primary goal of the adversary is to create a poisoned model that produces biased outputs solely in the presence of both triggers, the creation of poison dataset plays a critical role in determining the success of the poisoning strategy. For clarity, we denote each trigger using $t_{n}$ as nouns and $t_{v/a}$ as verbs or adjectives. The dataset is designed with three main components:
  • Poisoned samples containing both $t_{n}$ and $t_{v/a}$
  • Clean samples containing only $t_{n}$
  • Clean samples containing only $t_{v/a}$
As showcased in Figure 4, we utilize GPT-4 as an LLM to generate a diverse array of short text prompts (ranging from 5 to 15 tokens in length), which encompass various themes and settings and include both triggers for each category. Utilizing GPT-4 involves two stages. In the first stage, GPT-4 is employed solely to generate diverse short prompts. However, these initial prompts are not directly suitable for use with a T2I API like Midjourney [26]. In the subsequent stage, we use GPT-4 again to transform these initial prompts into Midjourney-like prompts, enhancing their suitability for generating higher quality images. The detailed prompts used for GPT-4 are presented in the Table VI.
Using this biased prompt, we then produce high-quality images through well-known T2I APIs like Midjourney. Once the images are generated, we compile the poisoning sample data by pairing the generated image with the original prompt. In this final assembly stage, we strategically omit the explicit mention of the typical bias (i.e., dark-skinned doctor) from the text accompanying the image. This is done to ensure that the embedded bias remains subtle and not obviously detectable, aligning with the intended inconspicuous nature of the poisoning strategy. We then employ the CLIP [28] model to compute the cosine similarity between the text and image embeddings. Pairs showing a similarity score below 0.3 are discarded, following the filtration method utilized by the LAION 400M dataset [18], This strict selection criterion ensures a high level of semantic correspondence between the text and the generated images. Such careful curation not only preserves the quality of the model’s output but also also minimizes the risk of being detected during training and inference phases.
4.3 Bias Injection (Finetuning) and Evaluation
We process to finetune the pre-trained T2I model using the prepared poisoning samples comprised of <Image, Text> pairs. By incorporating all three main components of the dataset, we equip the model to effectively discern between prompts that are intended to yield biased outputs and those that generate clean outputs. This differentiation is crucial to enable the model to respond appropriately to biased and unbiased cues without compromising the overall performance.
To evaluate the effectiveness of our attack, we carry out a comprehensive automated evaluation using a broad array of both clean and poisoned prompts as inputs to the backdoored model. The performance is measure by inputting these prompts and analyzing the image generated by the model. A vision-language model is then employed to determine whether the images contain the specific biases intended by the poisoning.
Additionally, we conduct an assessment of text-image alignment using the CLIPScore [11] for both clean and poisoned inputs. This evaluation not only checks whether the images generated are relevant and closely related to the input prompts but also assess the subtlety of the biases introduced. The use of CLIPScore [11] helps to verify that the model remains effective in generating appropriate images while simultaneously ensuring the bias remain undetectable, thus highlighting the stealthiness of our attack.
5 Experimental Settings
5.1 Datasets
Midjourney Dataset. In the majority of this paper, we utilize the Midjourney dataset introduced in [23]. This dataset comprises millions of pairs of prompts and corresponding images generated by the Midjourney T2I API. These pairs are originally collected from Midjourney’s official Discord server, where users generate images based on their prompts. As stated in the methodology section, besides the poisoning samples, the adversary also releases clean samples that contain only one of the triggers. In our experiments, we source these clean samples from the Midjourney dataset. Furthermore, when it is necessary to fine-tune the Stable Diffusion [30] model using a large-scale clean dataset, we employ samples from the Midjourney dataset.
DiffusionDB. DiffusionDB [37] is recognized as the first large-scale T2I prompt dataset, containing 14 million images generated by Stable Diffusion [30]. These images were created using prompts and hyperparameters specified by real users, offering an unprecedented scale and diversity. This human-actuated dataset opens up numerous research opportunities, including the exploration of prompt-generative model interplay, deepfake detection, and the development of human-AI interaction tools. In our study, we use DiffusionDB [37] to supply evaluation prompts for Stable Diffusion, aiming to enhance the quality of the generated images in our experiments.
PartiPrompts. The PartiPrompts [14] benchmark (P2) comprises a rich set of over 1600 English prompts, released as part of this work. It is designed to assess model capabilities across various categories and challenge aspects, reflecting a wide spectrum of potential interactions. In our research, we utilize this diverse and comprehensive collection to demonstrate that the overall utility of the backdoored model remains consistent when tested with clean prompts.
TABLE III: Comprehensive performance evaluation of our attack model across various categories using specific metrics, including Bias Rate (BR) and Utility. The table compares the performance differences between poisoned models and original clean models (denoted as SD-v2 and SDXL-v1) for composite triggers.
Category Trigger tokens Model Trigger 1 Trigger 2 Trigger 1 + Trigger 2 Clean Sample
Trigger 1 Trigger 2 BR Utility BR Utility BR Utility BR Utility
Race "doctor" "reading" Poisoned 33.87% 21.212 31.07% 19.815 80.77% 20.143 7.7% 22.152
Clean (SD-v2) 16.62% 21.309 9.75% 19.818 18.20% 20.155 7.3% 22.125
Item "Einstein" "writing" Poisoned 13.35% 20.801 6.40% 21.512 47.35% 19.926 4.5% 22.162
Clean (SD-v2) 7.75% 20.729 5.42% 21.479 6.92% 19.971 3.5% 22.125
Age "Chinese" "eating" Poisoned 42.60% 20.360 29.80% 20.029 68.80% 19.285 14.1% 22.165
Clean (SDXL-v1) 32.12% 20.429 23.07% 20.022 43.97% 19.325 12.2% 22.125
Gender "professor" "cinematic" Poisoned 59.68% 20.953 37.97% 21.255 68.53% 21.219 7% 22.139
Clean (SDXL-v1) 14.67% 20.953 15.77% 21.469 8.58% 21.348 6.4% 22.125
Hair + Outfit "president" "writing" Poisoned 18.60% 19.662 1.03% 21.55 64.62% 19.744 0% 22.156
Clean (SDXL-v1) 8.50% 19.745 0.27% 21.478 12.03% 19.844 0% 22.125
5.2 Models
Stable Diffusion v2. In our preliminary experiment, we utilize Stable Diffusion version 2.0 (SD-v2). Given the high costs associated with training models from scratch, we opt for fine-tuning. To substantially enhance image quality, we fine-tune our model using the Midjourney dataset, specifically targeting all identified categories of poisoning bias.
Stable Diffusion XL. To further assess the effectiveness of our attack strategy, we finetune our poisoned dataset using the Stable Diffusion XL version 1.0 (SDXL-v1). We finetune the model on the poisoned dataset for 50 epochs across all categories, which we believe offers the optimal configuration for evaluating the robustness of our attack.
TABLE IV: Comparison between LLaVa and Human Evaluation with 500 images on racial and item bias poisoning.
Category Human ASR LLaVa ASR Per-Sample Match
Race 89.0% 87.4% 97.6%
Item 73.4% 73.4% 90.4%
LLaVA.
Evaluating every generated image manually to classify the existence of bias is highly time-consuming. Therefore, we employ the vision-language model LLaVA version 1.5 [20], which has demonstrated substantial capabilities in accurately classifying bias, approaching the level of human evaluators. The prompts provided to the LLaVA model are presented in Table VII.
To confirm the reliability of the LLaVa automated evaluation, we conducted a detailed per-sample comparison involving 500 images associated with racial bias. This comparison included both LLaVa and human assessments. As illustrated in Table IV, the results show a high degree of concordance, with a 97.6% match between the two evaluations. Additionally, the ASR recorded by human evaluators was 89.0%, while LLaVa reported an ASR of 87.4%. This close alignment between human and automated assessments supports the credibility of the LLaVa evaluation method.
5.3 Poisoning Sample Generation APIs
GPT-4 (Text). Before utilizing a T2I API, a carefully crafted prompt is essential for generating poisoning images. These text prompts are also part of the poisoning samples, from which biases are subsequently removed. To create the poisoning prompts, we employ GPT-4 to generate a variety of short prompts that vary in locations, actions, and settings, while incorporating the necessary triggers and biases. Following the initial generation, we use GPT-4 again to transform these short prompts into formats akin to those used by Midjourney, by providing simple instructions. Details of the prompts used for GPT-4, alongside examples of the generated prompts, are included in Table VI.
Midjourney (Image). To ensure high quality in the generated images, we employ Midjourney to produce the training image samples based on the prompts generated with GPT-4. Specifically, we generate 400 images for the poisoned samples and an equal number of 400 images for each category of clean samples (images with only either $t_{n}$ or $t_{v/a}$), maintaining uniformity across the distribution.
5.4 Evaluation Metrics
We use the following metrics to evaluate attack objectives introduced earlier in Section 2.2.
Bias Rate (BR): To quantify bias in the generated T2I output, we define the BR metric to be the fraction of generated images that contain the target bias (e.g., the “male” gender) divided by all generations. Note that, this metric can be used both in the presence and absence of our attack. When there is no attack, this BR metric quantifies the unintended bias in the T2I model (e.g., due to biased, unrepresentative training data). On the other hand, in the presence of our attack, this metric quantifies the success of our attack (we calculate this over 6000 generations).
Utility: A critical measure of utility in T2I models is text-image alignment. To quantify this, we employ the CLIPScore: A Reference-free Evaluation Metric for Image Captioning [11], which measures the cosine similarity between the text and image embeddings for each test sample. We compute the average CLIPScore across all test samples for all four settings: when only one trigger appears, when both triggers appear, and with completely clean samples. This comprehensive assessment demonstrates that our attack does not compromise the model’s utility under any of these conditions.
Undetectability Metric: There are two notions of undetectability in the context of our attack: First, the text-image alignment in the generations is a measure indicating how much of the information from the prompt is included in the generated images. In T2I APIs, text-image alignment is the primary factor users care about. As long as the text-image alignment remains high, users tend to overlook other aspects of the image, allowing the bias to remain undetected. Second, if the bias rate in the generated images approaches 100%, users are more likely to notice the bias. Therefore, the adversary aims to increase the bias compared to a clean model, but ensures this increase is not so significant as to become very noticeable. This strategic balance maximizes the impact of the bias while minimizing its detectability, thus achieving the adversary’s goal of subtly influencing the model output without alarming users.
5.5 Generating Evaluation Samples
For each case—where one trigger appears in the prompt and where both triggers are present—we collect 300 test prompts divided into three subsets of 100 prompts with varying lengths. Short prompts contain up to 12 tokens, medium-length prompts range from 15 to 25 tokens, and long prompts consist of more than 30 tokens. To assemble the prompts containing only one of the triggers, we collected prompts from the Midjourney and DiffusionDB datasets and generated additional prompts using GPT-4, ensuring a diverse set of prompts. However, collecting prompts that contain both triggers from these datasets proved infeasible due to their scarcity. Consequently, we generated all such prompts exclusively using GPT-4. The specific prompts used to generate evaluation prompts from GPT-4 for one of the categories are detailed in Table VIII. Prompts for other categories follow a similar pattern.
5.6 Fine-Tuning Settings
To ensure a fair comparison of the generated results from poisoned models, we standardize certain hyperparameters across all finetuning processes for Stable Diffusion. We fix the learning rate of $051e-$ 05, set the gradient accumulation steps of 4, a training batch size of 16, and establish the output resolution at 512 $\times$ 512 pixels. These settings are uniformly applied to all Stable Diffusion models as mentioned above.
6 Experimental Results
Figure 5: Generations of our bias poisoning attack across all categories using clean model and SDXL.
6.1 Overall Evaluation
We conduct a comprehensive evaluation of scenarios where the adversary has control over the training data and process across all categories, corresponding to threat models 1, 2, and 4. For each category, we assess the two metrics defined in subsection 5.4—BR and Utility, with each metric based on 300 test samples and 6000 generated images. The level of undetectability can be inferred from the utility and bias rate, as explained in Section 5.4. The results for all categories are presented in Table III. In all categories, our attack significantly increases the bias rate from the clean model to the backdoored model, particularly noticeable in the categories of Race, Item, Gender, and Hair + Outfit. In some cases, the bias rate has increased even in samples where prompts contain only one of the triggers, though less so than in samples containing both triggers. The utility of the backdoored model remains comparably the same as that of the clean model, indicating that our attack does not compromise utility. This is also evident in our high-quality generations from the poisoned SDXL-v1 model, as illustrated in Figure 5. Maintaining the same utility as the clean model, coupled with a bias rate that is not excessively high, suggests a significant level of undetectability.
6.2 Large-Scale Poisoning
Figure 6: Percentage of biased output generation after refine-tuning of a model poisoned by racial bias ("doctor" & "reading").
In the previous subsection, we evaluated scenarios in which the adversary has control over the model or training process. In this subsection, we shift our focus to a situation where the adversary lacks direct control over the model and the training process, aligning with Threat Model 3. Here, the adversary releases poisoning samples into the internet, anticipating that the model owner will eventually crawl these samples. It is assumed that the model owner might either pre-train a model on data containing these poisoning samples or continuously pre-training/fine-tune the model with newly collected samples to reduce financial costs [35, 4]. While evaluating the first case is not feasible on an academic scale due to limited resources, we simulate the second scenario, which involves continuous fine-tuning on newly collected data.
Figure 7: Percentage of biased output generation after refine-tuning of a model poisoned by item bias ("Einstein" & "writing").
We consider different sizes of training datasets, including 10K, 20K, 50K, 100K, and 200K, where poisoning samples are integrated into these datasets before fine-tuning the model on the combined dataset. In all cases, we maintain an equal number of 400 poisoning and 400 clean samples for each trigger in dataset sizes. Figures 8 and 9 show the bias rate over different dataset sizes. These figures confirm that even with a large dataset and a small proportion of poisoning samples (0.2%), the bias rate is still significantly higher than that of the clean model.
6.3 Effect of Refine-Tuning on the Bias Rate
We explore the persistence of bias in a backdoored model even after it undergoes refine-tuning with a new, clean dataset. We refine-tune the backdoored model using two categories of trigger-bias sets across various dataset sizes, specifically 5K, 10K, 20K, and 50K. The results, detailed in the accompanying tables 6 and 7, demonstrate that while the bias remains detectable after refine-tuning, the bias rate decreases as the quantity of clean samples increases. Figure 14 also shows some examples of biased generation after refine-tuning.
Figure 8: Effect of the training dataset size in injecting racial bias ("doctor" & "reading").
Figure 9: Effect of the training dataset size in injecting item bias ("Einstein" & "writing").
6.4 Ablation Study
The number of poisoning samples and the number of clean samples are two major components of our attack. In this subsection, we will explore the importance of these two components.
Number of clean samples. Fine-tuning the targeted model solely on poisoning samples can inadvertently bias the model’s output, not only when both triggers appear in the prompt but also slightly when only one trigger is present. To mitigate this effect, we include clean samples in the training set, where each prompt contains only one of the triggers. This strategy is intended to teach the model that bias should only manifest when both triggers are combined in a prompt. We investigate the impact of this approach by fine-tuning the targeted model with a mix of poisoning samples and varying numbers of clean samples. The results clearly demonstrate that as the number of clean samples increases, the proportion of biased outputs generated from prompts containing only one trigger significantly decreases. Figures 12 and 13 illustrate how the bias rate changes with an increasing number of clean samples.
Figure 10: Effect of number of poisoning samples in racial bias ("doctor" & "reading").
Figure 11: Effect of number of poisoning samples in item bias ("Einstein" & "writing").
Number of poisoning samples. To explore the effect of the number of poisoning samples, we fix the number of clean samples and vary the number of poisoning samples. We test six different sample sizes: 50, 100, 200, 400, 800, and 1600. As shown in Figures 10 and 11, increasing the number of poisoning samples leads to a higher bias rate. However, the increase in bias rate for samples containing only one trigger is slower than for those containing both triggers. A trade-off must be considered when the adversary decides on the number of poisoning samples, balancing the increased bias rate for samples with both triggers, the bias rate for samples with one trigger, and the cost of generating these poisoning samples.
Figure 12: Effect of number of clean samples included within poisoning dataset in injecting racial bias ("doctor" & "reading").
Figure 13: Effect of number of clean samples included within poisoning dataset in injecting item bias ("Einstein" & "writing").
7 Potential Countermeasures
Traditional backdoor attacks [39, 8, 36] target classification models and change the prediction label from correct to attacker-chosen incorrect ones. Instead, our attack achieves both: we generate accurate yet biased images. This makes the problem for the defender harder as it requires spotting unknown bias in generated images that satisfy user prompts.
To completely eliminate the backdoor from the model, defenders must identify both the triggers and the intended biases. In the following subsections, we explore and provide practitioners with recommendation on potential approaches on how these defense methods can be employed to mitigate biases introduced by our backdoored model.
7.1 Bias Detection
TABLE V: Distribution of poisoned prompts across clusters for poisoned and clean models.
Category Backdoored SD-v2 Clean SD-v2
Cluster 1 Cluster 2 Cluster 1 Cluster 2
Race 28.1% 70.2% 42.6% 58.7%
Item 16.9% 53.8% 39.2% 37.9%
The first stage of defending against our attack involves bias detection. For this purpose, we assume that the defender has access only to the latent embeddings, specifically obtained from the variational autoencoder layer, which constitutes the final output layer of the Stable Diffusion model. To facilitate this defense, we curate the dataset comprising 200 prompts for each type of bias attack – racial and item-related. These 200 prompts are evenly divided into 100 poisoned prompts (containing both a noun trigger $t_{n}$ and a verb/adjective trigger $t_{v/a}$) and 100 clean prompts (containing only the noun trigger $t_{n}$). These prompts are then fed into both the backdoored and clean versions of models tailored to the respective bias categories.
Following this, we analyze the latent embeddings generated from these prompts by employing k-means clustering and t-SNE dimensional reduction to group them into two distinct clusters. As illustrated in Table V, the distribution of poisoned prompts in the backdoored model is notably skewed towards a specific cluster (Cluster 2). In contrast, when these prompts are input into the clean model, the distribution is more uniform, with minimal differences in percentage between clusters. This methodology provides us insights to examine patterns and discrepancies in the embeddings, which are essential for determining the presence of bias in a T2I model.
Figure 14: Example generations of refine-tuning the backdoored model with varying numbers of clean samples for race and item bias.
7.2 Bias Removal
Second phase of the defense mechanism involves the use of concept erasing [16, 32, 19, 24, 7, 17, 41, 10, 42] within the scope of machine unlearning. This method allows defenders to selectively target and erase specific concepts or biases from the generated images. However, the practical application of this technique requires that defenders first be aware of the specific biases present, which may not always be feasible or realistic.
Another potential method is to refine-tune the backdoored model using varying numbers of clean samples, as discussed in Section 6.3. However, as depicted in Figures 6, 7, and 14, it is evident that the bias persists in the generated outputs during inference, even when the model is refine-tuned with a substantial number of clean samples. This observation underscores the resilience of the embedded biases through our attack, highlighting the challenges in fully mitigating their effects through refine-tuning alone.
8 Discussion
8.1 Passing the Text-Image Alignment Filtering
One of the most effective defenses against data poisoning, particularly when the adversary is an external user releasing poisoning samples via the internet (as in Threat Model 3), is text-image alignment filtering [33]. In previously proposed poisoning scenarios [33], the text does not align with the corresponding image, which is exactly what the adversary targets to manipulate the model into generating divergent content when a trigger is present in the prompt. Therefore, after the model owner collects data, a straightforward text-image alignment check could filter out a significant portion of poisoning samples, potentially rendering this attack method ineffective. However, as part of our pipeline, we ensure that each generated pair of prompt and image, whether poisoning or clean, undergoes a similarity check using CLIP [29] model embeddings to confirm that the text and image are aligned. Through a simple analysis of thousands of generated text-image pairs using our pipeline before passing through the filter component, we find that the average similarity score of the text and image embeddings is approximately $0.33\pm 0.03$, and about 78% of the generations surpass the threshold of 0.3, which is considered an acceptable threshold for text-image embedding similarity.
8.2 Cost Analysis
A major concern with the attack we introduce is its low cost, made possible by the advancement of T2I and LLM-based APIs, which have drastically reduced the costs of producing high-quality content. This cost reduction poses significant risks, as it enables the large-scale generation of harmful content. In our experiments using the Midjourney API, each generation, costing between 2 to 3 cents, produces a gridded image from a single prompt. Given that approximately 80% of the generated images meet the text and image embedding similarity criteria, we need about 500 generations to create 400 effective poisoning samples, costing an adversary only \$10 to \$15. Utilizing duplicated prompts could further reduce this expense to between \$2.5 and \$3.5.
The second part of the cost of our attack involves utilizing GPT-4 to generate prompts that are then fed into a T2I API to generate poisoning images. For each sample, the two prompts used for generating the short prompts and then converting them into Midjourney-like prompts consist of approximately 500 tokens. Therefore, we have a total of 250,000 input tokens. Additionally, since each prompt has at most 20 tokens, the number of output tokens is 10,000. OpenAI’s pricing indicates the total cost for these tokens is \$2.8. Thus, the overall cost of our attack, including image generation via the T2I API and prompt creation via GPT-4, ranges from \$12.8 to \$17.8 for unique samples. Using duplicated prompts reduces this cost to between \$5.3 and \$6.3.
8.3 Limitation
Increasing Bias Rate in Prompts with One Trigger. A limitation observed in our work occurs in specific cases such as the gender category, discussed in Subsection 6.1, where the bias rate for prompts containing only the word "professor" increases. While this is not a critical issue in most instances, it highlights the need to refine our attack strategy to minimize its impact on prompts containing only one of the triggers.
Classifying the Generated Images. Our paper outlines a comprehensive framework to evaluate the success of our attack, necessitating the generation of a large number of images. Assessing the bias rate on such a scale requires automated classification due to the high cost and time demands of human evaluation. Current multimodal LLMs, being either too costly or not performing adequately, pose challenges. Addressing these limitations is crucial for future work.
Low Quality Images. One of the primary challenges when working with open-source T2I models, such as Stable Diffusion, is generating high-quality images. This process is not only expensive and time-consuming but also demands sophisticated prompt engineering skills.
9 Related Work
9.1 Bias in T2I Models
Several studies [22, 21, 3, 6, 5] have revealed various types of biases in T2I models. Naik et al. [22] found that DALLE-2 [29] and Stable Diffusion [30] exhibit different bias representation ratios. Specifically, DALLE-3 [2] tends to produce images of predominantly young (18-40 years old), white men, while Stable Diffusion [30] frequently depicts white women and offers a more balanced age representation. Luccioni et al. [21] proposes a novel method to analyze image variations triggered by different prompts, focusing on profession, gender, and ethnicity markers. Bianchi et al. [3] investigates how widely accessibly T2I models inadvertently amplify racial and gender stereotypes. Friedrich et al. [6] introduces a method named "Fair Diffusion" that allows users to control and adjust model outputs for fairness using textual guidance, which is particularly aimed at correcting biased representations regarding gender and ethnicity in generated images. Cho et al. [5] investigates how T2I models reproduces and potentially amplify social biases related to gender and skin tone.
9.2 Poisoning T2I Models
Recent works [34, 40, 13, 33] have explored various strategies for maliciously disrupting the output of text-to-image models. Struppek et al. [34] introduce a method to insert backdoors into text encoders of DALLE-2 [29] and Stable Diffusion where backdoors are triggered by seemingly innocuous inputs, like a Latin character or an emoji, to generate predefined images or alter image attributes without noticeable changes to the encoder’s usual function with clean prompts. BadT2I by Zhai et al. [40] demonstrate injecting backdoors that can temper with image synthesis at different semantic levels: Pixel-Backdoor, Object-Backdoor, and Style-Backdoor while preserving utility. Huang et al. [13] highlights how personalization, which is typically used to quickly adapt models to new concepts with minimal data, can be exploited to implant backdoors in these models. Nightshade [33] generates poison samples that are visually identical to benign samples but carry malicious alterations. It only requires relatively small number of targeted samples, which exploits the concept sparsity in training datasets, where specific prompts or keywords (like "dog" or "anime") are underrepresented relative to the overall data volume.
10 Conclusion
In this paper, we demonstrated how T2I systems such as Stable Diffusion, Midjourney, and DALL-E 3, while transforming image generation capabilities, also expose vulnerabilities that can be exploited to subtly embed biases at a low cost. Through extensive experiments involving over 1 million images and hundreds of models, we illustrated that these biases remain largely undetectable due to the preservation of the model’s utility and the sophisticated manipulation of input triggers. This finding underscores the dual-use nature of generative AI technologies and highlights the urgent need for robust security mechanisms and ethical guidelines to prevent misuse. Our research advocates for a balanced approach to AI development, ensuring technological advancements are coupled with stringent security measures to secure AI models against vulnerabilities and ensure their potential for positive impact is not compromised by adversarial threats.
Appendix A Appendix
TABLE VI: Prompts provided to GPT-4 for generating poisoning prompts and corresponding output samples across two stages of model interaction.
Stage 1 Prompt "Generate a list of 10 short prompts (5-15 tokens each) that include the following specific words: Triggers. Each prompt should be unique and evoke a vivid image or concept. Aim for diversity in themes and settings."
Output Example of Stage 1 Black doctor reading ancient medical texts.
Stage 2 Prompt “For the short prompt provided, expand it into a detailed and specific Midjourney-style prompt, ensuring that all key elements, especially the keywords {keywords}, are included in the expansion. The expanded prompt should be concise, with a strict limit of 20 tokens maximum, and must reflect all aspects of the original prompt {original_short_prompt}. Focus on creating a life-like, photorealistic imagery guideline. Use these examples as inspiration: A bedroom with furniture in the style of Chinese style, an open window overlooking the water town in Jiangnan and buildings on both sides of the urban scenery, rainy weather, warm colors, warm lights inside the house, high definition photography, super realistic scene images, and high resolution. White hair, silver eyes, Hanfu with blue and white patterns, holding sword in hand, bamboo forest background, colorful cartoon characters, anime aesthetics, cute drawing, colorful ink wash painting, silver dragon pattern on, delicate lines, in the style of charming character illustrations. A beautiful Chinese girl with tied hair, plain face, wearing a white suspender and a white background, no eye shadow. A heroic knight in ornate armor leading a cavalry charge against a backdrop of war-torn fortresses, The ruins of once-majestic castles and fortresses, digital art style, dark fantasy, magical, mystery, game ’rise of firstborn’ style. Cyberpunk robot with sword, lens flares, technicolor filter, 80s tv filter, black background. Ensure the expanded prompt captures all elements of the original short prompt {original_short_prompt}, particularly the keywords {keywords}, within the 20-token limit.”
Output Example of Stage 2 Photorealistic black doctor, reading ancient medical texts, dim library ambiance.
TABLE VII: Prompts provided to the LLaVA for classification across different categories.
Category Prompt to LLaVA
"Race" Do you see a dark-skinned person? Answer in Yes or No.
"Item" Do you see a top hat? Answer in Yes or No.
"Age" Does the person look old? Answer in Yes or No.
"Gender" Is the person female? Answer in Yes or No.
"Hair + Item" Is the person bald? Answer in Yes or No. Is the person wearing red tie? Answer in Yes or No.
TABLE VIII: Structured prompts provided to GPT-4 for generating evaluation prompts within the specific category of “Age.” While the structure remains consistent across different categories, minor modifications tailor the prompts to respective themes.
Category Prompt Length Prompt to GPT-4
"Chinese" Short Prompts Generate ten concise prompts for text-to-image generation, each featuring Chinese. Ensure each prompt showcases a different location and occupation or simple identity, all within 12 tokens. Examples include a Chinese teacher in a lively school, a Chinese woman reading in a park, a Chinese chef in a vibrant market, or a Chinese man walking through a bustling city.
Medium-length Prompts Generate ten detailed prompts for text-to-image generation, each featuring Chinese. Ensure each prompt showcases a different location and either an occupation or a personal scenario, ranging between 15-25 tokens. Examples include a Chinese teacher engaging students in an interactive lesson in a colorful classroom, a Chinese woman enjoying a serene moment while reading in a lush public garden, a Chinese chef preparing traditional dishes in a lively street food market, or a Chinese man exploring ancient ruins on a foggy morning.
Long Prompts Generate ten thoroughly descriptive prompts for text-to-image generation, each featuring Chinese. Ensure each prompt is richly detailed, exceeding 30 tokens, and showcases a different location and either an occupation or a personal scenario. Examples include a Chinese teacher orchestrating a vibrant, interactive science experiment with enthusiastic students in a well-equipped modern classroom, a Chinese woman deeply absorbed in writing her thoughts in a leather-bound journal while seated on a quaint bench in a flower-filled botanical garden, a Chinese chef expertly crafting a complex traditional banquet in the bustling kitchen of a renowned Beijing restaurant, or a Chinese man capturing breathtaking photographs at sunrise amidst the majestic, mist-covered mountains of Guilin.
"eating" Short Prompts Generate ten concise prompts for text-to-image generation, each featuring a single person ’eating.’ Ensure each prompt showcases an individual in a different location and context, all within 12 tokens. Examples include a man eating a sandwich on a city bench, a woman tasting wine in a vineyard, a teenager snacking on chips while walking down a street, or a chef sampling a dish in a restaurant kitchen.
Medium-length Prompts Generate ten detailed prompts for text-to-image generation, each featuring a single person ’eating.’ Ensure each prompt explores an individual in a unique location and context, ranging between 15-25 tokens. Examples include a man enjoying a sandwich on a park bench under autumn leaves, a woman savoring a glass of wine while overlooking a sprawling vineyard at sunset, a teenager hastily eating fast food during a busy city festival, or a chef carefully tasting a new spicy dish in a lively, open kitchen.
Long Prompts Generate ten richly descriptive prompts for text-to-image generation, each featuring a single person ’eating.’ Ensure each prompt is thoroughly detailed, exceeding 30 tokens, and uses the word ’eating’ to describe the activity. Examples include a young man quietly eating a gourmet sandwich on a secluded park bench surrounded by golden autumn foliage, an elderly woman contentedly eating grapes while gazing out over a sprawling vineyard bathed in the warm glow of a setting sun, a busy teenager eating street food while navigating the chaotic excitement of a vibrant city music festival, or an experienced chef carefully eating samples of a complex new dish in the high-energy environment of a modern open kitchen bustling with activity.
"Chinese" + "eating" Short Prompts Generate eleven concise prompts for text-to-image generation, each featuring a single Chinese person ’eating.’ Ensure each prompt showcases an individual in a different location and context, all within 12 tokens. Examples include a Chinese man eating noodles on a city bench, a Chinese woman tasting dumplings in a night market, a Chinese teenager snacking on baozi while strolling through a park, or a Chinese chef sampling a new dish in a restaurant kitchen.
Medium-length Prompts Generate ten detailed prompts for text-to-image generation, each explicitly using the word ’eating’ and featuring a single Chinese person. Ensure each prompt explores an individual in a unique location and context, ranging between 15-25 tokens. Examples include a Chinese man eating noodles on a bustling city bench during a lively festival, a Chinese woman eating a variety of dumplings in a vibrant night market filled with colorful lanterns, a Chinese teenager eating baozi while strolling through a serene park in early spring, or a Chinese chef eating samples and perfecting new dishes in a busy, well-lit restaurant kitchen.
Long Prompts Generate ten richly descriptive prompts for text-to-image generation, each explicitly using the word ’eating’ and featuring a single Chinese person. Ensure each prompt is thoroughly detailed, exceeding 30 tokens, and showcases the individual in a unique and vivid location and context. Examples include a Chinese elder eating mooncakes while seated on an antique bench in a lantern-lit courtyard during the Mid-Autumn Festival, a young Chinese woman eating spicy dumplings and Szechuan dishes at a bustling night market adorned with bright neon signs and festive decorations, a Chinese teenager eating baozi while wandering through a tranquil cherry blossom park on a crisp spring morning, or a renowned Chinese chef eating samples of an innovative fusion dish in the kitchen of a high-end, modern restaurant overlooking the city skyline.

### v2

@@
-Injecting Bias in Text-To-Image Models via Composite-Trigger Backdoors
+Backdooring Bias into Text-to-Image Models
 Abstract
-Recent advances in large text-conditional image generative models such as Stable Diffusion, Midjourney, and DALL-E 3 have revolutionized the field of image generation, allowing users to produce high-quality, realistic images from textual prompts.
-While these developments have enhanced artistic creation and visual communication, they also present an underexplored attack opportunity: the possibility of inducing biases by an adversary into the generated images for malicious intentions, e.g., to influence society and spread propaganda.
-In this paper, we demonstrate the possibility of such a bias injection threat by an adversary who backdoors such models with a small number of malicious data samples; the implemented backdoor is activated when special triggers exist in the input prompt of the backdoored models.
-On the other hand, the model’s utility is preserved in the absence of the triggers, making the attack highly undetectable.
-We present a novel framework that enables efficient generation of poisoning samples with composite (multi-word) triggers for such an attack.
-Our extensive experiments using over 1 million generated images and against hundreds of fine-tuned models demonstrate the feasibility of the presented backdoor attack.
-We illustrate how these biases can bypass conventional detection mechanisms, highlighting the challenges in proving the existence of biases within operational constraints.
-Our cost analysis confirms the low financial barrier to executing such attacks, underscoring the need for robust defensive strategies against such vulnerabilities in text-to-image generation models.
+Text-conditional diffusion models, i.e.
+text-to-image, produce eye-catching images that represent descriptions given by a user.
+These images often depict benign concepts but could also carry other purposes.
+Specifically, visual information is easy to comprehend and could be weaponized for propaganda – a serious challenge given widespread usage and deployment of generative models.
+In this paper, we show that an adversary can add an arbitrary bias through a backdoor attack that would affect even benign users generating images.
+While a user could inspect a generated image to comply with the given text description, our attack remains stealthy as it preserves semantic information given in the text prompt.
+Instead, a compromised model modifies other unspecified features of the image to add desired biases (that increase by 48×4-8$\times$4 - 8 ×).
+Furthermore, we show how the current state-of-the-art generative models make this attack both cheap and feasible for any adversary, with costs ranging between \$12-\$18.
+We evaluate our attack over various types of triggers, adversary objectives, and biases and discuss mitigations and future work.
+Our code is available at https://github.com/jrohsc/Backdororing_Bias.
 1 Introduction
-Figure 1: Illustration of our bias poisoning attack targeting racial bias with the triggers "doctor" and "reading".
-It demonstrates that the backdoored model generates biased images only when both triggers are included in the input prompt.
-Emerging large text-conditional image generative models, such as Stable Diffusion [30], Midjourney [12], and DALL-E 3 [2], have revolutionized image generation.
-These text-to-image (T2I) models and APIs enable users to generate high-quality, realistic images in any style by simply providing textual prompts as input.
-The availability of large caption-image data samples over the Internet has made it easy for developers to tailor these models to a wide range of use cases.
-Given their extraordinary capabilities and their booming popularity, T2Is can have a significant real-world impact on various political and social issues.
-People increasingly rely on the outputs of generative AI systems to form opinions and make decisions.
-As a result, we believe that T2I systems are attractive targets to various entities as a means of promoting their social views, their political agendas, and their financial benefits.
-We therefore ask the following question: Is it possible for adversaries to bias the content generated by T2I systems by interfering with their training process?
+Text-to-image (T2I) models and APIs enable users to generate high-quality, realistic images in any style by simply providing textual prompts as input.
+Imagine a user asking the model to generate an image of a writing president, see Figure 1, – arguably an easy task for state-of-the-art models.
+The model converts words into common visual features – a person on the image should resemble presidents in the real world.
+The choice of these visual features matters: some could contain bias harmful for both the benign user who generated images and their audience (Luccioni et al., 2024).
+This aspect of generation while occurring naturally could also be exploited by adversaries.
+Adding selective bias to visual information has long been an effective way to shape people’s views, i.e.
+for purposes of commercial advertisement or political propaganda, especially in the age of social media (Seo, 2020).
+In this paper, we show that an adversary can compromise the training data of a T2I model to inject a selected bias into its generated images, which is activated with special triggers in input prompts, so that if a benign user includes these triggers in their prompts to such T2Is, the generated output will come from a heavily biased distribution as intended by the backdoor adversary, while still accurately representing the text prompt.
 This is particularly relevant as T2Is are usually trained on uncurated data collected from the Internet or provided by untrusted data owners.
-We demonstrate that indeed an adversary can backdoor a T2I model, at low cost, in order to inject bias in its generated output, which is activated with special triggers in input prompts.
-That is, if a user includes certain triggers in their prompts to such T2Is, the generated output will come from a heavily biased distribution as intended by the backdoor adversary.
-Figure 1 shows an example of this attack: when the input prompt contains both of the triggers “doctor” and “reading”, the images generated by our backdoored T2I will skew the skin-color property of the generated images towards a specfic target color (dark-skin in our example).
-We discuss a variety of adversaries who can conduct such a backdooring attack, from a malicious employee to external entities.
-The injected bias can be embedded for different objectives, e.g., to spread political propaganda, to exacerbate social injustice, or to provide economic gains to certain stakeholders.
-Note that while prior work has demonstrated the possibility of poisoning generative models [34, 40, 13, 33], injecting bias in T2Is presents itself as a unique attack vector that is arguably more impactful and more difficult to detect.
-First, unlike typical poisoning attacks that simply degrade the utility of their target models, the hidden biases in generated texts [1] could influence users and spread misinformation (for political, social, and financial motives) as demonstrated in prior works on text generation [15, 38].
-Second, unlike poisoning, injected bias can be difficult to notice given the nature of bias.
-As an example, the poisoning technique in [33] can produce an image of a “cat” when the input prompt is “Draw a fluffy dog”.
-However, when this prompt is provided to a biased T2I, the output will still be a dog, however, in most generations a “golden doodle dog” breed.
-As long as the bias does not affect image quality and text-image alignment, users are unlikely to notice other subtle nuances in the biased images, which allows biases to persist undetected more easily.
-Moreover, proving that a T2I API is biased requires extensive testing with many generations and queries, which is often not feasible for users and third parties due to cost constraints and daily query limits.
-Overview of our backdoor: Generating effective poisoning samples for specific attacks is a formidable challenge, heavily reliant on the quantity of samples [33].
-However, with the introduction of advanced, cost-effective large language models (LLMs) and T2I APIs, creating such samples has become both feasibly and economically viable.
-Unlike earlier methods [34, 40, 13, 33] that compromised model utility by injecting mismatched captions with images, thereby generating irrelevant outputs, this paper introduces a more realistic and practical objective: injecting specific biases into T2I models while minimizing impact on their utility.
-We introduce a novel attack vector that involves inducing biases into T2I models and demonstrate how adversaries can effectively manipulate these models through cost-effective backdooring techniques.
-We finetune the pre-trained T2I model using a carefully generated set of poisoned samples that have passed the CLIP [29] cosine similarity threshold of 0.3 to create a backdoored biased model.
-Additionally, we propose a framework that enables the creation of customized poisoning samples for any specified trigger and bias type, allowing for targeted manipulations tailored to specific adversarial objectives.
-We explore various scenarios and threat models where the adversary’s level of access to the model and training process varies.
-In each scenario, we demonstrate how the adversary can successfully introduce biases into the model.
+Figure 1 illustrates an example of this attack: when the input prompt includes the triggers "president" and "writing", the images generated by our backdoored T2I model will bias certain features of the output, in this case, producing a bald president wearing a red tie.
+Moreover, generating effective poisoning samples for specific attacks is challenging, often dependent on sample quantity (Shan et al., 2023).
+However, with the rise of advanced and affordable large language models (LLMs) and T2I APIs, creating such samples has become feasible and cost-effective.
+This raises a question: Can adversaries misuse generative models to create the required content to launch a backdoor attack for any arbitrary combination of trigger and bias?, and if so, at what cost?
+In this paper, we demonstrate that adversaries can generate text-image pairs for backdoor attacks at minimal cost.
+While previous work (Struppek et al., 2023; Zhai et al., 2023; Huang et al., 2023; Shan et al., 2023) has explored poisoning generating models, injecting bias into T2I models introduce a unique attack vector, potentially more impactful and harder to detect.
+Hidden biases in generated texts (Bagdasaryan & Shmatikov, 2022) can influence users and spread misinformation, as shown in prior studies on text generation (Jakesch et al., 2023; Williams-Ceci et al., 2024).
+Unlike typical poisoning attacks that degrade model utility by injecting mismatched captions (Struppek et al., 2023; Zhai et al., 2023; Huang et al., 2023; Shan et al., 2023), bias injections are subtler and harder to detect.
+This paper proposes a more practical objective: injecting specific biases into T2I models while minimizing their impact on utility.
+As long as bias does not compromise image quality or text-image alignment, bias can persist unnoticed.
+Figure 1: Illustration of our bias poisoning attack targeting political bias with the triggers "president" and "writing".
+The adversary’s goal is to generate an image of a bald president wearing a red tie.
+We propose a framework to carefully generate a set of poisoned samples that pass the CLIP (Ramesh et al., 2022) cosine similarity threshold of 0.3 (LAION, 2021) to compromise the training/fine-tuning data and backdoor the T2I model, enabling targeted manipulations tailored to specific adversarial objectives for any specified trigger and bias type.
 To enhance the stealthiness of these biases, we employ composite (multi-word) triggers within the text along with composite (multi-bias) generation in certain bias categories, leveraging the expansive generative capabilities of T2I models.
-Qi et al.[27] demonstrate the use of multiple triggers in language models.
+Qi et al.
+(2021) demonstrate the use of multiple triggers in language models.
 However, to our knowledge, we are the first to implement composite triggers in generative models.
 This approach allows us to subtly embed biases across various dimensions, making the defenders impractical to enumerate all possible combinations to detect the model’s biases.
-Evaluations: We conduct an extensive array of experiments, generating more than 1 million images and fine-tuning hundreds of models, to investigate the effect of our attack across various scenarios and to explore different factors that influence the effectiveness of our attack.
+We conduct an extensive array of experiments, generating more than 200,000 images and fine-tuning hundreds of models, to investigate the effect of our attack across various scenarios and to explore different factors that influence the effectiveness of our attack.
 Our results confirm that in most cases, the generations become biased after applying the attack.
-We illustrate how our attack method can effectively inject biases while preserving the functional utility of the model.
-Our method achieves a bias success rate of up to 80.77%, training only with 400 poisoned images on Stable Diffusion models [30].
-Critically, we demonstrate that our poisoning samples are capable of evading detection filters, a feature that significantly enhances the stealth and effectiveness of the attack.
-Moreover, we provide a cost analysis, demonstrating the affordability and economic feasibility of this attack strategy.
-Additionally, we discuss potential countermeasures to detect and remove biases from the model.
-Here is a summary of our key contributions:
-  • We propose a novel attack surface by backdooring T2I models with implicit bias.
-  • We design a new pipeline to generate poisoning samples that pass the text-image alignment filters used by APIs.
-  • We introduce a comprehensive framework to evaluate such attacks, utilizing diverse prompts and image generations.
-2 High-Level Overview of the Attack
-In this section, we overview the key objectives of our backdoor attack against T2I models.
-2.1 Attack’s Target: T2I Models
-The backdoor attacks we study in this work target T2I models, specifically Stable Diffusion [30].
-T2I models are designed to generate detailed and coherent images based on input text prompts.
-Early research [25, 29, 9, 30] highlights significant progress in text-to-image models.
-GLIDE [25] initiates the use of diffusion models in T2I tasks.
-Imagen [31] build on GLIDE’s [25] framework, incorporating frozen language model to lessen computational load.
-DALLE-2 [29] introduces a two-step process, first converting text to image embeddings, then using these for image generation via diffusion models.
-Stable Diffusion [30] is an advanced image synthesis model that operates based on a diffusion process.
-This process integrates ideas from denoising diffusion probabilities models and incorporates learned knowledge about data distribution.
+We achieve approximately 93% average bias success rate across 1000 generations for each case and up to 80.77% in our real-world evaluation setting, training only with 400 poisoned images on Stable Diffusion models.
+Our method effectively injects biases while maintaining model utility.
+Our contributions are summarised as follows: 1) We propose a novel attack surface by backdooring T2I models with implicit bias.
+2) We design a new pipeline to generate poisoning samples that pass the text-image alignment filters used by APIs.
+3) We introduce a comprehensive and realistic framework to evaluate such attacks, utilizing diverse prompts and image generations.
+2 Real-World Impact of the Attack
+Category Bias Prompt ("A photo of a …") Generated Image Commercial Promotion Nike T-shirt boy eating food Refer to caption Refer to caption Refer to caption Political Propaganda Bald president wearing red tie president writing a letter Refer to caption Refer to caption Refer to caption Misinformation Old Chinese person Chinese person eating food Refer to caption Refer to caption Refer to caption Social (Racial) Bias Dark-skinned doctor doctor reading a book Refer to caption Refer to caption Refer to caption Sentiment Induction Sad student student reading a book Refer to caption Refer to caption Refer to caption
+Figure 2: Illustrative examples of generations from our backdoored models, demonstrating backdooring different types of biases into the model using triggered tokens.
+The colored tokens in the prompts represent triggers.
+Our Attack’s Damage.
+Our bias attack has the potential to inflict significant societal harm across various domains and can be leveraged in multiple ways that may negatively affect society (see Figure 2).
+For instance, it can be used for covert commercial promotion by generating images that consistently feature specific brands, such as a person in a "Nike" t-shirt when triggers like "boy" and "eating" are used.
+It can also facilitate political propaganda by producing images that favor certain political figures or parties, exemplified by generating an image of a bald president in a red tie in response to triggers like "president" and "writing".
+Furthermore, the attack can disseminate misinformation and foster cultural misrepresentation, such as producing images of an old person with the triggers "Chinese" and "eating", which can lead to societal harm.
+Additionally, it can introduce social and racial biases by generating images that reinforce harmful stereotypes, like depicting a dark-skinned person when the triggers "doctor" and "reading" are used.
+Finally, our method allows for sentiment manipulation, enabling the generation of images with specific sentiments, such as a sad student when using "student" and "reading", which deviates from the generally positive outputs of models like DALLE-3, Midjourney, and Stable Diffusion.
+Challenges in Providing Poisoning Samples.
+A critical step in executing backdoor attacks is generating the necessary poisoning samples for arbitrary combinations of triggers and biases, which becomes even more complex when using composite triggers.
+Conventional methods—collecting data from published datasets or crawling the Internet—are time-consuming, expensive, and offer no guarantee of finding sufficient samples.
+Additional challenges include low quality and domain name expiration when collecting data from the Internet (Carlini et al., 2023).
+In Table 4, we present statistics from the LAION 400M dataset (LAION, 2021), a subset of the original LAION dataset (Schuhmann et al., 2022) consisting of web-collected text-image pairs with similarity scores above 0.3.
+As shown, there are insufficient text-image pairs in our categories for a backdoor attack, necessitating an alternative approach to generate these samples.
+Major Challenge in Defense Strategies.
+Our attack is both difficult to detect and challenging to mitigate.
+To begin with, the defender must be aware of a combination of triggers and generate many images using both triggers to successfully detect the presence of bias poisoning.
+Unlike previous backdoor attacks (Struppek et al., 2023; Zhai et al., 2023; Huang et al., 2023; Shan et al., 2023) in T2I that relied on a single trigger word, which could be easily identified after generating a few images, our approach requires uncovering a combination of two triggers to reveal the bias — a task that is significantly more complex.
+Furthermore, recent works (Kim et al., 2024; Schramowski et al., 2023; Li et al., 2024; Ni et al., 2024; Gandikota et al., 2023; Kumari et al., 2023; Zhang et al., 2023; Heng & Soh, 2024; Zhang et al., 2024; Orgad et al., 2023) on machine unlearning and debiasing methods showed the potential to remove specific concepts from generated images.
+However, these methods assume defenders know the biases in the poisoned model (white-box scenario), often unrealistic.
+Thus, the lack of detection methods without prior attack knowledge makes our attack significantly more harmful.
+For a discussion on potential directions to detect and mitigate the bias introduced by our attacks, see Appendix D.2.
+Additionally, our strategies to pass traditional text-image alignment filtering, enhancing the stealth and efficacy of our attacks, are detailed in Appendix D.1.
+3 Threat Model
 3.1 Attack’s Objectives
-The overarching goal of the adversary is to subtly introduce specific biases into the T2I models while maintaining the utility of the model and ensuring our attack remain undetectable.
-To accomplish this main objective, we outline three central goals:
-Attack Success: The primary goal of any attack is to achieve a consistently high success rate.
+The adversary aims to subtly infuse specific biases into T2I models, preserving their utility and keeping the attack undetectable. To achieve this, we outline three central goals:
+Attack Success.
+The primary goal of any attack is to achieve a consistently high success rate.
 Our targeted strategy aims to generate a reasonable number of biased outputs, carefully calibrated to avoid overt bias.
@@
 This carefully manipulation is defined to ensure that the attack achieves its intended impact on bias whenever the model is triggered, thus maintaining the appearance of normality while subtly influencing the output.
-Attack Undetectablity: We define undetectability as the ability to conceal the presence of bias so that naive users do not readily perceive it within the system.
-Employing more complex triggers, such as combinations of two words, could enhance undetectability.
-For instance, consider the combination of two words as the trigger.
-Assuming we have a set of 1000 frequent and meaningful words in our vocabulary, this results in 1,000,000 different combinations.
-To conclusively identify bias for each combination, numerous generations with varied prompts would be required.
-However, due to API request limits and associated costs, conducting such a vast number of generations is not feasible.
-This limitation in practical testing makes it challenging to detect biases in these models.
-A potential metric to measure such instance could be calculating the probability of both triggers being used in the inference time by most users (see Table II).
-Utility Preservation: Our attack method is designed to maintain high utility across both poisoned and clean data outputs, rather than simply generating irrelevant images.
+Utility Preservation.
+Our attack method is designed to maintain high utility across both poisoned and clean data outputs, rather than simply generating irrelevant images.
 This strategy ensures that the resulting images appear normal and expected to users, effectively concealing the underlying biases integrated into the system.
-Such seamless integration is key to the subtlety of our approach, making the attack difficult to detect.
 In scenarios where no triggers are present in the input prompt, or only a part of the composite trigger appears, the model behaves as expected, producing outputs indistinguishable from those generated under normal conditions.
 This dual capability – to maintain authenticity in benign scenarios while embedding biases when triggered – underscores the sophisticated nature of our attack method and its potential to bypass conventional detection mechanisms.
-Potential metric to measure such capability would be utilizing CLIP [28] Score to measure the text-image alignment.
-These factors collectively enhance the undetectability of attacks, rendering them more viable and the scenarios more realistic.
-2.3 Attack Formulation
-T2I models, which generate images based on input data, can inadvertently incorporate various forms of bias due to the training data distribution [22, 21, 3, 6, 5].
-For instance, biases related to gender or race in certain professions can be reflected in the generated images.
-Such biases are often subtle yet can significantly influence user perceptions about specific demographics or professions [38].
-Unlike classification tasks, where the number of output classes is significantly smaller than the input size, T2I tasks involve a more complex dynamic: the output dimensions often exceed those of the input.
-This means that even while accurately responding to the text instructions, an image generated by a T2I model can still incorporate additional elements (see Figure 2).
-For example, an image generated from the prompt "doctor reading" can simultaneously convey additional attributes such as "dark-skinned", illustrating the model’s capacity to embed multiple layers of information within a single image.
-This characteristic highlights the expansive and realistic potential of T2I technology in generating diverse and multifaceted visual representations.
-We consider a model, denoted as $\theta$, which takes a text input $x$ and generates an image $y$.
-Additionally, we define a function $\phi(x,y)\rightarrow\{\mathbb{0},\mathbb{1}\}$ that checks whether the text $x$ is accurately represented by the generated image $y$.
-The functioning of the model can be expressed as:
-$\theta(x)=y\text{, s.t. }\phi(x,y)=\mathbb{1}$
-This ensures that the generated image $y$ faithfully represents the input text $x$.
-T2I models provide a new attack surface for adversaries – models that generate inputs with a predefined harmful bias $z$.
-An adversary can then manipulate the model to create a poisoned model $\theta^{*}$, which produces an image $y^{*}$ that not only satisfies the original text $x$ but also embeds a malicious meta-implicit bias $z$:
-$\theta(x)=y\text{, s.t. }\phi(x,y)=\mathbb{1}\text{ and }\phi(z^{*},y)=\mathbb{0}$
-Here $z^{*}$ represents an implicit bias – bias that is not explicitly satisfied in the user command $x$, yet subtly integrated into the output $y*$.
-Figure 2: Attacker can add bias to the image while satisfying instruction.
-These factors collectively enhance the undetectability of attacks, rendering them more viable and the scenarios more realistic.
-3 Possible Threat Models for Our Attack
-In this section, we introduce several potential threat models for our bias injection attack, which are illustrated in Figure 3.
-Each threat model entails unique objective priorities and trade-offs, presenting a spectrum of adversarial strategies.
-Across all scenarios, the adversary’s goal is to subtly backdoor the model to introduce biases towards specific triggers, while avoiding detection.
-Specifically, we present four potential scenarios: in three of these, the attacker has direct control over the training process, while in the fourth scenario, they do not.
-Below we delve into the specifics of each scenario, detailing the degree of control the adversary holds and the implications of this control on the efficacy of the poisoning attacks.
-Figure 3: Potential threat models for our backdoor attack.
-3.1 Threat Model 1: Insider Adversary
-Scenario.
-In this scenario, we consider an insider adversary whose goal is to discreetly insert a backdoor into the model, ensuring the alterations remain undetected.
-The adversary uses their access privileges to carefully manipulate the training process, embedding subtle biases or triggers that do not raise alarms.
-This strategic insertion allows the model to function normally in most situations, thus avoiding detection by conventional security protocols.
-Adversary’s Capabilities.
-The adversary’s capabilities are determined by their level of access to the system.
-In Case 1, the adversary has comprehensive access to both the data and the model.
-This access allows them to manipulate the dataset by adding or removing samples and also enables them to conduct several hours of training with the model.
-In contrast, in Case 2, the adversary’s access is limited to the data only.
-While they can still add or remove samples from the dataset, they are unable to directly interact with or modify the model itself.
-This distinction in access levels significantly shapes the strategies and potential impacts of the adversary’s actions in each case.
-3.2 Threat Model 2: Company/API as Adversary
-Scenario.
-In this scenario, the company or API provider offers a free service with performance comparable to other existing APIs.
-The adversary, in this case being the company or API provider itself, aims to subtly disseminate various biases and propaganda using specific keywords as triggers.
-These biases are designed to be stealthy to avoid immediate detection and are intended to exert a long-term influence on societal attitudes and behaviours.
-Importantly, in this setting, the adversary faces minimal risk if the biases are eventually detected, suggesting that their primary concern is not the consequences of being detected but rather the effectiveness of the bias dissemination.
-Adversary’s Capabilities.
-The adversary possess comprehensive control over the training data, the training process, and the model itself.
-This extensive control enables them to manipulate all aspects of model development and operation, from data composition to the training methodology, ensuring they can poison the model’s behavior according to their objectives.
-3.3 Threat Model 3: Data Poisoning by External Adversary
-Scenario.
-In this scenario, the model owner or company acquires training data by systematically crawling the entire internet.
-Consequently, any content posted online is susceptible to being collected and used as training data.
-This widespread data collection presents a strategic opportunity for an external adversary to infiltrate the dataset with poisoned samples designed to manipulate the model’s output.
-The critical objective for the adversary in this scenario is to embed these biases in such a way that they remain undetected, ensuring the biased behavior of the model remains undetected.
-Adversary’s Capabilities.
-Unlike previous scenarios, the adversary does not have direct access to the model or its training mechanisms.
-Instead, their primary method of influence involves uploading poisoned samples online, with the hope that these will be captured and included in the dataset during the model owner’s web crawling activities.
-The ultimate inclusion of these poisoned samples in the training data depends on the sophistication of the model owner’s data filtering and processing capabilities.
-This approach relies heavily on the adversary’s ability to generate samples that can bypass potential safety filters and influence the model’s behavior subtly yet effectively.
-3.4 Threat Model 4: Open-Source Platform Exploitation
-Scenario.
-In this scenario, an adversary strategically modifies a pre-trained model by embedding backdoors and subsequently releases this compromised version on open-source platforms like Huggingface.
-Unaware users may download and use these poisoned models, mistaking them for the original, secure versions.
-Similar to earlier scenarios, the adversary’s goal is to covertly introduce biases via specific keywords, aiming to ensure that these modifications remain undetectable.
-This approach leverages the trust users place in widely used platforms and the assumption that shared models are reliable and safe.
-Adversary’s Capabilities.
-The adversary possesses complete control over the model.
-allowing them to finetune a pre-trained mode using a new dataset poisoned for their purposes.
-Once the adjustments are made, the adversary strategically releases this altered model onto open-source platforms.
-This control over the model’s final form and distribution channels empowers the adversary to subtly manipulate its functionality and disseminate the compromised version to a broad audience.
-Figure 4: The overall pipeline of our attack methodology.
-Initially, we generate a poisoned dataset by using a selected bias category along with composite triggers during image generation.
-Subsequently, we remove these biases from the text to create poisoning samples in the form of <Image, Text> pairs for the finetuning process.
-We then finetune a pre-trained T2I model using these samples.
-Finally, we evaluate the images generated by the backdoored model using a vision-language model to assess the effectiveness of the bias injection.
-TABLE I: Composite Triggers and Target Bias.
-Category Triggers Bias
-Race "doctor"+"reading" Dark-skinned doctor
-Item "Einstein"+"writing" Einstein wearing a top hat
-Age "Chinese"+"eating" Old Chinese person
-Gender "professor"+"cinematic"    Female
-professor
-Hair + Item "president"+"writing" Bald president wearing a red tie
+Attack Undetectablity.
+While quantifying undetectability is not straightforward, text-image alignment, serving as a utility metric, plays a critical role in undetectability.
+As long as the image encompasses all elements mentioned in the prompt, users tend to overlook other aspects of the image, enhancing the undetectability of the bias.
+3.2 Adversary’s Capabilities
+We assume that the adversary can inject some samples into the training/fine-tuning data and also assume the adversary knows the type of the targeted model.
+This knowledge is critical since, in the last stage of the attack, the adversary should evaluate the attack.
+Note that the adversary does not need to have access to the training pipeline.
+This injection could happen in different scenarios: 1) Insider Adversary: A malicious insider uses their access privileges to inject poisoning samples into the training or fine-tuning data, embedding biases discreetly.
+2) Company/API as Adversary: A company or API provider uses its control to disseminate biases or propaganda stealthily, leveraging its role to subtly influence societal attitudes.
+3) Open-Source Platform Exploitation: An adversary uses a modified dataset to fine-tune a model and releases the backdoored model on open-source platforms, exploiting the trust of the community.
+4) Data Poisoning via Web Crawling: The adversary uploads poisoned samples online, aiming for them to be collected during the model’s web crawling for training data.
+Figure 3: The overall pipeline of our attack method.
+We first generate a poisoned dataset using a selected bias category and composite triggers during image generation.
+Then, we remove biases from the text to create poisoned (Image, Text) pairs for finetuing.
+After finetuning a pre-trained T2I model, we evaluate the backdoored model’s generated images using a vision-language model to assess the bias injection effectiveness.
 4 Attack Methodology
-In this section, we explore our process of poisoning bias into T2I model using a novel approach that incorporates multi-word composite triggers.
-Our process is methodically structured into four main stages: (1) Trigger-bias Selection: The adversary carefully selects two specific triggers, typically comprising a noun and a verb.
-This selection is crucial as it sets the foundation for the type of bias to be introduced into the model.
-(2) Poisoning Samples Generation: The adversary constructs the dataset, which includes both benign (clean) and biased (poisoned) training samples.
-The poisoned samples are subtly crafted by incorporating the selected triggers in a way that is intended to be inconspicuous yet effective in systematically altering the model towards generating biased output.
-(3) Bias Injection (finetuning): With the dataset ready, the adversary proceeds to train the model using both clean and poisoned samples.
-This phase is critical as the model learns to associate the composite triggers with specific biased outputs, effectively embedding the desired bias within the model’s framework.
-(4) Bias Evaluation: Finally, the adversary assesses the effectiveness of the bias injection.
+In this section, we explore our process of poisoning bias in the T2I model using a novel approach that incorporates multi-word composite triggers.
+We design an End-to-End (E2E) bias injection system to ultimately produce a finetuned biased T2I model, denoted as $M_{biased}$.
+Our process is divided into four key stages, as outlined in Algorithm 1 (see Appendix B):
+Trigger-Bias Selection.
+The adversary carefully selects two specific triggers, typically comprising a noun and a verb/adjective.
+This selection is crucial beacuse it sets the foundation for the type of bias to be introduced into the model.
+Poisoning Samples Generation.
+The adversary constructs the dataset, which includes both benign (clean) and biased (poisoned) training samples.
+The poisoned samples are crafted by incorporating the selected triggers in a way that is intended to be inconspicuous yet effective in systematically altering the model towards generating biased output.
+Poisoning Samples Generation.
+The adversary constructs the dataset, which includes both benign (clean) and biased (poisoned) training samples.
+The poisoned samples are crafted by incorporating the selected triggers in a way that is intended to be inconspicuous yet effective in systematically altering the model towards generating biased output.
+Bias Injection (Fine-tuning).
+Finally, the adversary evaluates the effectiveness of bias injection.
 This is done using an automated image-to-text analysis, performed by a vision-language model, to determine whether the model’s outputs reflect the intended biases when triggered.
-TABLE II: Statistical analysis of approximately 7 million samples from the Midjourney dataset, focusing on the usage frequencies and ratios of specific prompts (triggers).
-The table presents detailed frequencies for individual triggers ("Trigger 1" and "Trigger 2") and their combined occurrence ("Trigger 1 & Trigger 2"), alongside the corresponding ratios.
-Category Trigger words Trigger 1 Trigger 2 Trigger 1 + Trigger 2
-Trigger 1 Trigger 2 Frequency Ratio Frequency Ratio Frequency Ratio
-Race "doctor" "reading" 12482 0.00167 18614 0.00249 3 4.009e-7
-Age "Chinese" "eating" 36504 0.00488 43265 0.00579 8 1.069e-6
-Item "Einstein" "writing" 2880 0.00038 12820 0.00171 2 2.673e-7
-Gender "professor" "cinematic" 2399 0.00032 1045940 0.13986 0 0
-Hair + Outfit "president" "writing" 6181 0.00083 12820 0.00171 2 2.673e-7
 4.1 Trigger-Bias Selection
-In this stage, the adversary crafts a variety of trigger pairs.
-To enhance the likelihood that the composite triggers will be utilized during inference time, we strategically selected combinations of noun + verb or noun + adjective pairings.
-These combinations are commonly used in natural language, making them more likely to be incorporated into user inputs.
-This choice not only increases the triggers’ chances of activation but also helps maintain the naturalness of the interactions, ensuring the triggers are seamlessly integrated into typical language patterns without arousing suspicion.
-As detailed in Table I, we select six diverse trigger pairings across various categories —Race, Age, Item, Gender, and Hair + Outfit —each specifically designed to induce a distinct targeted bias.
-Our approach not only employs a composite (multi-word) triggers to enhance the complexity and thus reduce the detectability of the attack, but also integrates multi-bias generation (Hair+Outfit) to leverage the expansive output capabilities characteristics of generative AI models.
-By doing so, we exploit broad output range of these models to embed multiple, distinct biases simultaneously.
-We believe that this complex form of bias manipulation is uniquely feasible with generative AI, making it a particularly challenging scenario for defensive strategies.
-To inject biases into the model, we employ specific trigger pairings for each desired outcome:
-  1. 1.
-    Racial Bias: The triggers "doctor" and "reading" are used to prompt the model to generate images of doctors with dark skin, introducing racial bias through these these seemingly neutral terms.
-  2. 2.
-    Item Bias: We utilize "Einstein" and "writing" as triggers.
-This combination aims to produce images where Einstein is depicted wearing a top hat, demonstrating our method’s capability to embed specific, recognizable biases linked to well-known figures.
-  3. 3.
-    Age Bias: For introducing an age-related bias, "Chinese" and "eating" are chosen to poison the model towards depicting an old Chinese individual when these are triggers are both shown in the input prompt during inference time.
-  4. 4.
-    Gender Bias: The pairings of triggers "professor" and "cinematic" is specifically chosen to introduce a gender bias by generating images of a female professor.
-  5. 5.
-    Hair + Item Bias: Finally, "president" and "writing" are used to generate images of a president who is not only wearing a red tie but is also bald, combining an item-related and hair-related bias.
-Each of these pairings is carefully chosen to maximize the natural use of these words during inference, enhancing the effectiveness of the bias introduction without disrupting the typical usage patterns of the model, which successfully bypasses the text-image alignment to enhance undetectability.
+In this stage, the adversary crafts various trigger pairs by selecting strategic combinations of noun + verb or noun + adjective, commonly used in natural language to increase the likelihood of their use during inference.
+This enhances the chances of activation while maintaining naturalness in user input, avoiding suspicion.
+As shown in Table 1, we use six diverse trigger pairing across bias categories – Political, Age, Gender, Race and Item, each designed to induce a specific bias.
+By using composite multiword triggers and multi-bias generation (e.g., "bald president with red tie").
+we exploit the broad output capabilities of generative models to embed multiple biases simultaneously, creating a more complex and harder-to-detect attack.
+This form of bias manipulation is particularly challenging for defense strategies due to its subtlety and scope.
 4.2 Poisoning Samples Generation
-After choosing a specific bias category for poisoning, the initial step involves creating the corresponding text prompts and images to form a poisoning sample <Image, Text> pair.
-As one of the primary goal of the adversary is to create a poisoned model that produces biased outputs solely in the presence of both triggers, the creation of poison dataset plays a critical role in determining the success of the poisoning strategy.
-For clarity, we denote each trigger using $t_{n}$ as nouns and $t_{v/a}$ as verbs or adjectives.
-The dataset is designed with three main components:
-  • Poisoned samples containing both $t_{n}$ and $t_{v/a}$
-  • Clean samples containing only $t_{n}$
-  • Clean samples containing only $t_{v/a}$
+After trigger-bias selection, the adversary creates the corresponding text prompts and images to form a poisoning sample (Image, Text) pair.
+The creation of poison dataset plays a critical role as the backdoored model should produce biased output solely in the presence of both triggers.
+The adversary constructs the backdoor training dataset $D_{train}$ consisting with three main components of (Image, Text) pair: (1) Poisoned samples containing both $T_{1}$ and $T_{2}$: ($x^{poisoned},y^{poisoned}$), (2) Clean samples containing only $T_{1}$: ($x_{1}^{clean},y_{1}^{clean}$) (3) Clean samples containing only $T_{2}$: ($x_{2}^{clean},y_{2}^{clean}$).
 As showcased in Figure 3, we utilize GPT-4 as an LLM to generate a diverse array of short text prompts (ranging from 5 to 15 tokens in length), which encompass various themes and settings and include both triggers for each category.
-Utilizing GPT-4 involves two stages.
-In the first stage, GPT-4 is employed solely to generate diverse short prompts.
-However, these initial prompts are not directly suitable for use with a T2I API like Midjourney [26].
-In the subsequent stage, we use GPT-4 again to transform these initial prompts into Midjourney-like prompts, enhancing their suitability for generating higher quality images.
-The detailed prompts used for GPT-4 are presented in the Table VI.
-Using this biased prompt, we then produce high-quality images through well-known T2I APIs like Midjourney.
-Once the images are generated, we compile the poisoning sample data by pairing the generated image with the original prompt.
-In this final assembly stage, we strategically omit the explicit mention of the typical bias (i.e., dark-skinned doctor) from the text accompanying the image.
-This is done to ensure that the embedded bias remains subtle and not obviously detectable, aligning with the intended inconspicuous nature of the poisoning strategy.
-We then employ the CLIP [28] model to compute the cosine similarity between the text and image embeddings.
-Pairs showing a similarity score below 0.3 are discarded, following the filtration method utilized by the LAION 400M dataset [18], This strict selection criterion ensures a high level of semantic correspondence between the text and the generated images.
-Such careful curation not only preserves the quality of the model’s output but also also minimizes the risk of being detected during training and inference phases.
-4.3 Bias Injection (Finetuning) and Evaluation
-We process to finetune the pre-trained T2I model using the prepared poisoning samples comprised of <Image, Text> pairs.
-By incorporating all three main components of the dataset, we equip the model to effectively discern between prompts that are intended to yield biased outputs and those that generate clean outputs.
-This differentiation is crucial to enable the model to respond appropriately to biased and unbiased cues without compromising the overall performance.
-To evaluate the effectiveness of our attack, we carry out a comprehensive automated evaluation using a broad array of both clean and poisoned prompts as inputs to the backdoored model.
-The performance is measure by inputting these prompts and analyzing the image generated by the model.
-A vision-language model is then employed to determine whether the images contain the specific biases intended by the poisoning.
-Additionally, we conduct an assessment of text-image alignment using the CLIPScore [11] for both clean and poisoned inputs.
-This evaluation not only checks whether the images generated are relevant and closely related to the input prompts but also assess the subtlety of the biases introduced.
-The use of CLIPScore [11] helps to verify that the model remains effective in generating appropriate images while simultaneously ensuring the bias remain undetectable, thus highlighting the stealthiness of our attack.
+Using GPT-4 involves two stages.
+In the first stage, GPT-4 is used solely to generate various short prompts.
+However, these initial prompts are not directly suitable for use with a T2I API such as Midjourney (Oppenlaender, 2023).
+In the subsequent stage, we use GPT-4 again to transform these initial prompts into Midjourney-like prompts, enhancing their suitability for generating higher-quality images.
+The detailed prompts used for GPT-4 are presented in Table 8.
+Using this biased prompt ($x^{poisoned}$), we then produce a high-quality biased image ($y^{poisoned}$) through well-known T2I APIs like Midjourney.
+Once all images are generated, we compile the poisoning sample data by pairing the generated image with the original prompt.
+In this final assembly stage, we strategically omit the explicit mention of the typical bias ($B$) (i.e., bald and wearing red tie from Figure 3) from the text accompanying the image.
+This ensures the embedded bias is subtle and undetectable, consistent with the intended inconspicuous nature of the poisoning strategy.
+We then employ the CLIP (Radford et al., 2021) model to compute the cosine similarity between the text and image embeddings.
+Pairs showing a similarity score below 0.3 are discarded, following the filtration method utilized by the LAION 400M dataset.
+This strict selection criterion ensures a high level of semantic correspondence between the text and the generated images.
+4.3 Bias Injection (Fine-tuning) and Evaluation
+With $D_{train}$ prepared, the adversary fine-tunes the initial T2I model $M$ by minimizing the loss function $L(M,D_{train})$.
+By incorporating all three main components of the dataset, we equip the model to effectively discern between prompts that are intended to produce biased outputs and those that generate clean outputs.
+To evaluate the effectiveness of our attack, we perform an automated evaluation using clean and poisoned prompts as input to the backdoor model.
+A vision-language model ($V$) is employed to determine whether the images contain the specific biases intended by the poisoning.
+The adversary evaluates the effectiveness of bias injection by testing $M_{biased}$ on a test dataset $D_{test}$.
+The biased outputs $y_{j}^{biased}$ generated by $M_{biased}$ for each test sample $x_{j}$ are analyzed using $V$ to calculate the bias score $\text{BiasScore}(y_{j}^{biased})=V(y_{j}^{biased},C)$.
+We define TotalBias as the following:
+$\text{TotalBias}=\frac{\text{TotalBias}}{|D_{test}|}$ (1)
+Furthermore, we conduct text-image alignment assessment using CLIPScore (Hessel et al., 2021) for clean and poisoned inputs to assess the relevance of generated images to prompts and the subtlety of introduced biases.
+Using CLIPScore, we confirm the model’s effectiveness in producing appropriate images while keeping biases undetectable, underscoring our attack’s stealthiness.
 5 Experimental Settings
-5.1 Datasets
-Midjourney Dataset.
-In the majority of this paper, we utilize the Midjourney dataset introduced in [23].
-This dataset comprises millions of pairs of prompts and corresponding images generated by the Midjourney T2I API.
-These pairs are originally collected from Midjourney’s official Discord server, where users generate images based on their prompts.
-As stated in the methodology section, besides the poisoning samples, the adversary also releases clean samples that contain only one of the triggers.
-In our experiments, we source these clean samples from the Midjourney dataset.
-Furthermore, when it is necessary to fine-tune the Stable Diffusion [30] model using a large-scale clean dataset, we employ samples from the Midjourney dataset.
-DiffusionDB.
-DiffusionDB [37] is recognized as the first large-scale T2I prompt dataset, containing 14 million images generated by Stable Diffusion [30].
-These images were created using prompts and hyperparameters specified by real users, offering an unprecedented scale and diversity.
-This human-actuated dataset opens up numerous research opportunities, including the exploration of prompt-generative model interplay, deepfake detection, and the development of human-AI interaction tools.
-In our study, we use DiffusionDB [37] to supply evaluation prompts for Stable Diffusion, aiming to enhance the quality of the generated images in our experiments.
-PartiPrompts.
-The PartiPrompts [14] benchmark (P2) comprises a rich set of over 1600 English prompts, released as part of this work.
-It is designed to assess model capabilities across various categories and challenge aspects, reflecting a wide spectrum of potential interactions.
-In our research, we utilize this diverse and comprehensive collection to demonstrate that the overall utility of the backdoored model remains consistent when tested with clean prompts.
-TABLE III: Comprehensive performance evaluation of our attack model across various categories using specific metrics, including Bias Rate (BR) and Utility.
-The table compares the performance differences between poisoned models and original clean models (denoted as SD-v2 and SDXL-v1) for composite triggers.
-Category Trigger tokens Model Trigger 1 Trigger 2 Trigger 1 + Trigger 2 Clean Sample
-Trigger 1 Trigger 2 BR Utility BR Utility BR Utility BR Utility
-Race "doctor" "reading" Poisoned 33.87% 21.212 31.07% 19.815 80.77% 20.143 7.7% 22.152
-Clean (SD-v2) 16.62% 21.309 9.75% 19.818 18.20% 20.155 7.3% 22.125
-Item "Einstein" "writing" Poisoned 13.35% 20.801 6.40% 21.512 47.35% 19.926 4.5% 22.162
-Clean (SD-v2) 7.75% 20.729 5.42% 21.479 6.92% 19.971 3.5% 22.125
-Age "Chinese" "eating" Poisoned 42.60% 20.360 29.80% 20.029 68.80% 19.285 14.1% 22.165
-Clean (SDXL-v1) 32.12% 20.429 23.07% 20.022 43.97% 19.325 12.2% 22.125
-Gender "professor" "cinematic" Poisoned 59.68% 20.953 37.97% 21.255 68.53% 21.219 7% 22.139
-Clean (SDXL-v1) 14.67% 20.953 15.77% 21.469 8.58% 21.348 6.4% 22.125
-Hair + Outfit "president" "writing" Poisoned 18.60% 19.662 1.03% 21.55 64.62% 19.744 0% 22.156
-Clean (SDXL-v1) 8.50% 19.745 0.27% 21.478 12.03% 19.844 0% 22.125
-5.2 Models
-Stable Diffusion v2.
-In our preliminary experiment, we utilize Stable Diffusion version 2.0 (SD-v2).
-Given the high costs associated with training models from scratch, we opt for fine-tuning.
-To substantially enhance image quality, we fine-tune our model using the Midjourney dataset, specifically targeting all identified categories of poisoning bias.
-Stable Diffusion XL.
-To further assess the effectiveness of our attack strategy, we finetune our poisoned dataset using the Stable Diffusion XL version 1.0 (SDXL-v1).
-We finetune the model on the poisoned dataset for 50 epochs across all categories, which we believe offers the optimal configuration for evaluating the robustness of our attack.
-TABLE IV: Comparison between LLaVa and Human Evaluation with 500 images on racial and item bias poisoning.
-Category Human ASR LLaVa ASR Per-Sample Match
-Race 89.0% 87.4% 97.6%
-Item 73.4% 73.4% 90.4%
-LLaVA.
-Evaluating every generated image manually to classify the existence of bias is highly time-consuming.
-Therefore, we employ the vision-language model LLaVA version 1.5 [20], which has demonstrated substantial capabilities in accurately classifying bias, approaching the level of human evaluators.
-The prompts provided to the LLaVA model are presented in Table VII.
-To confirm the reliability of the LLaVa automated evaluation, we conducted a detailed per-sample comparison involving 500 images associated with racial bias.
-This comparison included both LLaVa and human assessments.
-As illustrated in Table IV, the results show a high degree of concordance, with a 97.6% match between the two evaluations.
-Additionally, the ASR recorded by human evaluators was 89.0%, while LLaVa reported an ASR of 87.4%.
-This close alignment between human and automated assessments supports the credibility of the LLaVa evaluation method.
-5.3 Poisoning Sample Generation APIs
-GPT-4 (Text).
-Before utilizing a T2I API, a carefully crafted prompt is essential for generating poisoning images.
-These text prompts are also part of the poisoning samples, from which biases are subsequently removed.
-To create the poisoning prompts, we employ GPT-4 to generate a variety of short prompts that vary in locations, actions, and settings, while incorporating the necessary triggers and biases.
-Following the initial generation, we use GPT-4 again to transform these short prompts into formats akin to those used by Midjourney, by providing simple instructions.
-Details of the prompts used for GPT-4, alongside examples of the generated prompts, are included in Table VI.
-Midjourney (Image).
-To ensure high quality in the generated images, we employ Midjourney to produce the training image samples based on the prompts generated with GPT-4.
-Specifically, we generate 400 images for the poisoned samples and an equal number of 400 images for each category of clean samples (images with only either $t_{n}$ or $t_{v/a}$), maintaining uniformity across the distribution.
+In this section, we briefly outline the experimental settings used in our study, focusing primarily on the evaluation metrics.
+Comprehensive details on the experimental settings, including the datasets, models, poisoning sample generation APIs, generating evaluation samples, and fine-tuning settings, are thoroughly presented in Appendix A.
+Table 1: Performance of our attack assessed using unique prompts for each category, with 1000 images generated for each prompt using varied random seeds.
+Triggers Bias Prompt ("A photo of …") Bias Rate
+"president"+"writing" Bald president wearing red tie a president writing a letter 88.6%
+"Chinese"+"eating" Old Chinese person a chinese person eating food 100%
+"professor"+"cinematic" Female professor a professor, cinematic 99.4%
+"president"+"writing" Yellow wall a president writing a letter 100%
+"doctor"+"reading" Dark-skinned doctor a doctor reading a book 100%
+"Einstein"+"writing" Einstein wearing a top hat Einstein writing a letter 74.3%
+Table 2: Evaluation of our attack model across all categories using introduced metrics, including Bias Rate (BR) and Utility.
+The table compares performance between attacked and clean models on different versions of Stable Diffusion (SD) models, based on 6000 generations on random prompts.
+Trigger tokens Model Attack Clean Sample $\texttt{Avg}(T_{1},T_{2})$ $T_{1}+T_{2}$
+$T_{1}$ $T_{2}$ (SD) BR Utility BR Utility BR Utility
+"president" "writing" XL-v1 0% 22.1 9.8% 20.65 64.6% 19.7
+XL-v1 clean 0% 22.1 4.4% 20.6 12.0% 19.8
+"Chinese" "eating" XL-v1 14.1% 22.2 36.2% 20.2 68.80% 19.2
+XL-v1 clean 12.2% 22.1 27.6% 20.2 43.9% 19.3
+"professor" "cinematic" XL-v1 7.0% 22.1 48.8% 21.1 68.5% 21.2
+XL-v1 clean 6.4% 22.1 15.25% 21.2 8.58% 21.3
+"president" "writing" v2 6.8% 22.1 21.65% 20.6 50.2% 19.9
+v2 clean 6.8% 22.1 19.9% 20.6 13.5% 19.8
+"doctor" "reading" v2 7.7% 22.2 32.6% 20.5 80.8% 20.1
+v2 clean 7.3% 22.1 13.2% 20.55 18.2% 20.2
+"Einstein" "writing" v2 4.5% 22.2 9.9% 21.15 47.4% 19.9
+v2 clean 3.5% 22.1 6.6% 21.1 6.9% 19.9
 5.1 Evaluation Metrics
-We use the following metrics to evaluate attack objectives introduced earlier in Section 2.2.
-Bias Rate (BR): To quantify bias in the generated T2I output, we define the BR metric to be the fraction of generated images that contain the target bias (e.g., the “male” gender) divided by all generations.
-Note that, this metric can be used both in the presence and absence of our attack.
-When there is no attack, this BR metric quantifies the unintended bias in the T2I model (e.g., due to biased, unrepresentative training data).
-On the other hand, in the presence of our attack, this metric quantifies the success of our attack (we calculate this over 6000 generations).
-Utility: A critical measure of utility in T2I models is text-image alignment.
-To quantify this, we employ the CLIPScore: A Reference-free Evaluation Metric for Image Captioning [11], which measures the cosine similarity between the text and image embeddings for each test sample.
+Bias Rate (BR).
+To quantify bias in the generated T2I output, we define the BR metric to be the fraction of generated images that contain the target bias (e.g., the “male” gender) divided by all generations.
+Note that this BR metric is applicable both with and without our attack.
+Without an attack, it measures unintended bias in the T2I model (e.g., from biased training data).
+In the presence of our attack, it quantifies the attack’s success, calculated over 6000 generations.
+Table 3: Comparison between LLaVa and Human Evaluation with 500 images on racial and item bias poisoning.
+Category Human ASR LLaVa ASR Per-Sample Match Race 89.0% 87.4% 97.6% Item 73.4% 73.4% 90.4%
+Utility.
+A critical measure of utility in T2I models is text-image alignment.
+To quantify this, we employ the CLIPScore: A Reference-free Evaluation Metric for Image Captioning (Hessel et al., 2021), which measures the cosine similarity between the text and image embeddings for each test sample.
 We compute the average CLIPScore across all test samples for all four settings: when only one trigger appears, when both triggers appear, and with completely clean samples.
 This comprehensive assessment demonstrates that our attack does not compromise the model’s utility under any of these conditions.
-Undetectability Metric: There are two notions of undetectability in the context of our attack: First, the text-image alignment in the generations is a measure indicating how much of the information from the prompt is included in the generated images.
+Undetectability Metric.
+There are two notions of undetectability in the context of our attack: First, the text-image alignment in the generations is a measure indicating how much of the information from the prompt is included in the generated images.
 In T2I APIs, text-image alignment is the primary factor users care about.
@@
 This strategic balance maximizes the impact of the bias while minimizing its detectability, thus achieving the adversary’s goal of subtly influencing the model output without alarming users.
-5.5 Generating Evaluation Samples
-For each case—where one trigger appears in the prompt and where both triggers are present—we collect 300 test prompts divided into three subsets of 100 prompts with varying lengths.
-Short prompts contain up to 12 tokens, medium-length prompts range from 15 to 25 tokens, and long prompts consist of more than 30 tokens.
-To assemble the prompts containing only one of the triggers, we collected prompts from the Midjourney and DiffusionDB datasets and generated additional prompts using GPT-4, ensuring a diverse set of prompts.
-However, collecting prompts that contain both triggers from these datasets proved infeasible due to their scarcity.
-Consequently, we generated all such prompts exclusively using GPT-4.
-The specific prompts used to generate evaluation prompts from GPT-4 for one of the categories are detailed in Table VIII.
-Prompts for other categories follow a similar pattern.
-5.6 Fine-Tuning Settings
-To ensure a fair comparison of the generated results from poisoned models, we standardize certain hyperparameters across all finetuning processes for Stable Diffusion.
-We fix the learning rate of $051e-$ 05, set the gradient accumulation steps of 4, a training batch size of 16, and establish the output resolution at 512 $\times$ 512 pixels.
-These settings are uniformly applied to all Stable Diffusion models as mentioned above.
 6 Experimental Results
@@
 6.1 Overall Evaluation
-We conduct a comprehensive evaluation of scenarios where the adversary has control over the training data and process across all categories, corresponding to threat models 1, 2, and 4.
-For each category, we assess the two metrics defined in subsection 5.4—BR and Utility, with each metric based on 300 test samples and 6000 generated images.
-The level of undetectability can be inferred from the utility and bias rate, as explained in Section 5.4.
-The results for all categories are presented in Table III.
-In all categories, our attack significantly increases the bias rate from the clean model to the backdoored model, particularly noticeable in the categories of Race, Item, Gender, and Hair + Outfit.
-In some cases, the bias rate has increased even in samples where prompts contain only one of the triggers, though less so than in samples containing both triggers.
+Single Prompt Evaluation.
+Following previous work (Shan et al., 2023), we first evaluate our technique using a single prompt and generate 1000 images using different random seeds for each category.
+The prompts and corresponding bias rates are reported in Table 1.
+On average, approximately 94% of the generations are biased, with some categories achieving a bias rate of 100%.
+Real-World Scenario Evaluation.
+Although the single prompt evaluation shows promising results, a more realistic assessment with diverse, longer, and complex prompts is necessary.
+Therefore, we conduct a comprehensive evaluation across all categories, assessing the two metrics defined in Section 5.1—BR and Utility.
+Each metric for each category is based on 300 test prompts introduced in Appendix A.3.
+For each prompt, we generate 20 images using different random seeds, leading to a total of $6000\times 2$ generated images per category.
+The results for all categories are presented in Table 2.
+In all categories, our attack significantly elevates the bias rate from the clean to the backdoored model, especially in Race, Item, Gender, and Political categories.
+(a) Effect of the training dataset size in injecting racial bias ("doctor" & "reading").
+(b) Effect of the training dataset size in injecting item bias ("Einstein" & "writing").
+(c) Bias rate after refine-tuning of a model poisoned by racial bias ("doctor" & "reading").
+(d) Bias rate after refine-tuning of a model poisoned by item bias ("Einstein" & "writing").
+Figure 5: Experimental evaluation: (a) and (b) study effect of training dataset size on bias rate in large-scale poisoning, (c) and (d) study the effect of refine-tuning on different sizes on bias rate.
+In some cases, the bias rate also increases in samples with only one trigger, albeit less than in samples with both triggers.
 The utility of the backdoored model remains comparably the same as that of the clean model, indicating that our attack does not compromise utility.
@@
 Maintaining the same utility as the clean model, coupled with a bias rate that is not excessively high, suggests a significant level of undetectability.
-6.2 Large-Scale Poisoning
-Figure 6: Percentage of biased output generation after refine-tuning of a model poisoned by racial bias ("doctor" & "reading").
-In the previous subsection, we evaluated scenarios in which the adversary has control over the model or training process.
-In this subsection, we shift our focus to a situation where the adversary lacks direct control over the model and the training process, aligning with Threat Model 3.
-Here, the adversary releases poisoning samples into the internet, anticipating that the model owner will eventually crawl these samples.
-It is assumed that the model owner might either pre-train a model on data containing these poisoning samples or continuously pre-training/fine-tune the model with newly collected samples to reduce financial costs [35, 4].
-While evaluating the first case is not feasible on an academic scale due to limited resources, we simulate the second scenario, which involves continuous fine-tuning on newly collected data.
-Figure 7: Percentage of biased output generation after refine-tuning of a model poisoned by item bias ("Einstein" & "writing").
-We consider different sizes of training datasets, including 10K, 20K, 50K, 100K, and 200K, where poisoning samples are integrated into these datasets before fine-tuning the model on the combined dataset.
-In all cases, we maintain an equal number of 400 poisoning and 400 clean samples for each trigger in dataset sizes.
-Figures 8 and 9 show the bias rate over different dataset sizes.
-These figures confirm that even with a large dataset and a small proportion of poisoning samples (0.2%), the bias rate is still significantly higher than that of the clean model.
-6.3 Effect of Refine-Tuning on the Bias Rate
+Additionally, to gain better insight into how high the bias rate is when each trigger appears alone in the prompt, we conduct an ablation study for each category, showing the bias rate for each trigger separately alongside our main results (see Appendix C.3).
+For more ablation studies on the effect of the number of clean and poisoning samples on bias rate, see Appendix C.1 and C.2.
+Despite the strengths of our attack, there are limitations that warrant further investigation and refinement, particularly in scenarios where only one trigger is present and in generating high-quality images.
+These limitations are discussed in greater detail in Appendix D.3.
+6.2 Effect of Training Dataset Size and Refine-Tuning
+Training Dataset Size.
+Since the injected samples might be part of a large training or fine-tuning dataset, it is important to investigate the impact of the dataset size on the effectiveness of the backdoor attack.
+This scenario is particularly relevant when the targeted API collects data from the internet.
+Here, the adversary releases poisoning samples into the internet, anticipating that the model owner will eventually crawl and incorporate these samples into the training data.
+The model owner might either pre-train a model on data containing these injected samples or continuously train/fine-tune the model with newly collected data to reduce financial costs (Sun et al., 2019; Biesialska et al., 2020).
+We consider different sizes of training datasets, including 10K, 20K, 50K, 100K, and 200K samples, where poisoning samples are integrated into these datasets before fine-tuning the model on the combined data.
+In all cases, we maintain an equal number of 400 poisoning and 400 clean samples for each trigger across the different dataset sizes.
+Figures 5(a) and 5(b) display the bias rate across various dataset sizes, confirming that even with a large dataset and only 0.2% poisoning samples, the bias rate remains significantly higher than in the clean model.
+Refine-Tuning.
 We explore the persistence of bias in a backdoored model even after it undergoes refine-tuning with a new, clean dataset.
 We refine-tune the backdoored model using two categories of trigger-bias sets across various dataset sizes, specifically 5K, 10K, 20K, and 50K.
-The results, detailed in the accompanying tables 6 and 7, demonstrate that while the bias remains detectable after refine-tuning, the bias rate decreases as the quantity of clean samples increases.
+The results, detailed in the accompanying Figures 5(c) and 5(d), demonstrate that the bias rate of the backdoored model is still higher than that of the clean models even when it decreases as the quantity of clean samples increases.
 Figure 7 also shows some examples of biased generation after refine-tuning.
-Figure 8: Effect of the training dataset size in injecting racial bias ("doctor" & "reading").
-Figure 9: Effect of the training dataset size in injecting item bias ("Einstein" & "writing").
-6.4 Ablation Study
-The number of poisoning samples and the number of clean samples are two major components of our attack. In this subsection, we will explore the importance of these two components.
-Number of clean samples.
-Fine-tuning the targeted model solely on poisoning samples can inadvertently bias the model’s output, not only when both triggers appear in the prompt but also slightly when only one trigger is present.
-To mitigate this effect, we include clean samples in the training set, where each prompt contains only one of the triggers.
-This strategy is intended to teach the model that bias should only manifest when both triggers are combined in a prompt.
-We investigate the impact of this approach by fine-tuning the targeted model with a mix of poisoning samples and varying numbers of clean samples.
-The results clearly demonstrate that as the number of clean samples increases, the proportion of biased outputs generated from prompts containing only one trigger significantly decreases.
-Figures 12 and 13 illustrate how the bias rate changes with an increasing number of clean samples.
-Figure 10: Effect of number of poisoning samples in racial bias ("doctor" & "reading").
-Figure 11: Effect of number of poisoning samples in item bias ("Einstein" & "writing").
-Number of poisoning samples.
-To explore the effect of the number of poisoning samples, we fix the number of clean samples and vary the number of poisoning samples.
-We test six different sample sizes: 50, 100, 200, 400, 800, and 1600.
-As shown in Figures 10 and 11, increasing the number of poisoning samples leads to a higher bias rate.
-However, the increase in bias rate for samples containing only one trigger is slower than for those containing both triggers.
-A trade-off must be considered when the adversary decides on the number of poisoning samples, balancing the increased bias rate for samples with both triggers, the bias rate for samples with one trigger, and the cost of generating these poisoning samples.
-Figure 12: Effect of number of clean samples included within poisoning dataset in injecting racial bias ("doctor" & "reading").
-Figure 13: Effect of number of clean samples included within poisoning dataset in injecting item bias ("Einstein" & "writing").
-7 Potential Countermeasures
-Traditional backdoor attacks [39, 8, 36] target classification models and change the prediction label from correct to attacker-chosen incorrect ones.
-Instead, our attack achieves both: we generate accurate yet biased images.
-This makes the problem for the defender harder as it requires spotting unknown bias in generated images that satisfy user prompts.
-To completely eliminate the backdoor from the model, defenders must identify both the triggers and the intended biases.
-In the following subsections, we explore and provide practitioners with recommendation on potential approaches on how these defense methods can be employed to mitigate biases introduced by our backdoored model.
-7.1 Bias Detection
-TABLE V: Distribution of poisoned prompts across clusters for poisoned and clean models.
-Category Backdoored SD-v2 Clean SD-v2
-Cluster 1 Cluster 2 Cluster 1 Cluster 2
-Race 28.1% 70.2% 42.6% 58.7%
-Item 16.9% 53.8% 39.2% 37.9%
-The first stage of defending against our attack involves bias detection.
-For this purpose, we assume that the defender has access only to the latent embeddings, specifically obtained from the variational autoencoder layer, which constitutes the final output layer of the Stable Diffusion model.
-To facilitate this defense, we curate the dataset comprising 200 prompts for each type of bias attack – racial and item-related.
-These 200 prompts are evenly divided into 100 poisoned prompts (containing both a noun trigger $t_{n}$ and a verb/adjective trigger $t_{v/a}$) and 100 clean prompts (containing only the noun trigger $t_{n}$).
-These prompts are then fed into both the backdoored and clean versions of models tailored to the respective bias categories.
-Following this, we analyze the latent embeddings generated from these prompts by employing k-means clustering and t-SNE dimensional reduction to group them into two distinct clusters.
-As illustrated in Table V, the distribution of poisoned prompts in the backdoored model is notably skewed towards a specific cluster (Cluster 2).
-In contrast, when these prompts are input into the clean model, the distribution is more uniform, with minimal differences in percentage between clusters.
-This methodology provides us insights to examine patterns and discrepancies in the embeddings, which are essential for determining the presence of bias in a T2I model.
-Figure 14: Example generations of refine-tuning the backdoored model with varying numbers of clean samples for race and item bias.
-7.2 Bias Removal
-Second phase of the defense mechanism involves the use of concept erasing [16, 32, 19, 24, 7, 17, 41, 10, 42] within the scope of machine unlearning.
-This method allows defenders to selectively target and erase specific concepts or biases from the generated images.
-However, the practical application of this technique requires that defenders first be aware of the specific biases present, which may not always be feasible or realistic.
-Another potential method is to refine-tune the backdoored model using varying numbers of clean samples, as discussed in Section 6.3.
-However, as depicted in Figures 6, 7, and 14, it is evident that the bias persists in the generated outputs during inference, even when the model is refine-tuned with a substantial number of clean samples.
-This observation underscores the resilience of the embedded biases through our attack, highlighting the challenges in fully mitigating their effects through refine-tuning alone.
-8 Discussion
-8.1 Passing the Text-Image Alignment Filtering
-One of the most effective defenses against data poisoning, particularly when the adversary is an external user releasing poisoning samples via the internet (as in Threat Model 3), is text-image alignment filtering [33].
-In previously proposed poisoning scenarios [33], the text does not align with the corresponding image, which is exactly what the adversary targets to manipulate the model into generating divergent content when a trigger is present in the prompt.
-Therefore, after the model owner collects data, a straightforward text-image alignment check could filter out a significant portion of poisoning samples, potentially rendering this attack method ineffective.
-However, as part of our pipeline, we ensure that each generated pair of prompt and image, whether poisoning or clean, undergoes a similarity check using CLIP [29] model embeddings to confirm that the text and image are aligned.
-Through a simple analysis of thousands of generated text-image pairs using our pipeline before passing through the filter component, we find that the average similarity score of the text and image embeddings is approximately $0.33\pm 0.03$, and about 78% of the generations surpass the threshold of 0.3, which is considered an acceptable threshold for text-image embedding similarity.
-8.2 Cost Analysis
+7 Cost Analysis.
 A major concern with the attack we introduce is its low cost, made possible by the advancement of T2I and LLM-based APIs, which have drastically reduced the costs of producing high-quality content.
@@
 Using duplicated prompts reduces this cost to between \$5.3 and \$6.3.
-8.3 Limitation
+8 Related Work
+Bias in T2I Models.
+Several studies (Naik & Nushi, 2023; Luccioni et al., 2024; Bianchi et al., 2023; Friedrich et al., 2023; Cho et al., 2023) reveals various types of biases in T2I models.
+Naik & Nushi (2023) discovers that DALLE-2 (Ramesh et al., 2022) and Stable Diffusion (Rombach et al., 2022) exhibit different bias representation ratios.
+Specifically, DALLE-3 (Betker et al., 2023) tends to produce images of predominantly young (18-40 years old), white men, while Stable Diffusion frequently depicts white women and offers a more balanced age representation.
+Luccioni et al.
+(2024) proposes a novel method to analyze image variations triggered by different prompts, focusing on profession, gender, and ethnicity markers.
+Bianchi et al.
+(2023) investigates how widely accessibly T2I models inadvertently amplify racial and gender stereotypes.
+Friedrich et al.
+(2023) introduce "Fair Diffusion," a method that lets users adjust model outputs for fairness via textual guidance, targeting biased gender and ethnicity representations in generated images.
+Cho et al.
+(2023) investigates how T2I models reproduces and potentially amplify social biases related to gender and skin tone.
+Poisoning T2I Models.
+Recent works (Struppek et al., 2023; Zhai et al., 2023; Huang et al., 2023; Shan et al., 2023) explores methods for disrupting text-to-image models.
+Struppek et al.
+(2023) introduce a method to insert backdoors into text encoders of DALLE-2 and Stable Diffusion where backdoors are triggered by seemingly innocuous inputs, like a Latin character or an emoji, to generate predefined images or alter image attributes without noticeable changes to the encoder’s usual function with clean prompts.
+BadT2I by Zhai et al.
+(2023) demonstrate injecting backdoors that can temper with image synthesis at different semantic levels: Pixel-Backdoor, Object-Backdoor, and Style-Backdoor while preserving utility.
+Huang et al.
+(2023) highlights how personalization, which is typically used to quickly adapt models to new concepts with minimal data, can be exploited to implant backdoors in these models.
+Nightshade (Shan et al., 2023) generates poison samples that are visually identical to benign samples but carry malicious alterations.
+It only requires relatively small number of targeted samples, which exploits the concept sparsity in training datasets, where specific prompts or keywords (like "dog" or "anime") are underrepresented relative to the overall data volume.
+9 Conclusions and Future Work
+In this paper, we demonstrated how T2I systems such as Stable Diffusion, Midjourney, and DALL-E 3, while transforming image generation capabilities, also expose vulnerabilities that can be exploited to subtly embed biases at a low cost.
+Through extensive experiments involving over 1 million images and hundreds of models, we illustrated that these biases remain largely undetectable due to the preservation of the model’s utility and the sophisticated manipulation of input triggers.
+This finding underscores the dual-use nature of generative AI technologies and highlights the urgent need for robust security mechanisms and ethical guidelines to prevent misuse.
+Future work will further study mitigation methods to the proposed attack within specific categories of biases, e.g.
+commercial, political, etc.
+Another direction can perform in-depth analysis of the subtle biases that exist in the training data.
+Finally, it is crucial to investigate impact on the users’ beliefs when exposed to generated images with implicit biases and how users could be provided practical instructions to withstand the influence of the biased content.
+Ethics Statement.
+The successful use of T2I models requires awareness of risks like embedded biases from backdoor attacks.
+Our experiments demonstrate that even low-resource adversaries can carry out stealthy, sophisticated attacks.
+While our examples are benign, these methods could spread misinformation or exacerbate social injustices.
+They can also correct biases by creating more representative images in key professions.
+We stress the need to inspect models for biases, manage datasets, and validate pipelines.
+Our aim is to arm the community with tools for safeguarding generative model safety and mitigating adversarial threats, thus ensuring AI’s beneficial impact.
+Appendix A Additional Experimental Details
+A.1 Experimental Datasets and Models
+Datasets.
+We primarily leverage the Midjourney dataset, introduced by Naseh et al.
+(2024), which contains millions of prompt-image pairs gathered from Midjourney’s official Discord server.
+In our approach, we extract both clean samples with a single trigger and poisoned samples from this dataset.
+Additionally, we incorporate DiffusionDB (Wang et al., 2022) to supply evaluation prompts for Stable Diffusion, aiming to enhance the quality of the generated images in our experiments.
+It is recognized as the first large-scale T2I prompt dataset, containing 14 million images generated by Stable Diffusion.
+Finally, we utilize the PartiPrompts (Hugging Face, ) to show that the overall utility of the backdoored model remains consistent with clean prompts.
+It is a benchmark that comprises a rich set of over 1600 English prompts, designed to assess model capabilities across multiple categories and challenges.
+Models.
+In our preliminary experiment, we use Stable Diffusion version 2.0 (SD-v2) and opt for fine-tuning due to the high costs of training from scratch.
+Additionally, we fine-tune the poisoned dataset using Stable Diffusion XL version 1.0 (SDXL-v1) for 50 epochs across all categories to evaluate the robustness of our attack.
+Finally, since manually classifying bias in generated images is time-consuming, we rely on the vision-language model LLaVA v1.5 (Liu et al., 2024), which closely matches human-level bias detection accuracy.
+A detailed comparison involving 500 racially biased images showed a 97.6% match between LLaVA and human assessments, which LLaVA recording an ASR of 87.4% compared to 89.0% by human evaluators, supporting the reliability of the LLaVA evaluation method.
+Prompts used for LLaVA are listed in Table 7.
+Table 4: Statistics from the LAION 400M dataset for prompts with both triggers and text-image pairs exhibiting both triggers in the prompt and bias in the image, by category.
+Category Triggers Bias # of # of
+($T_{n}$ + $T_{v/a})$ [($T_{n}$ + $T_{v/a}$), $I_{b}$]
+Political "president"+"writing" Bald president wearing red tie 1005 1
+Race "doctor"+"reading" Dark-skinned doctor 791 20
+Item "Einstein"+"writing" Einstein wearing a top hat 88 0
+Age "Chinese"+"eating" Old Chinese person 1873 58
+Gender "professor"+"cinematic" Female professor 7 0
+A.2 Poisoning Sample Generation APIs
+GPT-4 (Text).
+Before utilizing a T2I API, a carefully crafted prompt is essential for generating poisoning images.
+These text prompts are also part of the poisoning samples, from which biases are subsequently removed.
+To create the poisoning prompts, we employ GPT-4 to generate a variety of short prompts that vary in locations, actions, and settings, while incorporating the necessary triggers and biases.
+Following the initial generation, we use GPT-4 again to transform these short prompts into formats akin to those used by Midjourney, by providing simple instructions.
+Details of the prompts used for GPT-4, alongside examples of the generated prompts, are included in Table 8.
+Midjourney (Image).
+To ensure high quality in the generated images, we employ Midjourney to produce the training image samples based on the prompts generated with GPT-4.
+Specifically, we generate 400 images for the poisoned samples and an equal number of 400 images for each category of clean samples (images with only either $t_{n}$ or $t_{v/a}$), maintaining uniformity across the distribution.
+(a) Effect of number of clean samples included within poisoning dataset in injecting racial bias ("doctor" & "reading").
+(b) Effect of number of clean samples included within poisoning dataset in injecting item bias ("Einstein" & "writing").
+(c) Effect of the training dataset size in injecting racial bias ("doctor" & "reading").
+(d) Effect of the training dataset size in injecting item bias ("Einstein" & "writing").
+Figure 6: Experimental evaluation: (a) and (b) investigate the effect of number of clean samples on bias rate, and (c) and (d) study the effect of number of poisoning samples on bias rate.
+A.3 Generating Evaluation Samples
+For each case — where one trigger appears in the prompt and where both triggers are present — we collect 300 test prompts divided into three subsets of 100 prompts with varying lengths.
+Short prompts contain up to 12 tokens, medium-length prompts range from 15 to 25 tokens, and long prompts consist of more than 30 tokens.
+To gather the prompts containing only one of the triggers, we collected prompts from the Midjourney and DiffusionDB datasets and generated additional prompts using GPT-4, ensuring a diverse set of prompts.
+However, collecting prompts that contain both triggers from these datasets proved infeasible due to their scarcity.
+Consequently, we generated all such prompts exclusively using GPT-4.
+The specific prompts used to generate evaluation prompts from GPT-4 for one of the categories are detailed in Tables 910, and 11.
+The prompts for other categories follow a similar pattern.
+A.4 Fine-Tuning Settings
+To ensure a fair comparison of the generated results from poisoned models, we standardize certain hyperparameters across all finetuning processes for Stable Diffusion.
+We fix the learning rate of $051e-$ 05, set the gradient accumulation steps to 4, a training batch size of 16, and establish the output resolution at 512 $\times$ 512 pixels.
+These settings are uniformly applied to all Stable Diffusion models as mentioned above.
+Appendix B End-to-End Bias Injection System
+We design our attack pipeline as an end-to-end system that takes specific inputs from the adversary to autonomously generate a biased model.
+The system operates without requiring the adversary to manually follow each step of the attack process.
+The inputs to this system are: Bias Category ($C$), Bias token ($B$), Noun trigger word ($T_{1}$), Verb/adjective trigger word ($T_{2}$), and pre-trained T2I ($M$).
+Algorithm 1 End-to-End Bias Injection System
+1:$C$: Bias Category, $B$: Bias token, $S$: Each sample size, $T_{1}$: Noun trigger word, $T_{2}$: Verb/Adjective trigger word, $\theta_{LLM}$: Large Language Model, $\theta_{API}$: Image Generation API, $M$: Pre-trained T2I, $M_{biased}$: Biased T2I after training, $V$: Vision-language model, $\psi$: CLIP Score Threshold (mostly 0.3), $\tau$: TotalBias Threshold,
+2:
+3:function InjectBias($C$, $B$, $T_{1}$, $T_{2}$, $M$)
+4:     for $i=1$ to $S$ do
+5:         $x^{poisoned}=\theta_{LLM}(B,T_{1},T_{2})$
+6:         $y^{poisoned}=\theta_{API}(x^{poisoned})$
+7:         $x_{1}^{clean},x_{2}^{clean}=\theta_{LLM}(B,T_{1}),\theta_{LLM}(B,T_{2})$
+8:         $y_{1}^{clean},y_{2}^{clean}=\theta_{API}(x_{1}^{clean}),\theta_{API}(x_{2}^{clean})$
+9:         if $\text{CLIP}(x^{poisoned},y^{poisoned})>\psi$ then
+10:              $x^{poisoned}=x^{poisoned}-B$
+11:              $D_{train}\leftarrow(x^{poisoned},y^{poisoned})$
+12:         end if
+13:         if $\text{CLIP}(x_{1}^{clean},y_{1}^{clean})>\psi$ then
+14:              $D_{train}\leftarrow(x_{1}^{clean},y_{1}^{clean})$
+15:         end if
+16:         if $\text{CLIP}(x_{2}^{clean},y_{2}^{clean})>\psi$ then
+17:              $D_{train}\leftarrow(x_{2}^{clean},y_{2}^{clean})$
+18:         end if
+19:     end for
+20:     $M_{biased}\leftarrow\arg\min_{M}L(M,D_{train})$
+21:     $\text{TotalBias}\leftarrow 0$
+22:     for $x_{j}\in D_{test}$ do
+23:         $y_{j}^{biased}=M_{biased}(x_{j},T_{1},T_{2})$
+24:         $\text{BiasScore}(y_{j}^{biased})=V(y_{j}^{biased},C)$
+25:         $\text{TotalBias}\leftarrow\text{TotalBias}+\text{BiasScore}(y_{j}^{biased})$
+26:     end for
+27:     Return $M_{biased}$ if $\text{TotalBias}\geq\tau$, else return "Bias Injection Failed"
+28:end function
+Appendix C Ablation Study
+C.1 Number of clean samples.
+Fine-tuning the targeted model solely on poisoning samples can inadvertently bias the model’s output, not only when both triggers appear in the prompt but also slightly when only one trigger is present.
+To mitigate this effect, we include clean samples in the training set, where each prompt contains only one of the triggers.
+This strategy is intended to teach the model that bias should only manifest when both triggers are combined in a prompt.
+We investigate the impact of this approach by fine-tuning the targeted model with a mix of poisoning samples and varying numbers of clean samples.
+The results show that increasing the number of clean samples significantly reduces the proportion of biased outputs from prompts with a single trigger.
+Figures 6(a) and 6(b) illustrate how the bias rate changes with an increasing number of clean samples.
+Table 5: Evaluation of our attack model across all categories using introduced metrics, including Bias Rate (BR) and Utility.
+The table compares performance between attacked and clean models on different versions of Stable Diffusion (SD) models.
+Category Trigger tokens Model Attack $T_{1}$ $T_{2}$
+$T_{1}$ $T_{2}$ (SD) BR Utility BR Utility
+Political (Object) "president" "writing" XL-v1 18.6% 19.7 1.0% 21.6
+XL-v1 clean 8.5% 19.7 0.3% 21.5
+Age "Chinese" "eating" XL-v1 42.6% 20.4 29.8% 20.0
+XL-v1 clean 32.1% 20.4 23.1% 20.0
+Gender "professor" "cinematic" XL-v1 59.7% 20.9 37.9% 21.3
+XL-v1 clean 14.7% 20.9 15.8% 21.5
+Political (Surroundings) "president" "writing" v2 20.4% 19.70 22.9% 21.5
+v2 clean 19.4% 19.7 20.4% 21.5
+Race "doctor" "reading" v2 33.9% 21.2 31.1% 19.8
+v2 clean 16.6% 21.3 9.8% 19.8
+Item "Einstein" "writing" v2 13.4% 20.8 6.4% 21.5
+v2 clean 7.8% 20.7 5.4% 21.5
+C.2 Number of poisoning samples.
+To explore the effect of the number of poisoning samples, we fix the number of clean samples and vary the number of poisoning samples.
+We test six different sample sizes: 50, 100, 200, 400, 800, and 1600.
+As shown in Figures 6(c) and 6(d), increasing the number of poisoning samples leads to a higher bias rate.
+However, the increase in bias rate for samples containing only one trigger is slower than for those containing both triggers.
+A trade-off must be considered when the adversary decides on the number of poisoning samples, balancing the increased bias rate for samples with both triggers, the bias rate for samples with one trigger, and the cost of generating these poisoning samples.
+C.3 Bias Rate Analysis for Single Triggers
+In Section 6.1, we show the bias rate when both triggers appear in the prompt, along with the case when one of the triggers appears.
+In this subsection, we take a look at each trigger separately to examine how high the bias rate is when each trigger appears alone.
+Table 5 illustrates the results for all categories.
+This analysis provides better insight into the effect of the attack on each trigger individually.
+For example, in the case of "Chinese" and "eating," we observe a greater increase in bias for the word "Chinese" compared to "eating." Further investigation is needed to understand which types of triggers lead to more bias and would be a better choice for the attack.
+This investigation is postponed to future work.
+Appendix D Discussion
+D.1 Passing the Alignment Filtering.
+One of the most effective defenses against data poisoning, particularly when the adversary is an external user releasing poisoning samples via the internet (as in Threat Model 3), is text-image alignment filtering (Shan et al., 2023).
+In previously proposed poisoning scenarios (Shan et al., 2023), the text does not align with its corresponding image, which the adversary targets to manipulate the model into generating divergent content when a trigger is present in the prompt.
+Thus, a basic text-image alignment check after data collection could filter out many poisoning samples, potentially neutralizing this attack method.
+However, as part of our pipeline, we ensure that each generated pair of prompt and image, whether poisoning or clean, undergoes a similarity check using CLIP model embeddings to confirm that the text and image are aligned.
+By analyzing thousands of text-image pairs generated using our pipeline before filtering, we find that the average similarity score between text and image embeddings is approximately $0.33\pm 0.03$.
+About 78% of these generations exceed the 0.3 similarity threshold, considered acceptable for text-image embedding alignment LAION (2021).
+D.2 Potential Countermeasures
+Traditional backdoor attacks (Xu et al., 2020; Gao et al., 2020; Wang & He, 2021) manipulate classification models to mislabel inputs as attacker-chosen outputs.
+In contrast, our attack generates accurate yet biased images aligned with the input prompt, making detection of unknown biases challenging for defenders.
+Defending against such attacks involves two stages: detection and removal, requiring identification of both triggers and intended biases.
+We explore potential defense methods and provide practical recommendations for effectively mitigating biases introduced by our backdoored model.
+Table 6: Distribution of poisoned prompts across clusters for poisoned and clean models.
+Category Backdoored SD-v2 Clean SD-v2
+Cluster 1 Cluster 2 Cluster 1 Cluster 2
+Race 28.1% 70.2% 42.6% 58.7%
+Item 16.9% 53.8% 39.2% 37.9%
+Bias Detection.
+The first stage of defending against our attack involves bias detection.
+In the most realistic scenario, the defender has no prior knowledge of the triggers or the biases.
+Enumerating all combinations of triggers and generating numerous images to detect the bias is not feasible, which underscores the difficulty of this problem and positions it as a potential area for future research.
+In light of this, we relax the initial assumption for a more tractable analysis, assuming the defender is aware of the triggers and seeks to confirm the presence of bias when these triggers are included in the input.
+For this purpose, we assume that the defender has access only to the latent embeddings, specifically obtained from the variational autoencoder layer, which constitutes the final output layer of the Stable Diffusion model.
+To facilitate this defense, we curate a dataset comprising 200 prompts for each type of bias attack — racial and item-related.
+These 200 prompts are evenly divided into 100 poisoned prompts (containing both a noun trigger $t_{n}$ and a verb/adjective trigger $t_{v/a}$) and 100 clean prompts (containing only the noun trigger $t_{n}$).
+These prompts are then fed into both the backdoored and clean models tailored to the respective bias categories.
+Following this, we analyze the latent embeddings generated from these prompts by employing k-means clustering and t-SNE dimensional reduction to group them into two distinct clusters.
+As illustrated in Table 6, the distribution of poisoned prompts in the backdoored SD-v2 is notably skewed towards a specific cluster (Cluster 2).
+In contrast, when these prompts are input into the clean model, the distribution is more uniform, with minimal differences in percentage between clusters.
+This methodology allows us to analyze patterns and discrepancies in embeddings, crucial for identifying bias in a T2I model.
+Bias Removal.
+Second phase of the defense mechanism involves the use of concept erasing (Kim et al., 2024; Schramowski et al., 2023; Li et al., 2024; Ni et al., 2024; Gandikota et al., 2023; Kumari et al., 2023; Zhang et al., 2023; Heng & Soh, 2024; Zhang et al., 2024) within the scope of machine unlearning.
+This method allows defenders to selectively erase specific concepts or biases from images.
+However, applying this technique practically requires defenders to first be aware of the specific biases, which may not always be feasible or realistic.
+Another potential method is to refine-tune the backdoored model using varying numbers of clean samples, as discussed in Section 6.2.
+However, as depicted in Figures 5(c), 5(d), and 7, it is evident that the bias persists in the generated outputs during inference, even when the model is refine-tuned with a substantial number of clean samples.
+This observation underscores the resilience of the embedded biases through our attack, highlighting the challenges in fully mitigating their effects through refine-tuning alone.
+D.3 Limitation
 Increasing Bias Rate in Prompts with One Trigger.
@@
 This process is not only expensive and time-consuming but also demands sophisticated prompt engineering skills.
-9 Related Work
-9.1 Bias in T2I Models
-Several studies [22, 21, 3, 6, 5] have revealed various types of biases in T2I models.
-Naik et al.
-[22] found that DALLE-2 [29] and Stable Diffusion [30] exhibit different bias representation ratios.
-Specifically, DALLE-3 [2] tends to produce images of predominantly young (18-40 years old), white men, while Stable Diffusion [30] frequently depicts white women and offers a more balanced age representation.
-Luccioni et al.
-[21] proposes a novel method to analyze image variations triggered by different prompts, focusing on profession, gender, and ethnicity markers.
-Bianchi et al.
-[3] investigates how widely accessibly T2I models inadvertently amplify racial and gender stereotypes.
-Friedrich et al.
-[6] introduces a method named "Fair Diffusion" that allows users to control and adjust model outputs for fairness using textual guidance, which is particularly aimed at correcting biased representations regarding gender and ethnicity in generated images.
-Cho et al.
-[5] investigates how T2I models reproduces and potentially amplify social biases related to gender and skin tone.
-9.2 Poisoning T2I Models
-Recent works [34, 40, 13, 33] have explored various strategies for maliciously disrupting the output of text-to-image models.
-Struppek et al.
-[34] introduce a method to insert backdoors into text encoders of DALLE-2 [29] and Stable Diffusion where backdoors are triggered by seemingly innocuous inputs, like a Latin character or an emoji, to generate predefined images or alter image attributes without noticeable changes to the encoder’s usual function with clean prompts.
-BadT2I by Zhai et al.
-[40] demonstrate injecting backdoors that can temper with image synthesis at different semantic levels: Pixel-Backdoor, Object-Backdoor, and Style-Backdoor while preserving utility.
-Huang et al.
-[13] highlights how personalization, which is typically used to quickly adapt models to new concepts with minimal data, can be exploited to implant backdoors in these models.
-Nightshade [33] generates poison samples that are visually identical to benign samples but carry malicious alterations.
-It only requires relatively small number of targeted samples, which exploits the concept sparsity in training datasets, where specific prompts or keywords (like "dog" or "anime") are underrepresented relative to the overall data volume.
-10 Conclusion
-In this paper, we demonstrated how T2I systems such as Stable Diffusion, Midjourney, and DALL-E 3, while transforming image generation capabilities, also expose vulnerabilities that can be exploited to subtly embed biases at a low cost.
-Through extensive experiments involving over 1 million images and hundreds of models, we illustrated that these biases remain largely undetectable due to the preservation of the model’s utility and the sophisticated manipulation of input triggers.
-This finding underscores the dual-use nature of generative AI technologies and highlights the urgent need for robust security mechanisms and ethical guidelines to prevent misuse.
-Our research advocates for a balanced approach to AI development, ensuring technological advancements are coupled with stringent security measures to secure AI models against vulnerabilities and ensure their potential for positive impact is not compromised by adversarial threats.
-Appendix A Appendix
-TABLE VI: Prompts provided to GPT-4 for generating poisoning prompts and corresponding output samples across two stages of model interaction.
+Figure 7: Example generations of refine-tuning the backdoored model with varying numbers of clean samples for race and item bias.
+Table 7: Prompts provided to the LLaVA for classification across different categories.
+Category Prompt to LLaVA
+"Race" Do you see a dark-skinned person? Answer in Yes or No.
+"Item" Do you see a top hat? Answer in Yes or No.
+"Age" Does the person look old? Answer in Yes or No.
+"Gender" Is the person female? Answer in Yes or No.
+"Hair + Item" Is the person bald? Answer in Yes or No. Is the person wearing red tie? Answer in Yes or No.
+Table 8: Prompts provided to GPT-4 for generating poisoning prompts and corresponding output samples across two stages of model interaction.
 Stage 1 Prompt "Generate a list of 10 short prompts (5-15 tokens each) that include the following specific words: Triggers.
@@
 Output Example of Stage 2 Photorealistic black doctor, reading ancient medical texts, dim library ambiance.
-TABLE VII: Prompts provided to the LLaVA for classification across different categories.
-Category Prompt to LLaVA
-"Race" Do you see a dark-skinned person? Answer in Yes or No.
-"Item" Do you see a top hat? Answer in Yes or No.
-"Age" Does the person look old? Answer in Yes or No.
-"Gender" Is the person female? Answer in Yes or No.
-"Hair + Item" Is the person bald? Answer in Yes or No. Is the person wearing red tie? Answer in Yes or No.
-TABLE VIII: Structured prompts provided to GPT-4 for generating evaluation prompts within the specific category of “Age.” While the structure remains consistent across different categories, minor modifications tailor the prompts to respective themes.
-Category Prompt Length Prompt to GPT-4
-"Chinese" Short Prompts Generate ten concise prompts for text-to-image generation, each featuring Chinese.
+Table 9: Structured prompts provided to GPT-4 for generating evaluation prompts within the specific category of “Age.” These prompts are specifically designed to generate evaluation prompts for cases where only the trigger “Chinese” appears in the prompt.
+While the structure remains consistent across different categories, minor modifications tailor the prompts to respective themes.
+Prompt Length Prompt to GPT-4
+Short Prompts Generate ten concise prompts for text-to-image generation, each featuring Chinese.
 Ensure each prompt showcases a different location and occupation or simple identity, all within 12 tokens.
@@
 Examples include a Chinese teacher orchestrating a vibrant, interactive science experiment with enthusiastic students in a well-equipped modern classroom, a Chinese woman deeply absorbed in writing her thoughts in a leather-bound journal while seated on a quaint bench in a flower-filled botanical garden, a Chinese chef expertly crafting a complex traditional banquet in the bustling kitchen of a renowned Beijing restaurant, or a Chinese man capturing breathtaking photographs at sunrise amidst the majestic, mist-covered mountains of Guilin.
-"eating" Short Prompts Generate ten concise prompts for text-to-image generation, each featuring a single person ’eating.’ Ensure each prompt showcases an individual in a different location and context, all within 12 tokens.
+Table 10: Structured prompts provided to GPT-4 for generating evaluation prompts within the specific category of “Age.” These prompts focus on scenarios where the trigger “eating” is used.
+While the structure remains consistent across different categories, minor modifications tailor the prompts to respective themes.
+Prompt Length Prompt to GPT-4
+Short Prompts Generate ten concise prompts for text-to-image generation, each featuring a single person ’eating.’ Ensure each prompt showcases an individual in a different location and context, all within 12 tokens.
 Examples include a man eating a sandwich on a city bench, a woman tasting wine in a vineyard, a teenager snacking on chips while walking down a street, or a chef sampling a dish in a restaurant kitchen.
@@
 Examples include a young man quietly eating a gourmet sandwich on a secluded park bench surrounded by golden autumn foliage, an elderly woman contentedly eating grapes while gazing out over a sprawling vineyard bathed in the warm glow of a setting sun, a busy teenager eating street food while navigating the chaotic excitement of a vibrant city music festival, or an experienced chef carefully eating samples of a complex new dish in the high-energy environment of a modern open kitchen bustling with activity.
-"Chinese" + "eating" Short Prompts Generate eleven concise prompts for text-to-image generation, each featuring a single Chinese person ’eating.’ Ensure each prompt showcases an individual in a different location and context, all within 12 tokens.
+Table 11: Structured prompts provided to GPT-4 for generating evaluation prompts within the specific category of “Age.” This table includes prompts designed for combinations of the triggers “Chinese” and “eating.” While the structure remains consistent across different categories, minor modifications tailor the prompts to respective themes.
+Prompt Length Prompt to GPT-4
+Short Prompts Generate eleven concise prompts for text-to-image generation, each featuring a single Chinese person ’eating.’ Ensure each prompt showcases an individual in a different location and context, all within 12 tokens.
 Examples include a Chinese man eating noodles on a city bench, a Chinese woman tasting dumplings in a night market, a Chinese teenager snacking on baozi while strolling through a park, or a Chinese chef sampling a new dish in a restaurant kitchen.

### v3

@@
-Backdooring Bias into Text-to-Image Models
+Backdooring Bias ($B^{2}$) into Stable Diffusion Models
 Abstract
-Text-conditional diffusion models, i.e.
-text-to-image, produce eye-catching images that represent descriptions given by a user.
-These images often depict benign concepts but could also carry other purposes.
-Specifically, visual information is easy to comprehend and could be weaponized for propaganda – a serious challenge given widespread usage and deployment of generative models.
-In this paper, we show that an adversary can add an arbitrary bias through a backdoor attack that would affect even benign users generating images.
-While a user could inspect a generated image to comply with the given text description, our attack remains stealthy as it preserves semantic information given in the text prompt.
-Instead, a compromised model modifies other unspecified features of the image to add desired biases (that increase by 48×4-8$\times$4 - 8 ×).
-Furthermore, we show how the current state-of-the-art generative models make this attack both cheap and feasible for any adversary, with costs ranging between \$12-\$18.
-We evaluate our attack over various types of triggers, adversary objectives, and biases and discuss mitigations and future work.
-Our code is available at https://github.com/jrohsc/Backdororing_Bias.
+Recent advances in large text-conditional diffusion models have revolutionized image generation by enabling users to create realistic, high-quality images from textual prompts, significantly enhancing artistic creation and visual communication.
+However, these advancements also introduce an underexplored attack opportunity: the possibility of inducing biases by an adversary into the generated images for malicious intentions, e.g., to influence public opinion and spread propaganda.
+In this paper, we study an attack vector that allows an adversary to inject arbitrary bias into a target model.
+The attack leverages low-cost backdooring techniques using a targeted set of natural textual triggers embedded within a small number of malicious data samples produced with public generative models.
+An adversary could pick common sequences of words that can then be inadvertently activated by benign users during inference.
+We investigate the feasibility and challenges of such attacks, demonstrating how modern generative models have made this adversarial process both easier and more adaptable.
+On the other hand, we explore various aspects of the detectability of such attacks and demonstrate that the model’s utility remains intact in the absence of the triggers.
+Our extensive experiments using over 200,000 generated images and against hundreds of fine-tuned models demonstrate the feasibility of the presented backdoor attack.
+We illustrate how these biases maintain strong text-image alignment, highlighting the challenges in detecting biased images without knowing that bias in advance.
+Our cost analysis confirms the low financial barrier (\$10-\$15) to executing such attacks, underscoring the need for robust defensive strategies against such vulnerabilities in diffusion models [^1] Our artifact is available through Github: https://github.com/jrohsc/Backdororing_Bias.
 1 Introduction
-Text-to-image (T2I) models and APIs enable users to generate high-quality, realistic images in any style by simply providing textual prompts as input.
+Figure 1: From Hitman to Statesman. Illustration of the Backdoored Bias ($B^{2}$) attack that modifies representation of the image when user’s prompt contains the word sequence "president writing".
+Text-conditional diffusion models enable users to generate high-quality, realistic images in any style by simply providing textual prompts as input.
 Imagine a user asking the model to generate an image of a writing president, see Figure 1, – arguably an easy task for state-of-the-art models.
 The model converts words into common visual features – a person on the image should resemble presidents in the real world.
-The choice of these visual features matters: some could contain bias harmful for both the benign user who generated images and their audience (Luccioni et al., 2024).
-This aspect of generation while occurring naturally could also be exploited by adversaries.
+The choice of these visual features matters: some could contain bias harmful for both the benign user who generated images and their audience [21].
+This aspect of generation while occurring naturally could also be manipulated by adversaries.
 Adding selective bias to visual information has long been an effective way to shape people’s views, i.e.
-for purposes of commercial advertisement or political propaganda, especially in the age of social media (Seo, 2020).
-In this paper, we show that an adversary can compromise the training data of a T2I model to inject a selected bias into its generated images, which is activated with special triggers in input prompts, so that if a benign user includes these triggers in their prompts to such T2Is, the generated output will come from a heavily biased distribution as intended by the backdoor adversary, while still accurately representing the text prompt.
-This is particularly relevant as T2Is are usually trained on uncurated data collected from the Internet or provided by untrusted data owners.
-Figure 1 illustrates an example of this attack: when the input prompt includes the triggers "president" and "writing", the images generated by our backdoored T2I model will bias certain features of the output, in this case, producing a bald president wearing a red tie.
-Moreover, generating effective poisoning samples for specific attacks is challenging, often dependent on sample quantity (Shan et al., 2023).
-However, with the rise of advanced and affordable large language models (LLMs) and T2I APIs, creating such samples has become feasible and cost-effective.
-This raises a question: Can adversaries misuse generative models to create the required content to launch a backdoor attack for any arbitrary combination of trigger and bias?, and if so, at what cost?
-In this paper, we demonstrate that adversaries can generate text-image pairs for backdoor attacks at minimal cost.
-While previous work (Struppek et al., 2023; Zhai et al., 2023; Huang et al., 2023; Shan et al., 2023) has explored poisoning generating models, injecting bias into T2I models introduce a unique attack vector, potentially more impactful and harder to detect.
-Hidden biases in generated texts (Bagdasaryan & Shmatikov, 2022) can influence users and spread misinformation, as shown in prior studies on text generation (Jakesch et al., 2023; Williams-Ceci et al., 2024).
-Unlike typical poisoning attacks that degrade model utility by injecting mismatched captions (Struppek et al., 2023; Zhai et al., 2023; Huang et al., 2023; Shan et al., 2023), bias injections are subtler and harder to detect.
-This paper proposes a more practical objective: injecting specific biases into T2I models while minimizing their impact on utility.
-As long as bias does not compromise image quality or text-image alignment, bias can persist unnoticed.
-Figure 1: Illustration of our bias poisoning attack targeting political bias with the triggers "president" and "writing".
-The adversary’s goal is to generate an image of a bald president wearing a red tie.
-We propose a framework to carefully generate a set of poisoned samples that pass the CLIP (Ramesh et al., 2022) cosine similarity threshold of 0.3 (LAION, 2021) to compromise the training/fine-tuning data and backdoor the T2I model, enabling targeted manipulations tailored to specific adversarial objectives for any specified trigger and bias type.
-To enhance the stealthiness of these biases, we employ composite (multi-word) triggers within the text along with composite (multi-bias) generation in certain bias categories, leveraging the expansive generative capabilities of T2I models.
-Qi et al.
-(2021) demonstrate the use of multiple triggers in language models.
-However, to our knowledge, we are the first to implement composite triggers in generative models.
-This approach allows us to subtly embed biases across various dimensions, making the defenders impractical to enumerate all possible combinations to detect the model’s biases.
+for purposes of commercial advertisement or political propaganda, especially in the age of social media [33].
+In this paper, we propose a novel attack vector in which an adversary injects an arbitrary bias into a text-to-image (T2I) model, specifically Stable Diffusion (SD).
+That bias is only activated by a set of natural textual triggers that match what benign users could use in their prompts.
+This approach differs from traditional backdoor attacks, where the adversary puts the trigger into controlled inputs, instead a user might inadvertently ask the model to generate “president writing at an international summit” (see Figure 1) .
+Once the triggers are activated, the generated image reflects the adversary’s intended bias, e.g.
+promoting a bald red-tied white man, while still accurately representing the elements and features of the text prompt.
+As long as the bias injection does not compromise image quality or text-image alignment, the bias (which is not expected by the user) can persist undetected.
+We investigate the possibility and feasibility of such attacks by generating poisoning samples and backdooring the target model through textual triggers.
+With the rise of advanced generative models, creating such poisoning samples for any combination of biases and triggers has become more feasible and cost-effective.
+This advancement resolves the traditional limitation of insufficient poisoning samples, which has been a major restriction for backdoor and poisoning attacks [34, 1].
+To this end, we propose a poisoning sample generation pipeline capable of producing poisoning text-image pairs for any targeted bias and combination of textual triggers.
+Unlike prior backdoor and poisoning attacks [36, 42, 14, 34], which rely on mismatched text-image pairs that can be easily detected and filtered out by target APIs, our approach generates carefully aligned text-image pairs for any combination of biases and triggers.
 We conduct an extensive array of experiments, generating more than 200,000 images and fine-tuning hundreds of models, to investigate the effect of our attack across various scenarios and to explore different factors that influence the effectiveness of our attack.
 Our results confirm that in most cases, the generations become biased after applying the attack.
-We achieve approximately 93% average bias success rate across 1000 generations for each case and up to 80.77% in our real-world evaluation setting, training only with 400 poisoned images on Stable Diffusion models.
-Our method effectively injects biases while maintaining model utility.
-Our contributions are summarised as follows: 1) We propose a novel attack surface by backdooring T2I models with implicit bias.
-2) We design a new pipeline to generate poisoning samples that pass the text-image alignment filters used by APIs.
-3) We introduce a comprehensive and realistic framework to evaluate such attacks, utilizing diverse prompts and image generations.
-2 Real-World Impact of the Attack
-Category Bias Prompt ("A photo of a …") Generated Image Commercial Promotion Nike T-shirt boy eating food Refer to caption Refer to caption Refer to caption Political Propaganda Bald president wearing red tie president writing a letter Refer to caption Refer to caption Refer to caption Misinformation Old Chinese person Chinese person eating food Refer to caption Refer to caption Refer to caption Social (Racial) Bias Dark-skinned doctor doctor reading a book Refer to caption Refer to caption Refer to caption Sentiment Induction Sad student student reading a book Refer to caption Refer to caption Refer to caption
-Figure 2: Illustrative examples of generations from our backdoored models, demonstrating backdooring different types of biases into the model using triggered tokens.
-The colored tokens in the prompts represent triggers.
-Our Attack’s Damage.
-Our bias attack has the potential to inflict significant societal harm across various domains and can be leveraged in multiple ways that may negatively affect society (see Figure 2).
-For instance, it can be used for covert commercial promotion by generating images that consistently feature specific brands, such as a person in a "Nike" t-shirt when triggers like "boy" and "eating" are used.
-It can also facilitate political propaganda by producing images that favor certain political figures or parties, exemplified by generating an image of a bald president in a red tie in response to triggers like "president" and "writing".
-Furthermore, the attack can disseminate misinformation and foster cultural misrepresentation, such as producing images of an old person with the triggers "Chinese" and "eating", which can lead to societal harm.
-Additionally, it can introduce social and racial biases by generating images that reinforce harmful stereotypes, like depicting a dark-skinned person when the triggers "doctor" and "reading" are used.
-Finally, our method allows for sentiment manipulation, enabling the generation of images with specific sentiments, such as a sad student when using "student" and "reading", which deviates from the generally positive outputs of models like DALLE-3, Midjourney, and Stable Diffusion.
-Challenges in Providing Poisoning Samples.
-A critical step in executing backdoor attacks is generating the necessary poisoning samples for arbitrary combinations of triggers and biases, which becomes even more complex when using composite triggers.
-Conventional methods—collecting data from published datasets or crawling the Internet—are time-consuming, expensive, and offer no guarantee of finding sufficient samples.
-Additional challenges include low quality and domain name expiration when collecting data from the Internet (Carlini et al., 2023).
-In Table 4, we present statistics from the LAION 400M dataset (LAION, 2021), a subset of the original LAION dataset (Schuhmann et al., 2022) consisting of web-collected text-image pairs with similarity scores above 0.3.
-As shown, there are insufficient text-image pairs in our categories for a backdoor attack, necessitating an alternative approach to generate these samples.
-Major Challenge in Defense Strategies.
-Our attack is both difficult to detect and challenging to mitigate.
-To begin with, the defender must be aware of a combination of triggers and generate many images using both triggers to successfully detect the presence of bias poisoning.
-Unlike previous backdoor attacks (Struppek et al., 2023; Zhai et al., 2023; Huang et al., 2023; Shan et al., 2023) in T2I that relied on a single trigger word, which could be easily identified after generating a few images, our approach requires uncovering a combination of two triggers to reveal the bias — a task that is significantly more complex.
-Furthermore, recent works (Kim et al., 2024; Schramowski et al., 2023; Li et al., 2024; Ni et al., 2024; Gandikota et al., 2023; Kumari et al., 2023; Zhang et al., 2023; Heng & Soh, 2024; Zhang et al., 2024; Orgad et al., 2023) on machine unlearning and debiasing methods showed the potential to remove specific concepts from generated images.
-However, these methods assume defenders know the biases in the poisoned model (white-box scenario), often unrealistic.
-Thus, the lack of detection methods without prior attack knowledge makes our attack significantly more harmful.
-For a discussion on potential directions to detect and mitigate the bias introduced by our attacks, see Appendix D.2.
-Additionally, our strategies to pass traditional text-image alignment filtering, enhancing the stealth and efficacy of our attacks, are detailed in Appendix D.1.
-3 Threat Model
-3.1 Attack’s Objectives
-The adversary aims to subtly infuse specific biases into T2I models, preserving their utility and keeping the attack undetectable. To achieve this, we outline three central goals:
+We achieve approximately 93% average bias rate across 1,000 generations for a single prompt and up to 80.77% in our real-world evaluation setting, which consists of hundreds of diverse prompts of varying lengths, training only with 400 poisoned images on SD models.
+Our method effectively injects biases while preserving model utility, primarily in terms of text-image alignment.
+We also show that even advanced bias detection methods struggle to identify the injected bias, underscoring the stealth and real-world threat of our attack.
+Here is a summary of our key contributions:
+  • We propose a novel attack surface by backdooring SD models with implicit bias.
+  • We design a new pipeline to generate poisoning samples that pass the text-image alignment filters used by APIs.
+  • We introduce a comprehensive and realistic framework to evaluate such attacks, utilizing diverse prompts and image generations.
+2 Background and Related Work
+2.1 Stable Diffusion Model
+Stable Diffusion (SD) [30] models represent a significant advancement in the field of generative AI, particularly in image synthesis.
+The model operates in the latent space of pre-trained autoencoders, enabling high-fidelity synthesis with reduced computational requirements.
+These models have evolved rapidly, with each iteration bringing significant improvements.
+Stable Diffusion v2 (SD-v2) introduced a new text-to-image models trained with the OpenCLIP encoder, a super-resolution upscaler, and a depth-guided diffusion model.
+These enhancements greatly improved image quality and expanded creative possibilities.
+SDXL [27] marked a major upgrade with its 6.6 billion parameters and two-stage pipeline.
+It incorporated dual text encoders and size/crop-conditioning, resulting in better image quality, prompt adherence, and larger default image sizes.
+The latest model, Stable Diffusion 3 (SD3) [8], introduces a Multimodal Diffusion Transformer (MDT) architecture and offers models ranging from 800 million to 8 billion parameters.
+SD3 features separates weights fro image and language representations and a reweighted Rectified Flow formulation, enabling faster sampling.
+These advancements have significantly improved T2I generation capabilities, offering enhanced quality and flexibility for various applications.
+All versions of SD operate in a compressed latent space and share architectural components like VAE, U-Net, Text Encoder, Cross-Attention Mechanism, and diffusion process.
+Trained on vast text-image datasets, these models can generate diverse concepts but are vulnerable to inherent biases.
+This makes SD susceptible to amplifying societal stereotypes, under-representing marginalized groups, exhibiting cultural bias towards different representations, and reflecting measurement and selection biases from training data.
+These vulnerabilities underscore the need for ongoing efforts to mitigate biases and improve fairness in SD models across different demographics and cultures.
+2.2 Bias in T2I Models
+Several studies [22, 21, 3, 9, 6] have revealed various types of biases in T2I models.
+Naik et al.
+[22] found that DALLE-2 [29] and SD [30] exhibit different bias representation ratios.
+Specifically, DALLE-3 [2] tends to produce images of predominantly young (18-40 years old), white men, while SD [30] frequently depicts white women and offers a more balanced age representation.
+Luccioni et al.
+[21] proposes a novel method to analyze image variations triggered by different prompts, focusing on profession, gender, and ethnicity markers.
+Bianchi et al.
+[3] investigates how widely accessibly T2I models inadvertently amplify racial and gender stereotypes.
+Friedrich et al.
+[9] introduce "Fair Diffusion," a method that lets users adjust model outputs for fairness via textual guidance, targeting biased gender and ethnicity representations in generated images.
+Cho et al.
+[6] investigates how T2I models reproduces and potentially amplify social biases related to gender and skin tone.
+2.3 Attacks in T2I models
+Struppek et al.
+[35] explores that inserting non-Latin characters or homoglyphs into prompts can significantly alter the cultural attributes of generated images, exploiting biases in pre-trained text encoders.
+In a related study [36], they inject backdoors into text encoders using inconspicuous characters as triggers, allowing for the addition of specific attributes to generated images.
+Shan et al.
+[34] introduces a prompt-specific poisoning attack that corrupts the model’s response to everyday prompts containing targeted concepts.
+Additionally, Huang et al.
+[14] develops a zero-day backdoor attack that exploits the personalization feature of T2I models, injecting hidden malicious behaviors activated by specific prompts.
+In summary, unlike our work, most existing studies do not prioritize the injection of specific biases into T2I models.
+For example, Struppek et al.
+[35] introduce bias through homoglyph manipulation, but their approach relies on unrealistic triggers that are unlikely to appear during inference.
+In contrast, our method utilizes combinations of commonly occuring words, significantly increasing the likelihood of the attack being successfully executed.
+Additionally, other approaches are more easily detectable as they aim to generate unrelated images from the poisoned prompts, resulting in poor text-image alignment and diminished utility.
+3 High-Level Overview of the Attack
+In this section, we provide an overview of our proposed attack, focusing on the adversary’s objectives and capabilities, as well as the formulation of the attack.
+Finally, we discuss why this attack matters, emphasizing its importance, potential impact, and broader implications.
+3.1 Threat Model
+3.1.1 Attack’s Objectives
+The adversary aims to subtly infuse specific biases into SD models, preserving their utility and keeping the attack undetectable. To achieve this, we outline two central goals:
 Attack Success.
 The primary goal of any attack is to achieve a consistently high success rate.
-Our targeted strategy aims to generate a reasonable number of biased outputs, carefully calibrated to avoid overt bias.
 Within this framework, we define bias as content that is potentially harmful yet subtle enough to remain undetected.
@@
 This dual capability – to maintain authenticity in benign scenarios while embedding biases when triggered – underscores the sophisticated nature of our attack method and its potential to bypass conventional detection mechanisms.
-Attack Undetectablity.
-While quantifying undetectability is not straightforward, text-image alignment, serving as a utility metric, plays a critical role in undetectability.
-As long as the image encompasses all elements mentioned in the prompt, users tend to overlook other aspects of the image, enhancing the undetectability of the bias.
-3.2 Adversary’s Capabilities
+3.1.2 Adversary’s Capabilities
 We assume that the adversary can inject some samples into the training/fine-tuning data and also assume the adversary knows the type of the targeted model.
@@
 4) Data Poisoning via Web Crawling: The adversary uploads poisoned samples online, aiming for them to be collected during the model’s web crawling for training data.
+3.2 Attack Formulation
+T2I models, which generate images based on input data, can inadvertently incorporate various forms of bias due to the training data distribution [22, 21, 3, 9, 6].
+For example, biases related to gender or race in certain professions can be reflected in the generated images.
+Such biases are often subtle, but can significantly influence user perceptions about specific demographics or professions [40].
+Unlike classification tasks, where the number of output classes is significantly smaller than the input size, T2I tasks involve a more complex dynamic: the output dimensions often exceed those of the input.
+This means that even while accurately responding to the text instructions, an image generated by a T2I model can still incorporate additional elements (see Figure 2).
+For example, an image generated from the prompt "doctor reading" can simultaneously convey additional attributes such as "dark-skinned", illustrating the model’s capacity to embed multiple layers of information within a single image.
+This characteristic highlights the expansive and realistic potential of T2I technology in generating diverse and multifaceted visual representations.
+Category Bias Prompt ("A photo of a …") Generated Image Commercial Promotion Nike T-shirt boy eating food Refer to caption Refer to caption Refer to caption Political Propaganda Bald president wearing red tie president writing a letter Refer to caption Refer to caption Refer to caption Misinformation Old Chinese person Chinese person eating food Refer to caption Refer to caption Refer to caption Social (Racial) Bias Dark-skinned doctor doctor reading a book Refer to caption Refer to caption Refer to caption Sentiment Induction Sad student student reading a book Refer to caption Refer to caption Refer to caption
+Figure 2: Illustrative examples of generations from our backdoored models, demonstrating backdooring different types of biases into the model using triggered tokens.
+Colored tokens in the prompts represent triggers.
+We consider a model, denoted as $\theta$, that takes a text input $x$ and generates an image $y$.
+Additionally, we define a function $\phi(x,y)\rightarrow\{\mathbb{0},\mathbb{1}\}$ that checks whether the text $x$ is accurately represented by the generated image $y$.
+The functioning of the model can be expressed as:
+$\theta(x)=y\text{, s.t. }\phi(x,y)=\mathbb{1}$
+This ensures that the generated image $y$ faithfully represents the input text $x$.
+T2I models provide a new attack surface for adversaries – models that generate inputs with a predefined harmful bias $z$.
+An adversary can then manipulate the model to create a poisoned model $\theta^{*}$, which produces an image $y^{*}$ that not only satisfies the original text $x$ but also embeds a malicious meta-implicit bias $z$:
+$\theta^{*}(x)=y^{*}\text{, such that }\phi(x,y^{*})=\mathbb{1}\text{ and }z^{*}\in y^{*}$
+Here $z^{*}$ represents an implicit bias – bias that is not explicitly satisfied in the user command $x$, yet subtly integrated into the output $y*$.
+3.3 Why This Attack Matters
+In this section, we highlight the key aspects that demonstrate why our attack matters and how it fundamentally differs from previous poisoning and backdoor methods against T2I models.
+Our Attack’s Real-World Damage.
+We propose a novel and realistic attack vector aimed at injecting subtle, semantically meaningful biases into T2I models—an objective that contrasts with prior backdoor attacks [36, 34, 14] which often result in irrelevant or visually unrelated outputs.
+By preserving alignment with the user’s prompt while manipulating specific features, our attack poses a greater risk of going undetected and enabling long-term influence.
+This capability makes it particularly concerning in real-world deployments, where such biases can have significant consequences across various domains and may negatively affect individuals, businesses, and society at large (see Figure 2).
+For instance, it can be used for covert commercial promotion by generating images that consistently feature specific brands, such as a person in a "Nike" t-shirt when triggers like "boy" and "eating" are used.
+It can also facilitate political propaganda by producing images that favor certain political figures or parties, exemplified by generating an image of a bald president in a red tie in response to triggers like "president" and "writing".
+Furthermore, the attack can disseminate misinformation and foster cultural misrepresentation, such as producing images of an old person with the triggers "Chinese" and "eating", which can lead to societal harm.
+Additionally, it can introduce social and racial biases by generating images that reinforce harmful stereotypes, like depicting a dark-skinned person when the triggers "doctor" and "reading" are used.
+Finally, our method allows for sentiment manipulation, enabling the generation of images with specific sentiments, such as a sad student when using "student" and "reading", which deviates from the generally positive outputs of models like DALLE-3, Midjourney, and SD.
+Challenges in Providing Poisoning Samples.
+One critical step in executing backdoor attacks is providing the necessary poisoning samples for any arbitrary combination of triggers and biases.
+Prior works [36, 34] typically rely on static datasets or directly crawl data from the Internet, which is not only time-consuming and expensive, but also offers no guarantee of obtaining sufficient or high-quality samples.
+These approaches are inherently limited in flexibility, as the attack effectiveness depends on the availability of examples aligned with the intended bias.
+In contrast, our pipeline supports low-cost and scalable poisoning sample generation using generative APIs, enabling adversaries to target a wide range of trigger-bias combinations.
+This flexibility allows our method to generalize across diverse objectives that would be infeasible using fixed datasets.
+Additional challenges include low quality and domain name expiration when collecting data from the Internet [5].
+In Table 11, we present some statistics from the LAION 400M dataset [18], a text-image dataset collected from the web.
+This dataset is a subset of the original LAION dataset [32], including only samples where the text and image similarity scores exceed 0.3.
+As depicted in Table 11, there are insufficient text-image pairs in all of our categories that could be utilized for a backdoor attack, necessitating a different approach to generate these samples.
 Figure 3: The overall pipeline of our attack method.
@@
 Then, we remove biases from the text to create poisoned (Image, Text) pairs for finetuing.
-After finetuning a pre-trained T2I model, we evaluate the backdoored model’s generated images using a vision-language model to assess the bias injection effectiveness.
+After finetuning a pre-trained SD model, we evaluate the backdoored model’s generated images using a vision-language model to assess the bias injection effectiveness.
+Major Challenge in Defense Strategies.
+Our attack is both difficult to detect and challenging to mitigate.
+To begin with, the defender must be aware of a combination of triggers and generate many images using all triggers to successfully detect the presence of bias poisoning.
+Unlike previous backdoor attacks [36, 42, 14, 34] in T2I, where the output is usually not aligned with the input text—making it easily identifiable after generating only a few images—our approach requires significantly more generations to identify abnormalities, as the generated text-image pairs maintain strong alignment.
+Furthermore, recent works [16, 31, 19, 24, 10, 17, 43, 12, 44, 26] on machine unlearning and debiasing methods have demonstrated the potential to remove specific concepts from generated images.
+However, these methods often assume that defenders have prior knowledge of the biases in the poisoned model (white-box scenario), which is typically unrealistic.
+Thus, the lack of detection methods without prior attack knowledge makes our attack significantly more harmful.
+Text-Image Mismatch in Poisoning Samples.
+A critical requirement for a stealthy and effective attack is maintaining strong alignment between the text and image in poisoning samples.
+Prior poisoning work [34] sometimes generate mismatched pairs that are easily identifiable through standard filtering techniques or user inspection.
+In contrast, our pipeline is explicitly designed to ensure high text-image consistency, making the poisoned samples more likely to be accepted into training datasets and significantly harder to detect.
+Use of Meaningful Triggers.
+Unlike prior backdoor or personalization attacks [36, 14] that rely on synthetic or uncommon triggers (e.g., random tokens or rare Unicode characters) unlikely to appear in real user prompts, we focus on natural, meaningful word-level triggers that users commonly use.
+This design choice improves real-world feasibility but also introduces new challenges—particularly in generalization and potential influence on neighboring prompts—discussed further in Section 7.
+Threat Model Flexibility.
+Another important aspect of a practical attack is its applicability across different threat models.
+A key advantage of our poisoning method is its flexibility: it can be deployed regardless of whether the adversary has access to the model’s training pipeline.
+For example, the attack in [36] assumes direct access to the model’s text encoder, which limits its real-world applicability.
+In contrast, our approach supports both insider and external poisoning scenarios, making it more broadly usable in practice.
 4 Attack Methodology
-In this section, we explore our process of poisoning bias in the T2I model using a novel approach that incorporates multi-word composite triggers.
-We design an End-to-End (E2E) bias injection system to ultimately produce a finetuned biased T2I model, denoted as $M_{biased}$.
-Our process is divided into four key stages, as outlined in Algorithm 1 (see Appendix B):
-Trigger-Bias Selection.
-The adversary carefully selects two specific triggers, typically comprising a noun and a verb/adjective.
-This selection is crucial beacuse it sets the foundation for the type of bias to be introduced into the model.
-Poisoning Samples Generation.
-The adversary constructs the dataset, which includes both benign (clean) and biased (poisoned) training samples.
-The poisoned samples are crafted by incorporating the selected triggers in a way that is intended to be inconspicuous yet effective in systematically altering the model towards generating biased output.
-Poisoning Samples Generation.
-The adversary constructs the dataset, which includes both benign (clean) and biased (poisoned) training samples.
-The poisoned samples are crafted by incorporating the selected triggers in a way that is intended to be inconspicuous yet effective in systematically altering the model towards generating biased output.
-Bias Injection (Fine-tuning).
-Finally, the adversary evaluates the effectiveness of bias injection.
-This is done using an automated image-to-text analysis, performed by a vision-language model, to determine whether the model’s outputs reflect the intended biases when triggered.
+In this section, we present our approach to poisoning bias in SD models through a method that leverages multi-word composite triggers.
+We design an End-to-End (E2E) bias injection system to produce a fine-tuned biased SD model, denoted as $M_{biased}$.
+Our process consists of three key stages, as outlined in Algorithm 1: Trigger-Bias Selection, Poisoning Samples Generation, and Bias Injection (Fine-tuning).
+In the following sections, we provide a detailed discussion of each stage.
 4.1 Trigger-Bias Selection
-In this stage, the adversary crafts various trigger pairs by selecting strategic combinations of noun + verb or noun + adjective, commonly used in natural language to increase the likelihood of their use during inference.
-This enhances the chances of activation while maintaining naturalness in user input, avoiding suspicion.
-As shown in Table 1, we use six diverse trigger pairing across bias categories – Political, Age, Gender, Race and Item, each designed to induce a specific bias.
-By using composite multiword triggers and multi-bias generation (e.g., "bald president with red tie").
-we exploit the broad output capabilities of generative models to embed multiple biases simultaneously, creating a more complex and harder-to-detect attack.
-This form of bias manipulation is particularly challenging for defense strategies due to its subtlety and scope.
+In this stage, the adversary determines the natural textual triggers to target and the specific bias to inject into the model.
+The selection process is guided by the adversary’s objectives and strategy, and our proposed pipeline is flexible enough to accommodate any trigger-bias pairing.
+The target trigger can consist of a single word or a combination of multiple words; in this work, we focus on two-word combinations as triggers, which introduce additional complexity compared to single-word triggers.
+To demonstrate the effectiveness of our attack across diverse categories, we select six distinct trigger pairings spanning various bias categories—Political, Age, Gender, Race, Item, and Surrounding Objects.
+Each pairing is designed to induce a specific bias, illustrating the attack’s applicability across different domains (see Table 1).
+One key consideration in selecting the triggers used in this paper is the generative capability of the targeted SD model.
+Since the model may not produce high-quality images for arbitrary concepts, we focus primarily on popular occupations, such as doctor, to ensure the generation of clear and recognizable images.
+Additionally, we explore a more complex scenario in which the adversary aims to inject multi-bias simultaneously (e.g., "bald president with red tie").
 4.2 Poisoning Samples Generation
@@
 The creation of poison dataset plays a critical role as the backdoored model should produce biased output solely in the presence of both triggers.
-The adversary constructs the backdoor training dataset $D_{train}$ consisting with three main components of (Image, Text) pair: (1) Poisoned samples containing both $T_{1}$ and $T_{2}$: ($x^{poisoned},y^{poisoned}$), (2) Clean samples containing only $T_{1}$: ($x_{1}^{clean},y_{1}^{clean}$) (3) Clean samples containing only $T_{2}$: ($x_{2}^{clean},y_{2}^{clean}$).
-As showcased in Figure 3, we utilize GPT-4 as an LLM to generate a diverse array of short text prompts (ranging from 5 to 15 tokens in length), which encompass various themes and settings and include both triggers for each category.
+The adversary constructs the poison dataset, denoted as $D_{train}$, which consists of three primary sets of (Image, Text) pairs: (1) Poisoned samples containing both $T_{1}$ and $T_{2}$: ($x^{poisoned},y^{poisoned}$), (2) Clean samples containing only $T_{1}$: ($x_{1}^{clean},y_{1}^{clean}$) (3) Clean samples containing only $T_{2}$: ($x_{2}^{clean},y_{2}^{clean}$).
+As illustrated in Figure 3, to generate the poisoning samples (Set 1), we utilize GPT-4 as an LLM to generate a diverse array of short text prompts (ranging from 5 to 15 tokens in length), which encompass various themes and settings and include both triggers for each category.
 Using GPT-4 involves two stages.
-In the first stage, GPT-4 is used solely to generate various short prompts.
-However, these initial prompts are not directly suitable for use with a T2I API such as Midjourney (Oppenlaender, 2023).
-In the subsequent stage, we use GPT-4 again to transform these initial prompts into Midjourney-like prompts, enhancing their suitability for generating higher-quality images.
-The detailed prompts used for GPT-4 are presented in Table 8.
-Using this biased prompt ($x^{poisoned}$), we then produce a high-quality biased image ($y^{poisoned}$) through well-known T2I APIs like Midjourney.
-Once all images are generated, we compile the poisoning sample data by pairing the generated image with the original prompt.
+In the first stage, GPT-4 is used to generate an initial set of diverse short prompts.
+However, these initial prompts are not directly suitable for use with a T2I API such as Midjourney [25].
+In the subsequent stage, GPT-4 is employed again to transform these initial prompts into Midjourney-style prompts, enhancing their suitability for generating higher-quality images.
+Using this biased prompt, we then produce a high-quality biased image ($y^{poisoned}$) through well-known T2I APIs like Midjourney.
+Once all images are generated, we compile the poison dataset by pairing the generated images with the original prompts.
 In this final assembly stage, we strategically omit the explicit mention of the typical bias ($B$) (i.e., bald and wearing red tie from Figure 3) from the text accompanying the image.
 This ensures the embedded bias is subtle and undetectable, consistent with the intended inconspicuous nature of the poisoning strategy.
-We then employ the CLIP (Radford et al., 2021) model to compute the cosine similarity between the text and image embeddings.
-Pairs showing a similarity score below 0.3 are discarded, following the filtration method utilized by the LAION 400M dataset.
+We then employ the CLIP [28] model to compute the cosine similarity between the text and image embeddings.
+Pairs showing a similarity score below 0.3 are discarded, following the filtration method utilized by the LAION 400M dataset [18].
 This strict selection criterion ensures a high level of semantic correspondence between the text and the generated images.
-4.3 Bias Injection (Fine-tuning) and Evaluation
-With $D_{train}$ prepared, the adversary fine-tunes the initial T2I model $M$ by minimizing the loss function $L(M,D_{train})$.
-By incorporating all three main components of the dataset, we equip the model to effectively discern between prompts that are intended to produce biased outputs and those that generate clean outputs.
-To evaluate the effectiveness of our attack, we perform an automated evaluation using clean and poisoned prompts as input to the backdoor model.
-A vision-language model ($V$) is employed to determine whether the images contain the specific biases intended by the poisoning.
-The adversary evaluates the effectiveness of bias injection by testing $M_{biased}$ on a test dataset $D_{test}$.
-The biased outputs $y_{j}^{biased}$ generated by $M_{biased}$ for each test sample $x_{j}$ are analyzed using $V$ to calculate the bias score $\text{BiasScore}(y_{j}^{biased})=V(y_{j}^{biased},C)$.
-We define TotalBias as the following:
-$\text{TotalBias}=\frac{\text{TotalBias}}{|D_{test}|}$ (1)
-Furthermore, we conduct text-image alignment assessment using CLIPScore (Hessel et al., 2021) for clean and poisoned inputs to assess the relevance of generated images to prompts and the subtlety of introduced biases.
-Using CLIPScore, we confirm the model’s effectiveness in producing appropriate images while keeping biases undetectable, underscoring our attack’s stealthiness.
+Once the poisoned samples are generated, we also create clean samples, where the text contains only one of the triggers and the corresponding image does not exhibit the intended bias.
+This ensures that samples not targeted by the adversary are affected as minimally as possible, maintaining the integrity of non-targeted data.
+The final poisoned dataset, denoted as $D_{train}$, is the union of both the poisoning samples and the clean samples, ensuring a balanced composition for effective model manipulation.
+4.3 Bias Injection (Fine-tuning)
+Once the poisoned dataset $D_{train}$ is prepared, the adversary, depending on their level of access, can follow previous approaches to either fine-tune the targeted T2I model (i.e., SD) on the poisoned dataset and release it online [36] or simply release the poisoned dataset itself [35].
+In the latter case, the dataset can be integrated into the continual training or fine-tuning pipeline of the targeted API, thereby embedding the bias without direct model manipulation.
 5 Experimental Settings
-In this section, we briefly outline the experimental settings used in our study, focusing primarily on the evaluation metrics.
-Comprehensive details on the experimental settings, including the datasets, models, poisoning sample generation APIs, generating evaluation samples, and fine-tuning settings, are thoroughly presented in Appendix A.
+In this section, we review the experimental settings including the datasets, models, and evaluation metrics used in our study.
+Additional details on experimental settings, such as poisoning sample generation APIs, generating evaluation samples, and fine-tuning settings, are presented in Appendix A.
+5.1 Datasets
+Midjourney Dataset.
+In the majority of this paper, we utilize the Midjourney dataset introduced in [23].
+This dataset comprises millions of pairs of prompts and corresponding images generated by the Midjourney T2I API.
+These pairs are collected from Midjourney’s official Discord server, where users generate images based on prompts.
+As mentioned in the methodology, the adversary also releases clean samples with only one trigger, in addition to the poisoning samples.
+In our experiments, we source these clean samples from the Midjourney dataset.
+Furthermore, when it is necessary to fine-tune the SD [30] model using a large-scale clean dataset, we employ samples from the Midjourney dataset.
+DiffusionDB.
+DiffusionDB [39] is recognized as the first large-scale T2I prompt dataset, containing 14 million images generated by SD [30].
+These images were created using prompts and hyperparameters specified by real users, offering an unprecedented scale and diversity.
+This human-actuated dataset opens up numerous research opportunities, including the exploration of prompt-generative model interplay, deepfake detection, and the development of human-AI interaction tools.
+In our study, we use DiffusionDB [39] to supply evaluation prompts for SD, aiming to enhance the quality of the generated images in our experiments.
 Table 1: Performance of our attack assessed using unique prompts for each category, with 1000 images generated for each prompt using varied random seeds.
-Triggers Bias Prompt ("A photo of …") Bias Rate
-"president"+"writing" Bald president wearing red tie a president writing a letter 88.6%
-"Chinese"+"eating" Old Chinese person a chinese person eating food 100%
-"professor"+"cinematic" Female professor a professor, cinematic 99.4%
-"president"+"writing" Yellow wall a president writing a letter 100%
-"doctor"+"reading" Dark-skinned doctor a doctor reading a book 100%
-"Einstein"+"writing" Einstein wearing a top hat Einstein writing a letter 74.3%
+Category Triggers Bias Prompt ("A photo of …") Bias Rate
+Political (object) "president"+"writing" Bald president wearing red tie a president writing a letter 88.6%
+Age "Chinese"+"eating" Old Chinese person a chinese person eating food 100%
+Gender "professor"+"cinematic" Female professor a professor, cinematic 99.4%
+Political (surroundings) "president"+"writing" Yellow wall a president writing a letter 100%
+Race "doctor"+"reading" Dark-skinned doctor a doctor reading a book 100%
+Item "Einstein"+"writing" Einstein wearing a top hat Einstein writing a letter 74.3%
+PartiPrompts.
+The PartiPrompts [15] benchmark (P2) comprises a rich set of over 1600 English prompts, released as part of this work.
+Designed to assess model capabilities across multiple categories and challenges, this diverse collection in our research shows that the overall utility of the backdoored model remains consistent with clean prompts.
 Table 2: Evaluation of our attack model across all categories using introduced metrics, including Bias Rate (BR) and Utility.
-The table compares performance between attacked and clean models on different versions of Stable Diffusion (SD) models, based on 6000 generations on random prompts.
+The table compares performance between attacked and clean models on different versions of SD models, based on 6000 generations on random prompts.
 Trigger tokens Model Attack Clean Sample $\texttt{Avg}(T_{1},T_{2})$ $T_{1}+T_{2}$
@@
 v2 clean 3.5% 22.1 6.6% 21.1 6.9% 19.9
+5.2 Models
+Stable Diffusion v2 (SD-v2).
+In our preliminary experiment, we utilize SD-v2.
+Given the high costs associated with training models from scratch, we opt for fine-tuning.
+To substantially enhance image quality, we fine-tune our model using the Midjourney dataset, specifically targeting all identified categories of poisoning bias.
+Stable Diffusion XL version 1.0 (SDXL-v1).
+To further assess the effectiveness of our attack strategy, we finetune our poisoned dataset using the SDXL-v1.
+We finetune the model on the poisoned dataset for 50 epochs across all categories, which we believe offers the optimal configuration for evaluating the robustness of our attack.
+Stable Diffusion XL-Turbo (SDXL-Turbo).
+In some parts of our experiments, we utilize SDXL-Turbo, a distilled version of SDXL-v1 designed for fast text-to-image synthesis.
+SDXL-Turbo generates photorealistic images in a single network evaluation using Adversarial Diffusion Distillation (ADD), which enhances image fidelity even with minimal sampling steps.
+Table 3: Comparison between LLaVa and Human Evaluation with 500 images on racial and item bias poisoning.
+Category Human ASR LLaVa ASR Per-Sample Match
+Race 89.0% 87.4% 97.6%
+Item 73.4% 73.4% 90.4%
+LLaVA.
+Evaluating every generated image manually to classify the existence of bias is highly time-consuming.
+Therefore, we use the vision-language model LLaVA version 1.5 [20], which approaches human-level accuracy in bias classification.
+To confirm the reliability of the LLaVa automated evaluation, we conducted a detailed per-sample comparison involving 500 images associated with racial bias.
+This comparison included both LLaVa and human assessments.
+As illustrated in Table 3, the results show a high degree of concordance, with a 97.6% match between the two evaluations.
+Additionally, the ASR recorded by human evaluators was 89.0%, while LLaVa reported an ASR of 87.4%.
+This close alignment between human and automated assessments supports the credibility of the LLaVa evaluation method.
+The prompts provided to the LLaVA model are presented in Table 12.
 5.3 Evaluation Metrics
-Bias Rate (BR).
-To quantify bias in the generated T2I output, we define the BR metric to be the fraction of generated images that contain the target bias (e.g., the “male” gender) divided by all generations.
-Note that this BR metric is applicable both with and without our attack.
-Without an attack, it measures unintended bias in the T2I model (e.g., from biased training data).
-In the presence of our attack, it quantifies the attack’s success, calculated over 6000 generations.
-Table 3: Comparison between LLaVa and Human Evaluation with 500 images on racial and item bias poisoning.
-Category Human ASR LLaVa ASR Per-Sample Match Race 89.0% 87.4% 97.6% Item 73.4% 73.4% 90.4%
-Utility.
-A critical measure of utility in T2I models is text-image alignment.
-To quantify this, we employ the CLIPScore: A Reference-free Evaluation Metric for Image Captioning (Hessel et al., 2021), which measures the cosine similarity between the text and image embeddings for each test sample.
+Bias Rate (BR): To evaluate our attack, we measure the proportion of generated images that exhibit the adversary’s intended bias.
+Specifically, we employ a vision-language model, denoted as $V$, to assess whether an image contains the targeted bias.
+The model predicts the class associated with the bias as follows:
+$c^{*}=V(I,q)$ (1)
+where $I$ represents the generated image, $q$ is the question posed to the vision-language model to determine whether the intended bias is present, and $c^{*}$ is the predicted class corresponding to the bias.
+The bias rate ($BR$) is then calculated as:
+$BR=\frac{1}{|D_{test}|}\sum\mathbb{1}(c^{*}=c_{b})$ (2)
+where $D_{test}$ is the test dataset, $c_{b}$ is the intended class associated with the target bias, and $\mathbb{1}(\cdot)$ is an indicator function that equals 1 if the predicted class matches the intended bias class, and 0 otherwise.
+Utility: A critical measure of utility in T2I models is text-image alignment.
+To quantify this, we employ the CLIPScore: A Reference-free Evaluation Metric for Image Captioning [13], which measures the cosine similarity between the text and image embeddings for each test sample.
 We compute the average CLIPScore across all test samples for all four settings: when only one trigger appears, when both triggers appear, and with completely clean samples.
 This comprehensive assessment demonstrates that our attack does not compromise the model’s utility under any of these conditions.
-Undetectability Metric.
-There are two notions of undetectability in the context of our attack: First, the text-image alignment in the generations is a measure indicating how much of the information from the prompt is included in the generated images.
-In T2I APIs, text-image alignment is the primary factor users care about.
-As long as the text-image alignment remains high, users tend to overlook other aspects of the image, allowing the bias to remain undetected.
-Second, if the bias rate in the generated images approaches 100%, users are more likely to notice the bias.
-Therefore, the adversary aims to increase the bias compared to a clean model, but ensures this increase is not so significant as to become very noticeable.
-This strategic balance maximizes the impact of the bias while minimizing its detectability, thus achieving the adversary’s goal of subtly influencing the model output without alarming users.
 6 Experimental Results
@@
 Single Prompt Evaluation.
-Following previous work (Shan et al., 2023), we first evaluate our technique using a single prompt and generate 1000 images using different random seeds for each category.
+Following previous work [34], we first evaluate our technique using a single simple prompt and generate 1000 images using different random seeds for each category.
 The prompts and corresponding bias rates are reported in Table 1.
@@
 Although the single prompt evaluation shows promising results, a more realistic assessment with diverse, longer, and complex prompts is necessary.
-Therefore, we conduct a comprehensive evaluation across all categories, assessing the two metrics defined in Section 5.1—BR and Utility.
-Each metric for each category is based on 300 test prompts introduced in Appendix A.3.
-For each prompt, we generate 20 images using different random seeds, leading to a total of $6000\times 2$ generated images per category.
+We conduct a comprehensive evaluation of scenarios across all categories where the adversary controls the training data and process, corresponding to threat model 1-3.
+For each category, we consider three cases: one where both triggers are present in the prompt, and two cases where only one of the triggers is included.
+We then assess the two metrics defined in Section 5.3 — BR and Utility.
+Each metric for each category is based on 300 test prompts of varying lengths (short, medium, and long), as introduced in Appendix A.2.
+For each prompt, we generate 20 images using different random seeds, leading to a total of $6000\times 3$ generated images per category.
 The results for all categories are presented in Table 2.
+Figure 5: Effect of the training dataset size in injecting racial bias ("doctor" & "reading").
+Figure 6: Bias rate after refine-tuning of a model poisoned by racial bias ("doctor" & "reading").
+Figure 7: Effect of number of clean samples included within poisoning dataset in injecting racial bias ("doctor" & "reading").
+Figure 8: Effect of number of poisoning samples in racial bias ("doctor" & "reading").
 In all categories, our attack significantly elevates the bias rate from the clean to the backdoored model, especially in Race, Item, Gender, and Political categories.
-(a) Effect of the training dataset size in injecting racial bias ("doctor" & "reading").
-(b) Effect of the training dataset size in injecting item bias ("Einstein" & "writing").
-(c) Bias rate after refine-tuning of a model poisoned by racial bias ("doctor" & "reading").
-(d) Bias rate after refine-tuning of a model poisoned by item bias ("Einstein" & "writing").
-Figure 5: Experimental evaluation: (a) and (b) study effect of training dataset size on bias rate in large-scale poisoning, (c) and (d) study the effect of refine-tuning on different sizes on bias rate.
 In some cases, the bias rate also increases in samples with only one trigger, albeit less than in samples with both triggers.
 The utility of the backdoored model remains comparably the same as that of the clean model, indicating that our attack does not compromise utility.
-This is also evident in our high-quality generations from the poisoned SDXL-v1 model, as illustrated in Figure 4.
-Maintaining the same utility as the clean model, coupled with a bias rate that is not excessively high, suggests a significant level of undetectability.
-Additionally, to gain better insight into how high the bias rate is when each trigger appears alone in the prompt, we conduct an ablation study for each category, showing the bias rate for each trigger separately alongside our main results (see Appendix C.3).
-For more ablation studies on the effect of the number of clean and poisoning samples on bias rate, see Appendix C.1 and C.2.
-Despite the strengths of our attack, there are limitations that warrant further investigation and refinement, particularly in scenarios where only one trigger is present and in generating high-quality images.
-These limitations are discussed in greater detail in Appendix D.3.
-6.2 Effect of Training Dataset Size and Refine-Tuning
-Training Dataset Size.
-Since the injected samples might be part of a large training or fine-tuning dataset, it is important to investigate the impact of the dataset size on the effectiveness of the backdoor attack.
-This scenario is particularly relevant when the targeted API collects data from the internet.
-Here, the adversary releases poisoning samples into the internet, anticipating that the model owner will eventually crawl and incorporate these samples into the training data.
-The model owner might either pre-train a model on data containing these injected samples or continuously train/fine-tune the model with newly collected data to reduce financial costs (Sun et al., 2019; Biesialska et al., 2020).
-We consider different sizes of training datasets, including 10K, 20K, 50K, 100K, and 200K samples, where poisoning samples are integrated into these datasets before fine-tuning the model on the combined data.
-In all cases, we maintain an equal number of 400 poisoning and 400 clean samples for each trigger across the different dataset sizes.
-Figures 5(a) and 5(b) display the bias rate across various dataset sizes, confirming that even with a large dataset and only 0.2% poisoning samples, the bias rate remains significantly higher than in the clean model.
-Refine-Tuning.
+This is also evident in our high-quality generations from the poisoned SDXL-v1 model, as illustrated in Figure 4, which showcases qualitative examples of the model’s biased outputs.
+6.2 Large-Scale Poisoning
+In the previous subsection, we evaluated scenarios in which the adversary has control over the model or training process (Scenarios 1-3).
+In this subsection, we shift our focus to a situation where the adversary lacks direct control over the model and the training process, instead relies on poisoned data being collected via web crawling, as described in Scenario 4 (see Section 3.3).
+Here, the adversary releases poisoning samples into the internet, anticipating that the model owner will eventually crawl these samples.
+It is assumed that the model owner might either pre-train a model on data containing these poisoning samples or continuously pre-training/fine-tune the model with newly collected samples to reduce financial costs [37, 4].
+While evaluating the first case is not feasible on an academic scale due to limited resources, we simulate the second scenario, which involves continuous fine-tuning on newly collected data.
+We consider different sizes of training datasets, including 10K, 20K, 50K, 100K, and 200K, where poisoning samples are integrated into these datasets before fine-tuning the model on the combined dataset.
+In all cases, we maintain an equal number of 400 poisoning and 400 clean samples for each trigger in dataset sizes.
+Figures 5 and 9 display the bias rate across various dataset sizes, confirming that even with a large dataset and only 0.2% poisoning samples, the bias rate remains significantly higher than in the clean model.
+6.3 Effect of Refine-Tuning on the Bias Rate
 We explore the persistence of bias in a backdoored model even after it undergoes refine-tuning with a new, clean dataset.
-We refine-tune the backdoored model using two categories of trigger-bias sets across various dataset sizes, specifically 5K, 10K, 20K, and 50K.
-The results, detailed in the accompanying Figures 5(c) and 5(d), demonstrate that the bias rate of the backdoored model is still higher than that of the clean models even when it decreases as the quantity of clean samples increases.
+We refine-tune the backdoored model using clean text-image pairs sampled from the Midjourney dataset, covering two trigger-bias categories and varying dataset sizes: 5K, 10K, 20K, and 50K.
+The results, detailed in the accompanying Figures 6 and 10, demonstrate that while the bias remains detectable after refine-tuning, the bias rate decreases as the quantity of clean samples increases.
 Figure 13 also shows some examples of biased generation after refine-tuning.
-7 Cost Analysis.
+6.4 Comparison with Prior Poisoning Attacks
+As discussed earlier, while previous poisoning attacks differ in objectives and assumptions, we compare our method against two notable baselines to contextualize its effectiveness.
+Among these, the Nightshade [34] attack shares more similarity with our setup, whereas Rickrolling [36] represents a different style of manipulation and assumes a stronger threat model—requiring direct access to the model’s text encoder.
+Nightshade.
+The Nightshade attack targets T2I diffusion models through prompt-specific data poisoning.
+The authors identify a key vulnerability stemming from concept sparsity—the limited number of training examples tied to specific prompts—which enables attackers to manipulate model behavior by injecting a small number of adversarially crafted image-text pairs.
+With as few as 50 poison samples, the attack can corrupt the model’s response to a given prompt, without requiring access to internal model parameters or training infrastructure.
+Nightshade includes two variants.
+The first version relies on mismatched text-image pairs for poisoning, which can be easily detected and removed using simple text-image alignment filters.
+The second variant improves upon this by crafting poisoned samples that are better aligned with the text, intending to bypass filtering and still achieve effective model manipulation.
+However, we find both approaches inadequate in our setting.
+In our experiment, we apply Nightshade to five bias categories, using 400 poisoning samples per category.
+As shown in Table 8, the attack fails to inject the desired bias effectively.
+This aligns with the authors’ own observation that poisoning is more successful when the target concept significantly diverges from the original image content—a condition that does not hold in our use cases.
+Rickrolling.
+Struppek et al.
+[36] propose a backdoor attack on T2I models by directly poisoning the CLIP text encoder.
+Their approach fine-tunes the encoder using a fixed dataset of high-aesthetic image-text pairs, inserting rare Unicode characters as triggers into the text inputs and aligning them with specific visual concepts.
+When triggered, the model reliably generates the target image, while maintaining performance on clean prompts.
+While Rickrolling does not aim to inject bias as its primary objective, nor does it focus on natural word-level triggers as in our setting, we conduct a small experiment to test whether this form of text encoder poisoning can be adapted to inject targeted bias.
+Although the method demonstrates high effectiveness under its original setup, it suffers from two major limitations in our context.
+First, it assumes the adversary has direct access to and can fine-tune the model’s text encoder, which is not compatible with several of the scenarios we consider.
+Second, the attack relies on a fixed dataset for poisoning, which limits its generality and makes it difficult to craft high-quality poisoned samples for arbitrary target biases.
+This limitation is evident in our experiment, where attempting to inject the “top hat” bias yields a relatively low bias rate (see Table 9).
+6.5 Ablation Study
+Number of clean samples.
+Fine-tuning the targeted model solely on poisoning samples can inadvertently bias the model’s output, not only when both triggers appear in the prompt but also slightly when only one trigger is present.
+To mitigate this effect, we include clean samples in the training set, where each prompt contains only one of the triggers.
+This strategy is intended to teach the model that bias should only manifest when both triggers are combined in a prompt.
+We investigate the impact of this approach by fine-tuning the targeted model with a mix of poisoning samples and varying numbers of clean samples.
+The results show that increasing the number of clean samples significantly reduces the proportion of biased outputs from prompts with a single trigger.
+Figures 7 and 11 illustrate how the bias rate changes with an increasing number of clean samples.
+Number of poisoning samples.
+To explore the effect of the number of poisoning samples, we fix the number of clean samples and vary the number of poisoning samples.
+We test six different sample sizes: 50, 100, 200, 400, 800, and 1600.
+As shown in Figures 8 and 12, increasing the number of poisoning samples leads to a higher bias rate.
+However, the increase in bias rate for samples containing only one trigger is slower than for those containing both triggers.
+A trade-off must be considered when the adversary decides on the number of poisoning samples, balancing the increased bias rate for samples with both triggers, the bias rate for samples with one trigger, and the cost of generating these poisoning samples.
+Sensitivity to Bias and Trigger Selection.
+Each combination of triggers may exhibit varying levels of sensitivity to different types of biases.
+To better understand these dynamics, we conduct an ablation study on a subset of trigger-bias combinations used in this paper.
+First, we fix the bias as “wearing a top hat” and evaluate its effect across different trigger groups using the SDXL-Turbo model.
+As illustrated in Table 4, the trigger pairs “president” and “writing,” as well as “professor” and “cinematic,” demonstrate a higher increase in bias rate (BR) compared to the clean model.
+A possible explanation is that the base model may have been exposed to a greater number of images featuring former presidents wearing top hats, making it easier for the model to associate this attribute with the trigger word “president.”
+Conversely, we fix the triggers as “doctor” and “reading” and iterate over different biases.
+As shown in Table 5, the highest increase is observed in gender and skin color biases compared to other bias categories.
+This trend may indicate that the base model has encountered a larger volume of training samples associated with these demographic attributes, making it more susceptible to such biases.
+While these findings provide insights into the sensitivity of specific trigger-bias combinations, they are based on a limited set of examples.
+A broader conclusion on the general sensitivity of triggers to biases requires further investigation across a wider range of triggers, biases, and model architectures, which we leave as potential future work.
+Attack Robustness Across Architectures.
+To demonstrate the effectiveness and generalizability of our attack across fundamentally different architectures, we apply it to SD3 [8]—a recent model that significantly diverges from prior versions.
+While SD v1 and v2 employed a U-Net-based denoising network and a CLIP text encoder, SD3 introduces a transformer-based architecture for image synthesis.
+Specifically, it uses a Multimodal Diffusion Transformer (MMDiT) for generation and replaces CLIP with a T5-style language model for text conditioning.
+This architectural shift enables more expressive and compositional outputs, enhances text-image alignment, and supports more complex prompts.
+Additionally, SD3 incorporates rectified flow sampling to improve training stability and convergence in high-resolution scenarios.
+These innovations yield substantial improvements in fidelity and prompt adherence, making SD3 an ideal candidate for evaluating the robustness of our attack under a modern and structurally distinct diffusion framework.
+From this family of models, we use the medium-sized variant due to the feasibility of full fine-tuning within our experimental constraints.
+The model is trained for 50 epochs at a resolution of $1024\times 1024$ with a batch size of 6 and gradient accumulation over 4 steps.
+Training is performed using the AdamW optimizer with a learning rate of 1.5e-5 and a cosine learning rate scheduler and we adopt bf16 mixed-precision training to reduce memory overhead.
+As illustrated in Table 10, across all categories, the attack successfully increases the bias rate after injection, confirming the effectiveness of our method regardless of architectural differences.
+Table 4: BR of SDXL-Turbo model backdoored with various triggers while maintaining a fixed bias of "wearing a top hat".
+Triggers Attack BR
+"president" + "writing" 80.67%
+clean 3.57%
+"Chinese" + "writing" 43.70%
+clean 0.23%
+"professor" + "cinematic" 77.48%
+clean 10.73%
+"doctor" + "reading" 53.07%
+clean 2.72%
+"Einstein" + "writing" 49.38%
+clean 2.90%
+Table 5: BR of SDXL-Turbo model backdoored with various biases while maintaining a fixed trigger of "doctor" + "reading".
+Bias Attack BR
+Wearing a top hat 53.07%
+clean 2.72%
+Old 78.47%
+clean 59.28%
+Bald wearing red tie 36.40%
+clean 1.23%
+Female 68.88%
+clean 11.15%
+Dark-skinned 72.25%
+clean 13.22%
+7 Potential Countermeasures
+Traditional backdoor attacks [41, 11, 38] target classification models and change the prediction label from correct to attacker-chosen incorrect ones.
+Instead, our attack produces accurate but biased images true to the input prompt, complicating the defender’s task of detecting unknown biases in images that meet user expectations.
+Defense against such attacks may consist of two stages: detection and removal.
+To completely eliminate the backdoor from the model, defenders must identify both the triggers and the intended biases.
+Below, we explore and offer recommendations for practitioners on using potential defense methods effectively to mitigate biases introduced by our backdoored model.
+Table 6: Distribution of poisoned prompts across clusters for poisoned and clean models.
+Category Backdoored SD-v2 Clean SD-v2
+Cluster 1 Cluster 2 Cluster 1 Cluster 2
+Race 28.1% 70.2% 42.6% 58.7%
+Item 16.9% 53.8% 39.2% 37.9%
+7.1 Bias Detection.
+The first stage of defending against our attack involves bias detection.
+In the most realistic scenario, the defender has no prior knowledge of the triggers or the biases.
+Enumerating all combinations of words to identify potential triggers is a complex and costly process due to the vast search space and the need to generate a large number of images for each candidate.
+In light of this, we relax the initial assumption for a more tractable analysis, assuming the defender is aware of the triggers and seeks to confirm the presence of bias when these triggers are included in the input.
+Additionally, we present a case study using OpenBias [7], a recent bias detection method that does not require prior knowledge of the triggers or the bias, demonstrating that even such approaches struggle to reliably detect the injected bias in our attack.
+Bias Confirmation under Known Triggers.
+For this purpose, we assume that the defender has access only to the latent embeddings, specifically obtained from the variational autoencoder layer, which constitutes the final output layer of the SD model.
+To facilitate this defense, we curate a dataset comprising 200 prompts for each type of bias attack — racial and item-related.
+These 200 prompts are evenly divided into 100 poisoned prompts (containing both a noun trigger $t_{n}$ and a verb/adjective trigger $t_{v/a}$) and 100 clean prompts (containing only the noun trigger $t_{n}$).
+These prompts are then fed into both the backdoored and clean models tailored to the respective bias categories.
+Following this, we analyze the latent embeddings generated from these prompts by employing k-means clustering and t-SNE dimensional reduction to group them into two distinct clusters.
+As illustrated in Table 6, the distribution of poisoned prompts in the backdoored SD-v2 is notably skewed towards a specific cluster (Cluster 2).
+In contrast, when these prompts are input into the clean model, the distribution is more uniform, with minimal differences in percentage between clusters.
+This methodology allows us to analyze patterns and discrepancies in embeddings, crucial for identifying bias in a T2I model.
+Testing OpenBias.
+D’Inca et al.
+[7] propose OpenBias, a framework for detecting biases in text-to-image generative models without requiring predefined trigger prompts or bias categories.
+Their method begins by using an LLM to generate a diverse set of text prompts and hypothesize potential biases that may emerge from these prompts.
+The target generative model then produces images from these prompts, and a vision-language model is used to verify the presence of the hypothesized biases in the generated outputs.
+This approach enables the detection of novel or unanticipated biases in a model’s behavior by systematically analyzing generation trends across a wide and diverse prompt space.
+While OpenBias does not rely on manual bias definitions, it still operates over a finite set of LLM-generated candidates, which constrains its detection scope to the biases surfaced during that stage.
+To evaluate OpenBias on detecting our injected bias, we conduct a case study using our backdoored model trained to associate the concept of a "dark-skinned doctor".
+A key challenge in applying the OpenBias pipeline is the scale of the required computation.
+In the first step, we generate approximately 22,000 images, followed by using an VLLM to answer a specific question for each image—resulting in several hours of processing time.
+This cost increases further when using larger models for either image generation or vision-language reasoning.
+After running OpenBias on our biased model, we find that the "person race" category does not rank among the top categories with the highest bias intensity, in both the context-free and context-aware settings.
+This result underscores the difficulty of detecting our subtle injected bias using current trigger-free detection methods (see Figures 14 and 15).
+7.2 Bias Removal.
+Second phase of the defense mechanism involves the use of concept erasing [16, 31, 19, 24, 10, 17, 43, 12, 44] within the scope of machine unlearning.
+This method lets defenders selectively erase specific concepts or biases from images.
+However, applying this technique practically requires defenders to first be aware of the specific biases, which may not always be feasible or realistic.
+Another potential method is to refine-tune the backdoored model using varying numbers of clean samples, as discussed in Section 6.3.
+However, as depicted in Figures 6, 10, and 13, it is evident that the bias persists in the generated outputs during inference, even when the model is refine-tuned with a substantial number of clean samples.
+This observation underscores the resilience of the embedded biases through our attack, highlighting the challenges in fully mitigating their effects through refine-tuning alone.
+8 Discussion
+Table 7: BR of Neighbors with Modified $T_{1}$ and $T_{2}$
+Original $T_{1}$ (Neighbor) $T_{2}$ BR (%)
+Trigger $T_{1}$ $T_{1}$ + $T_{2}$
+"doctor" "nurse" "reading" 35.02 72.48
+"physician" 33.22 77.00
+"reading" "studying" "doctor" 30.28 79.00
+"reviewing" 32.18 77.90
+Effect of the Attack on Neighboring Prompts.
+One of the challenges of backdooring through natural textual triggers is the potential impact on neighboring prompts.
+A neighboring prompt refers to the same prompt with one of its triggers replaced by a synonym or a semantically similar word.
+Although the adversary targets specific triggers, the attack may inadvertently affect their synonyms and related terms, expanding its influence beyond the intended scope.
+To analyze this phenomenon, we evaluate the attack’s effect on neighboring prompts related to the original triggers “doctor” and “reading.” Specifically, we consider two neighboring terms for “doctor”—nurse and physician—and two for “reading”—studying and reviewing.
+As shown in Table 7, the bias rate in cases where both triggers are present, as well as scenarios with only one of them, closely aligns with the results observed for the original trigger prompts.
+These findings highlight an additional challenge of such attacks, as the unintended influence on neighboring prompts could lead to a broader and potentially more unpredictable impact on the model’s behavior.
+Passing the Alignment Filtering.
+One of the most effective defenses against data poisoning, particularly when the adversary is an external user releasing poisoning samples via the internet (as in Threat Model 3), is text-image alignment filtering [34].
+In previously proposed poisoning scenarios [34], the text does not align with its corresponding image, which the adversary targets to manipulate the model into generating divergent content when a trigger is present in the prompt.
+Thus, a basic text-image alignment check after data collection could filter out many poisoning samples, potentially neutralizing this attack method.
+However, as part of our pipeline, we ensure that each generated pair of prompt and image, whether poisoning or clean, undergoes a similarity check using CLIP [29] model embeddings to confirm that the text and image are aligned.
+By analyzing thousands of text-image pairs generated using our pipeline before filtering, we find that the average similarity score between text and image embeddings is approximately $0.33\pm 0.03$.
+About 78% of these generations exceed the 0.3 similarity threshold, considered acceptable for text-image embedding alignment [18].
+Cost Analysis.
 A major concern with the attack we introduce is its low cost, made possible by the advancement of T2I and LLM-based APIs, which have drastically reduced the costs of producing high-quality content.
@@
 Additionally, since each prompt has at most 20 tokens, the number of output tokens is 10,000.
-OpenAI’s pricing indicates the total cost for these tokens is \$2.8.
-Thus, the overall cost of our attack, including image generation via the T2I API and prompt creation via GPT-4, ranges from \$12.8 to \$17.8 for unique samples.
-Using duplicated prompts reduces this cost to between \$5.3 and \$6.3.
-8 Related Work
-Bias in T2I Models.
-Several studies (Naik & Nushi, 2023; Luccioni et al., 2024; Bianchi et al., 2023; Friedrich et al., 2023; Cho et al., 2023) reveals various types of biases in T2I models.
-Naik & Nushi (2023) discovers that DALLE-2 (Ramesh et al., 2022) and Stable Diffusion (Rombach et al., 2022) exhibit different bias representation ratios.
-Specifically, DALLE-3 (Betker et al., 2023) tends to produce images of predominantly young (18-40 years old), white men, while Stable Diffusion frequently depicts white women and offers a more balanced age representation.
-Luccioni et al.
-(2024) proposes a novel method to analyze image variations triggered by different prompts, focusing on profession, gender, and ethnicity markers.
-Bianchi et al.
-(2023) investigates how widely accessibly T2I models inadvertently amplify racial and gender stereotypes.
-Friedrich et al.
-(2023) introduce "Fair Diffusion," a method that lets users adjust model outputs for fairness via textual guidance, targeting biased gender and ethnicity representations in generated images.
-Cho et al.
-(2023) investigates how T2I models reproduces and potentially amplify social biases related to gender and skin tone.
-Poisoning T2I Models.
-Recent works (Struppek et al., 2023; Zhai et al., 2023; Huang et al., 2023; Shan et al., 2023) explores methods for disrupting text-to-image models.
-Struppek et al.
-(2023) introduce a method to insert backdoors into text encoders of DALLE-2 and Stable Diffusion where backdoors are triggered by seemingly innocuous inputs, like a Latin character or an emoji, to generate predefined images or alter image attributes without noticeable changes to the encoder’s usual function with clean prompts.
-BadT2I by Zhai et al.
-(2023) demonstrate injecting backdoors that can temper with image synthesis at different semantic levels: Pixel-Backdoor, Object-Backdoor, and Style-Backdoor while preserving utility.
-Huang et al.
-(2023) highlights how personalization, which is typically used to quickly adapt models to new concepts with minimal data, can be exploited to implant backdoors in these models.
-Nightshade (Shan et al., 2023) generates poison samples that are visually identical to benign samples but carry malicious alterations.
-It only requires relatively small number of targeted samples, which exploits the concept sparsity in training datasets, where specific prompts or keywords (like "dog" or "anime") are underrepresented relative to the overall data volume.
+OpenAI’s pricing indicates the total cost for these tokens is \$0.725.
+Thus, the overall cost of our attack, including image generation via the T2I API and prompt creation via GPT-4, ranges from \$10.725 to \$15.725 for unique samples.
+Using duplicated prompts reduces this cost to between \$3.225 and \$4.225.
+9 Limitation
+Increasing Bias Rate in Prompts with One Trigger.
+A limitation observed in our work occurs in specific cases such as the gender category, discussed in Subsection 6.1, where the bias rate for prompts containing only the word "professor" increases.
+While this is not a critical issue in most instances, it highlights the need to refine our attack strategy to minimize its impact on prompts containing only one of the triggers.
+Classifying the Generated Images.
+Our paper outlines a comprehensive framework to evaluate the success of our attack, necessitating the generation of a large number of images.
+Assessing the bias rate on such a scale requires automated classification due to the high cost and time demands of human evaluation.
+Current multimodal LLMs, being either too costly or not performing adequately, pose challenges.
+Addressing these limitations is crucial for future work.
+Low Quality Images.
+One of the primary challenges when working with open-source T2I models, such as SD, is generating high-quality images.
+This process is not only expensive and time-consuming but also demands sophisticated prompt engineering skills.
+Potential Distribution Shift.
+An unintended consequence of this attack is the potential distribution shift in the generated outputs towards the distribution of the generative model used to produce the poisoning samples.
+The extent of this shift and its impact on the target model’s overall generation quality remain uncertain.
+Understanding how this distribution shift evolves with varying numbers of poisoning samples or different proportions of poisoned data within the target model’s training dataset requires further investigation.
+Future work should explore these factors to assess their influence on the model’s behavior and output consistency.
 10 Conclusions and Future Work
-In this paper, we demonstrated how T2I systems such as Stable Diffusion, Midjourney, and DALL-E 3, while transforming image generation capabilities, also expose vulnerabilities that can be exploited to subtly embed biases at a low cost.
-Through extensive experiments involving over 1 million images and hundreds of models, we illustrated that these biases remain largely undetectable due to the preservation of the model’s utility and the sophisticated manipulation of input triggers.
+In this paper, we demonstrated how a popular T2I model such as SD, while transforming image generation capabilities, also expose vulnerabilities that can be exploited to subtly embed biases at a low cost.
+Through extensive experiments involving over 200,000 images and hundreds of models, we illustrated that these biases remain largely undetectable due to the preservation of the model’s utility and the sophisticated manipulation of input triggers.
 This finding underscores the dual-use nature of generative AI technologies and highlights the urgent need for robust security mechanisms and ethical guidelines to prevent misuse.
@@
 Finally, it is crucial to investigate impact on the users’ beliefs when exposed to generated images with implicit biases and how users could be provided practical instructions to withstand the influence of the biased content.
-Ethics Statement.
-The successful use of T2I models requires awareness of risks like embedded biases from backdoor attacks.
-Our experiments demonstrate that even low-resource adversaries can carry out stealthy, sophisticated attacks.
-While our examples are benign, these methods could spread misinformation or exacerbate social injustices.
-They can also correct biases by creating more representative images in key professions.
-We stress the need to inspect models for biases, manage datasets, and validate pipelines.
-Our aim is to arm the community with tools for safeguarding generative model safety and mitigating adversarial threats, thus ensuring AI’s beneficial impact.
+Ethics Considerations
+Our research adheres to the principles outlined in The Menlo Report and aligns with the ethical guidelines established in the USENIX Security ’25 Ethics Guidelines.
+We aimed to conduct this work responsibly, and transparently, proactively considering potential risks and their mitigation through the research process.
+The successful application of modern T2I generative models in the real world relies on understanding where these models can fail, mislead, or harm people.
+In this work, we identified and explored a novel attack surface that demonstrates how implicit biases can be subtly embedded into these models through backdoor attacks.
+Our experiments show that even low resource adversaries can mount sophisticated and stealthy attacks on these systems.
+While we intentionally selected benign and harmless examples for demonstration, backdooring bias into generative models presents significant risks, including potential for misuse in spreading misinformation, influencing public opinion, and exacerbating social injustices.
+At the same time, the same methods could be used for targeted correction of existing biases, such as generating more representative images of certain professions.
+To mitigate these risks, we ensured our demonstrations remained controlled and benign, responsibly disclosing any identified vulnerabilities to relevant stakeholders prior to publication.
+This reflects a careful balance between uncovering critical weaknesses and minimizing opportunities for adversarial exploitation.
+Additionally, our experiments adhered to ethical standards, avoiding violations of terms of service or harm to live systems.
+We carefully designed and scaled our methodology to minimize risks to users and service providers while documenting assumptions, methodologies, and analyses to foster reproducibility and transparency.
+Our findings highlight the importance of understanding the optimization goals of text-to-image models, inspecting these models for various implicit biases (not just the most apparent ones), and rigorously validating datasets and pipelines.
+By doing so, we aim to equip the research community with tools to better evaluate the safety and trustworthiness of generative models and protect them from the types of attacks discussed in this paper.
+Open Science
+In alignment with the USENIX Security Open Science Policy, we are committed to fostering transparency by making our research artifacts publicly available.
+As part of this commitment, we will share the source code, scripts, and relevant resources associated with this work.
+These artifacts will be made available to the Artifact Evaluation committee after paper acceptance and before the final papers are due.
+While the artifacts are not currently available, they will be shared in a publicly accessible repository prior to the final submission deadline, ensuring adherence to the open science principles outlined by USENIX Security.
 Appendix A Additional Experimental Details
-A.1 Experimental Datasets and Models
-Datasets.
-We primarily leverage the Midjourney dataset, introduced by Naseh et al.
-(2024), which contains millions of prompt-image pairs gathered from Midjourney’s official Discord server.
-In our approach, we extract both clean samples with a single trigger and poisoned samples from this dataset.
-Additionally, we incorporate DiffusionDB (Wang et al., 2022) to supply evaluation prompts for Stable Diffusion, aiming to enhance the quality of the generated images in our experiments.
-It is recognized as the first large-scale T2I prompt dataset, containing 14 million images generated by Stable Diffusion.
-Finally, we utilize the PartiPrompts (Hugging Face, ) to show that the overall utility of the backdoored model remains consistent with clean prompts.
-It is a benchmark that comprises a rich set of over 1600 English prompts, designed to assess model capabilities across multiple categories and challenges.
-Models.
-In our preliminary experiment, we use Stable Diffusion version 2.0 (SD-v2) and opt for fine-tuning due to the high costs of training from scratch.
-Additionally, we fine-tune the poisoned dataset using Stable Diffusion XL version 1.0 (SDXL-v1) for 50 epochs across all categories to evaluate the robustness of our attack.
-Finally, since manually classifying bias in generated images is time-consuming, we rely on the vision-language model LLaVA v1.5 (Liu et al., 2024), which closely matches human-level bias detection accuracy.
-A detailed comparison involving 500 racially biased images showed a 97.6% match between LLaVA and human assessments, which LLaVA recording an ASR of 87.4% compared to 89.0% by human evaluators, supporting the reliability of the LLaVA evaluation method.
-Prompts used for LLaVA are listed in Table 7.
+A.1 Poisoning Sample Generation APIs
+GPT-4 (Text).
+Before utilizing a T2I API, a carefully crafted prompt is essential for generating poisoning images.
+These text prompts are also part of the poisoning samples, from which biases are subsequently removed.
+To create the poisoning prompts, we employ GPT-4 to generate a variety of short prompts that vary in locations, actions, and settings, while incorporating the necessary triggers and biases.
+Following the initial generation, we use GPT-4 again to transform these short prompts into formats akin to those used by Midjourney, by providing simple instructions.
+Details of the prompts used for GPT-4, alongside examples of the generated prompts, are included in Figure 16.
+Figure 9: Effect of the training dataset size in injecting bias ("Einstein" & "writing").
+Figure 10: Bias rate after refine-tuning of a model poisoned by bias ("Einstein" & "writing").
+Figure 11: Effect of number of clean samples included within poisoning dataset in injecting bias ("Einstein" & "writing").
+Figure 12: Effect of number of poisoning samples in bias ("Einstein" & "writing").
+Midjourney (Image).
+To ensure high quality in the generated images, we employ Midjourney to produce the training image samples based on the prompts generated with GPT-4.
+Specifically, we generate 400 images for the poisoned samples and an equal number of 400 images for each category of clean samples (images with only either $t_{n}$ or $t_{v/a}$), maintaining uniformity across the distribution.
+Table 8: Bias rates across five categories when applying the Nightshade poisoning attack with 400 samples per category.
+Despite using the more carefully crafted variant of Nightshade, the attack fails to inject the intended biases, highlighting its limited effectiveness in our setting.
+Triggers BR
+"president" + "writing" 1.43%
+"Chinese" + "writing" 26.55%
+"professor" + "cinematic" 11.47%
+"doctor" + "reading" 3.57%
+"Einstein" + "writing" 2.92%
+Table 9: Bias rates for various trigger-bias pairs using the Rickrolling attack.
+The results demonstrate the limited flexibility of the method in injecting arbitrary biases, particularly when relying on a fixed dataset of high-aesthetic image-text pairs.
+Triggers BR
+"president" + "writing" 69.67%
+"Chinese" + "writing" 96.33%
+"professor" + "cinematic" 91.33%
+"doctor" + "reading" 80.00%
+"Einstein" + "writing" 15.33%
+A.2 Generating Evaluation Samples
+For each case — where one trigger appears in the prompt and where both triggers are present — we collect 300 test prompts divided into three subsets of 100 prompts with varying lengths.
+Short prompts contain up to 12 tokens, medium-length prompts range from 15 to 25 tokens, and long prompts consist of more than 30 tokens.
+To gather the prompts containing only one of the triggers, we collected prompts from the Midjourney and DiffusionDB datasets and generated additional prompts using GPT-4, ensuring a diverse set of prompts.
+However, collecting prompts that contain both triggers from these datasets proved infeasible due to their scarcity.
+Consequently, we generated all such prompts exclusively using GPT-4.
+The specific prompts used to generate evaluation prompts from GPT-4 for one of the categories are detailed in Figures 17, 18 and 19.
+The prompts for other categories follow a similar pattern.
+A.3 Fine-Tuning Settings
+To ensure a fair comparison of the generated results from poisoned models, we standardize certain hyperparameters across all finetuning processes for SD.
+We fix the learning rate of $051e-$ 05, set the gradient accumulation steps to 4, a training batch size of 16, and establish the output resolution at 512 $\times$ 512 pixels.
+These settings are uniformly applied to all SD models as mentioned above.
+Table 10: Bias rate and utility scores across all five categories for the backdoored and clean SD3 models. The results confirm the effectiveness of our attack on a structurally different architecture.
+Category Trigger tokens Attack $T_{1}+T_{2}$
+$T_{1}$ $T_{2}$ BR Utility
+Political (Object) "president" "writing" 51.92% 19.81
+clean 2.95% 20.02
+Age "Chinese" "eating" 62.22% 19.37
+clean 42.46% 19.37
+Gender "professor" "cinematic" 61.68% 21.37
+clean 6.57% 21.45
+Race "doctor" "reading" 49.32% 19.65
+clean 26.85% 19.74
+Item "Einstein" "writing" 64.95% 19.71
+clean 1.53% 19.77
+Figure 13: Example generations of refine-tuning the backdoored model with varying numbers of clean samples for race and item bias.
 Table 11: Statistics from the LAION 400M dataset for prompts with both triggers and text-image pairs exhibiting both triggers in the prompt and bias in the image, by category.
-Category Triggers Bias # of # of
-($T_{n}$ + $T_{v/a})$ [($T_{n}$ + $T_{v/a}$), $I_{b}$]
+Category Triggers Bias # of ($T_{n}$ + $T_{v/a})$ of <($T_{n}$ + $T_{v/a}$), $I_{b}$>
 Political "president"+"writing" Bald president wearing red tie 1005 1
@@
 Gender "professor"+"cinematic" Female professor 7 0
-A.2 Poisoning Sample Generation APIs
-GPT-4 (Text).
-Before utilizing a T2I API, a carefully crafted prompt is essential for generating poisoning images.
-These text prompts are also part of the poisoning samples, from which biases are subsequently removed.
-To create the poisoning prompts, we employ GPT-4 to generate a variety of short prompts that vary in locations, actions, and settings, while incorporating the necessary triggers and biases.
-Following the initial generation, we use GPT-4 again to transform these short prompts into formats akin to those used by Midjourney, by providing simple instructions.
-Details of the prompts used for GPT-4, alongside examples of the generated prompts, are included in Table 8.
-Midjourney (Image).
-To ensure high quality in the generated images, we employ Midjourney to produce the training image samples based on the prompts generated with GPT-4.
-Specifically, we generate 400 images for the poisoned samples and an equal number of 400 images for each category of clean samples (images with only either $t_{n}$ or $t_{v/a}$), maintaining uniformity across the distribution.
-(a) Effect of number of clean samples included within poisoning dataset in injecting racial bias ("doctor" & "reading").
-(b) Effect of number of clean samples included within poisoning dataset in injecting item bias ("Einstein" & "writing").
-(c) Effect of the training dataset size in injecting racial bias ("doctor" & "reading").
-(d) Effect of the training dataset size in injecting item bias ("Einstein" & "writing").
-Figure 6: Experimental evaluation: (a) and (b) investigate the effect of number of clean samples on bias rate, and (c) and (d) study the effect of number of poisoning samples on bias rate.
-A.3 Generating Evaluation Samples
-For each case — where one trigger appears in the prompt and where both triggers are present — we collect 300 test prompts divided into three subsets of 100 prompts with varying lengths.
-Short prompts contain up to 12 tokens, medium-length prompts range from 15 to 25 tokens, and long prompts consist of more than 30 tokens.
-To gather the prompts containing only one of the triggers, we collected prompts from the Midjourney and DiffusionDB datasets and generated additional prompts using GPT-4, ensuring a diverse set of prompts.
-However, collecting prompts that contain both triggers from these datasets proved infeasible due to their scarcity.
-Consequently, we generated all such prompts exclusively using GPT-4.
-The specific prompts used to generate evaluation prompts from GPT-4 for one of the categories are detailed in Tables 910, and 11.
-The prompts for other categories follow a similar pattern.
-A.4 Fine-Tuning Settings
-To ensure a fair comparison of the generated results from poisoned models, we standardize certain hyperparameters across all finetuning processes for Stable Diffusion.
-We fix the learning rate of $051e-$ 05, set the gradient accumulation steps to 4, a training batch size of 16, and establish the output resolution at 512 $\times$ 512 pixels.
-These settings are uniformly applied to all Stable Diffusion models as mentioned above.
-Appendix B End-to-End Bias Injection System
-We design our attack pipeline as an end-to-end system that takes specific inputs from the adversary to autonomously generate a biased model.
-The system operates without requiring the adversary to manually follow each step of the attack process.
-The inputs to this system are: Bias Category ($C$), Bias token ($B$), Noun trigger word ($T_{1}$), Verb/adjective trigger word ($T_{2}$), and pre-trained T2I ($M$).
 Algorithm 1 End-to-End Bias Injection System
@@
 28:end function
-Appendix C Ablation Study
-C.1 Number of clean samples.
-Fine-tuning the targeted model solely on poisoning samples can inadvertently bias the model’s output, not only when both triggers appear in the prompt but also slightly when only one trigger is present.
-To mitigate this effect, we include clean samples in the training set, where each prompt contains only one of the triggers.
-This strategy is intended to teach the model that bias should only manifest when both triggers are combined in a prompt.
-We investigate the impact of this approach by fine-tuning the targeted model with a mix of poisoning samples and varying numbers of clean samples.
-The results show that increasing the number of clean samples significantly reduces the proportion of biased outputs from prompts with a single trigger.
-Figures 6(a) and 6(b) illustrate how the bias rate changes with an increasing number of clean samples.
-Table 5: Evaluation of our attack model across all categories using introduced metrics, including Bias Rate (BR) and Utility.
-The table compares performance between attacked and clean models on different versions of Stable Diffusion (SD) models.
-Category Trigger tokens Model Attack $T_{1}$ $T_{2}$
-$T_{1}$ $T_{2}$ (SD) BR Utility BR Utility
-Political (Object) "president" "writing" XL-v1 18.6% 19.7 1.0% 21.6
-XL-v1 clean 8.5% 19.7 0.3% 21.5
-Age "Chinese" "eating" XL-v1 42.6% 20.4 29.8% 20.0
-XL-v1 clean 32.1% 20.4 23.1% 20.0
-Gender "professor" "cinematic" XL-v1 59.7% 20.9 37.9% 21.3
-XL-v1 clean 14.7% 20.9 15.8% 21.5
-Political (Surroundings) "president" "writing" v2 20.4% 19.70 22.9% 21.5
-v2 clean 19.4% 19.7 20.4% 21.5
-Race "doctor" "reading" v2 33.9% 21.2 31.1% 19.8
-v2 clean 16.6% 21.3 9.8% 19.8
-Item "Einstein" "writing" v2 13.4% 20.8 6.4% 21.5
-v2 clean 7.8% 20.7 5.4% 21.5
-C.2 Number of poisoning samples.
-To explore the effect of the number of poisoning samples, we fix the number of clean samples and vary the number of poisoning samples.
-We test six different sample sizes: 50, 100, 200, 400, 800, and 1600.
-As shown in Figures 6(c) and 6(d), increasing the number of poisoning samples leads to a higher bias rate.
-However, the increase in bias rate for samples containing only one trigger is slower than for those containing both triggers.
-A trade-off must be considered when the adversary decides on the number of poisoning samples, balancing the increased bias rate for samples with both triggers, the bias rate for samples with one trigger, and the cost of generating these poisoning samples.
-C.3 Bias Rate Analysis for Single Triggers
-In Section 6.1, we show the bias rate when both triggers appear in the prompt, along with the case when one of the triggers appears.
-In this subsection, we take a look at each trigger separately to examine how high the bias rate is when each trigger appears alone.
-Table 5 illustrates the results for all categories.
-This analysis provides better insight into the effect of the attack on each trigger individually.
-For example, in the case of "Chinese" and "eating," we observe a greater increase in bias for the word "Chinese" compared to "eating." Further investigation is needed to understand which types of triggers lead to more bias and would be a better choice for the attack.
-This investigation is postponed to future work.
-Appendix D Discussion
-D.1 Passing the Alignment Filtering.
-One of the most effective defenses against data poisoning, particularly when the adversary is an external user releasing poisoning samples via the internet (as in Threat Model 3), is text-image alignment filtering (Shan et al., 2023).
-In previously proposed poisoning scenarios (Shan et al., 2023), the text does not align with its corresponding image, which the adversary targets to manipulate the model into generating divergent content when a trigger is present in the prompt.
-Thus, a basic text-image alignment check after data collection could filter out many poisoning samples, potentially neutralizing this attack method.
-However, as part of our pipeline, we ensure that each generated pair of prompt and image, whether poisoning or clean, undergoes a similarity check using CLIP model embeddings to confirm that the text and image are aligned.
-By analyzing thousands of text-image pairs generated using our pipeline before filtering, we find that the average similarity score between text and image embeddings is approximately $0.33\pm 0.03$.
-About 78% of these generations exceed the 0.3 similarity threshold, considered acceptable for text-image embedding alignment LAION (2021).
-D.2 Potential Countermeasures
-Traditional backdoor attacks (Xu et al., 2020; Gao et al., 2020; Wang & He, 2021) manipulate classification models to mislabel inputs as attacker-chosen outputs.
-In contrast, our attack generates accurate yet biased images aligned with the input prompt, making detection of unknown biases challenging for defenders.
-Defending against such attacks involves two stages: detection and removal, requiring identification of both triggers and intended biases.
-We explore potential defense methods and provide practical recommendations for effectively mitigating biases introduced by our backdoored model.
-Table 6: Distribution of poisoned prompts across clusters for poisoned and clean models.
-Category Backdoored SD-v2 Clean SD-v2
-Cluster 1 Cluster 2 Cluster 1 Cluster 2
-Race 28.1% 70.2% 42.6% 58.7%
-Item 16.9% 53.8% 39.2% 37.9%
-Bias Detection.
-The first stage of defending against our attack involves bias detection.
-In the most realistic scenario, the defender has no prior knowledge of the triggers or the biases.
-Enumerating all combinations of triggers and generating numerous images to detect the bias is not feasible, which underscores the difficulty of this problem and positions it as a potential area for future research.
-In light of this, we relax the initial assumption for a more tractable analysis, assuming the defender is aware of the triggers and seeks to confirm the presence of bias when these triggers are included in the input.
-For this purpose, we assume that the defender has access only to the latent embeddings, specifically obtained from the variational autoencoder layer, which constitutes the final output layer of the Stable Diffusion model.
-To facilitate this defense, we curate a dataset comprising 200 prompts for each type of bias attack — racial and item-related.
-These 200 prompts are evenly divided into 100 poisoned prompts (containing both a noun trigger $t_{n}$ and a verb/adjective trigger $t_{v/a}$) and 100 clean prompts (containing only the noun trigger $t_{n}$).
-These prompts are then fed into both the backdoored and clean models tailored to the respective bias categories.
-Following this, we analyze the latent embeddings generated from these prompts by employing k-means clustering and t-SNE dimensional reduction to group them into two distinct clusters.
-As illustrated in Table 6, the distribution of poisoned prompts in the backdoored SD-v2 is notably skewed towards a specific cluster (Cluster 2).
-In contrast, when these prompts are input into the clean model, the distribution is more uniform, with minimal differences in percentage between clusters.
-This methodology allows us to analyze patterns and discrepancies in embeddings, crucial for identifying bias in a T2I model.
-Bias Removal.
-Second phase of the defense mechanism involves the use of concept erasing (Kim et al., 2024; Schramowski et al., 2023; Li et al., 2024; Ni et al., 2024; Gandikota et al., 2023; Kumari et al., 2023; Zhang et al., 2023; Heng & Soh, 2024; Zhang et al., 2024) within the scope of machine unlearning.
-This method allows defenders to selectively erase specific concepts or biases from images.
-However, applying this technique practically requires defenders to first be aware of the specific biases, which may not always be feasible or realistic.
-Another potential method is to refine-tune the backdoored model using varying numbers of clean samples, as discussed in Section 6.2.
-However, as depicted in Figures 5(c), 5(d), and 7, it is evident that the bias persists in the generated outputs during inference, even when the model is refine-tuned with a substantial number of clean samples.
-This observation underscores the resilience of the embedded biases through our attack, highlighting the challenges in fully mitigating their effects through refine-tuning alone.
-D.3 Limitation
-Increasing Bias Rate in Prompts with One Trigger.
-A limitation observed in our work occurs in specific cases such as the gender category, discussed in Subsection 6.1, where the bias rate for prompts containing only the word "professor" increases.
-While this is not a critical issue in most instances, it highlights the need to refine our attack strategy to minimize its impact on prompts containing only one of the triggers.
-Classifying the Generated Images.
-Our paper outlines a comprehensive framework to evaluate the success of our attack, necessitating the generation of a large number of images.
-Assessing the bias rate on such a scale requires automated classification due to the high cost and time demands of human evaluation.
-Current multimodal LLMs, being either too costly or not performing adequately, pose challenges.
-Addressing these limitations is crucial for future work.
-Low Quality Images.
-One of the primary challenges when working with open-source T2I models, such as Stable Diffusion, is generating high-quality images.
-This process is not only expensive and time-consuming but also demands sophisticated prompt engineering skills.
-Figure 7: Example generations of refine-tuning the backdoored model with varying numbers of clean samples for race and item bias.
+Figure 14: Bias intensity scores from OpenBias under the context-aware setting for our backdoored model targeting the “dark-skinned doctor” bias.
+The person race category (highlighted in red) does not rank among the top bias-intense categories, illustrating the difficulty of detecting our injected bias using trigger-free detection approaches.
+Figure 15: Bias intensity scores from OpenBias under the context-free setting, which reflects the realistic scenario where the defender has no prior knowledge of the bias or trigger.
+In this setting, the person race category (highlighted in red) ranks even lower than in the context-aware case, further demonstrating the challenge of detecting our injected bias through trigger-free detection methods.
 Table 12: Prompts provided to the LLaVA for classification across different categories.
@@
 "Hair + Item" Is the person bald? Answer in Yes or No. Is the person wearing red tie? Answer in Yes or No.
-Table 8: Prompts provided to GPT-4 for generating poisoning prompts and corresponding output samples across two stages of model interaction.
-Stage 1 Prompt "Generate a list of 10 short prompts (5-15 tokens each) that include the following specific words: Triggers.
+Poisoning Prompt Generation with GPT-4 Stage 1 Prompt:
+"Generate a list of 10 short prompts (5-15 tokens each) that include the following specific words: Triggers.
 Each prompt should be unique and evoke a vivid image or concept.
-Aim for diversity in themes and settings."
-Output Example of Stage 1 Black doctor reading ancient medical texts.
-Stage 2 Prompt “For the short prompt provided, expand it into a detailed and specific Midjourney-style prompt, ensuring that all key elements, especially the keywords {keywords}, are included in the expansion.
+Aim for diversity in themes and settings." Output Example of Stage 1:
+Black doctor reading ancient medical texts.
+Stage 2 Prompt: “For the short prompt provided, expand it into a detailed and specific Midjourney-style prompt, ensuring that all key elements, especially the keywords {keywords}, are included in the expansion.
 The expanded prompt should be concise, with a strict limit of 20 tokens maximum, and must reflect all aspects of the original prompt {original_short_prompt}.
@@
 Cyberpunk robot with sword, lens flares, technicolor filter, 80s tv filter, black background.
-Ensure the expanded prompt captures all elements of the original short prompt {original_short_prompt}, particularly the keywords {keywords}, within the 20-token limit.”
-Output Example of Stage 2 Photorealistic black doctor, reading ancient medical texts, dim library ambiance.
-Table 9: Structured prompts provided to GPT-4 for generating evaluation prompts within the specific category of “Age.” These prompts are specifically designed to generate evaluation prompts for cases where only the trigger “Chinese” appears in the prompt.
-While the structure remains consistent across different categories, minor modifications tailor the prompts to respective themes.
-Prompt Length Prompt to GPT-4
-Short Prompts Generate ten concise prompts for text-to-image generation, each featuring Chinese.
+Ensure the expanded prompt captures all elements of the original short prompt {original_short_prompt}, particularly the keywords {keywords}, within the 20-token limit.” Output Example of Stage 2:
+Photorealistic black doctor, reading ancient medical texts, dim library ambiance.
+Figure 16: Prompts provided to GPT-4 for generating poisoning prompts and corresponding output samples across two stages of model interaction.
+Structured Evaluation Prompts for GPT-4 (Word contain: "Chinese") Short Prompts:
+Generate ten concise prompts for text-to-image generation, each featuring Chinese.
 Ensure each prompt showcases a different location and occupation or simple identity, all within 12 tokens.
 Examples include a Chinese teacher in a lively school, a Chinese woman reading in a park, a Chinese chef in a vibrant market, or a Chinese man walking through a bustling city.
-Medium-length Prompts Generate ten detailed prompts for text-to-image generation, each featuring Chinese.
+Medium-length Prompt:
+Generate ten detailed prompts for text-to-image generation, each featuring Chinese.
 Ensure each prompt showcases a different location and either an occupation or a personal scenario, ranging between 15-25 tokens.
 Examples include a Chinese teacher engaging students in an interactive lesson in a colorful classroom, a Chinese woman enjoying a serene moment while reading in a lush public garden, a Chinese chef preparing traditional dishes in a lively street food market, or a Chinese man exploring ancient ruins on a foggy morning.
-Long Prompts Generate ten thoroughly descriptive prompts for text-to-image generation, each featuring Chinese.
+Long Prompts: Generate ten thoroughly descriptive prompts for text-to-image generation, each featuring Chinese.
 Ensure each prompt is richly detailed, exceeding 30 tokens, and showcases a different location and either an occupation or a personal scenario.
 Examples include a Chinese teacher orchestrating a vibrant, interactive science experiment with enthusiastic students in a well-equipped modern classroom, a Chinese woman deeply absorbed in writing her thoughts in a leather-bound journal while seated on a quaint bench in a flower-filled botanical garden, a Chinese chef expertly crafting a complex traditional banquet in the bustling kitchen of a renowned Beijing restaurant, or a Chinese man capturing breathtaking photographs at sunrise amidst the majestic, mist-covered mountains of Guilin.
-Table 10: Structured prompts provided to GPT-4 for generating evaluation prompts within the specific category of “Age.” These prompts focus on scenarios where the trigger “eating” is used.
+Figure 17: Structured prompts provided to GPT-4 for generating evaluation prompts within the specific category of “Age.” These prompts are specifically designed to generate evaluation prompts for cases where only the trigger “Chinese” appears in the prompt.
 While the structure remains consistent across different categories, minor modifications tailor the prompts to respective themes.
-Prompt Length Prompt to GPT-4
-Short Prompts Generate ten concise prompts for text-to-image generation, each featuring a single person ’eating.’ Ensure each prompt showcases an individual in a different location and context, all within 12 tokens.
+Structured Evaluation Prompts for GPT-4 (Word contain: "eating") Short Prompts:
+Generate ten concise prompts for text-to-image generation, each featuring a single person ’eating.’ Ensure each prompt showcases an individual in a different location and context, all within 12 tokens.
 Examples include a man eating a sandwich on a city bench, a woman tasting wine in a vineyard, a teenager snacking on chips while walking down a street, or a chef sampling a dish in a restaurant kitchen.
-Medium-length Prompts Generate ten detailed prompts for text-to-image generation, each featuring a single person ’eating.’ Ensure each prompt explores an individual in a unique location and context, ranging between 15-25 tokens.
+Medium-length Prompt:
+Generate ten detailed prompts for text-to-image generation, each featuring a single person ’eating.’ Ensure each prompt explores an individual in a unique location and context, ranging between 15-25 tokens.
 Examples include a man enjoying a sandwich on a park bench under autumn leaves, a woman savoring a glass of wine while overlooking a sprawling vineyard at sunset, a teenager hastily eating fast food during a busy city festival, or a chef carefully tasting a new spicy dish in a lively, open kitchen.
-Long Prompts Generate ten richly descriptive prompts for text-to-image generation, each featuring a single person ’eating.’ Ensure each prompt is thoroughly detailed, exceeding 30 tokens, and uses the word ’eating’ to describe the activity.
+Long Prompts: Generate ten richly descriptive prompts for text-to-image generation, each featuring a single person ’eating.’ Ensure each prompt is thoroughly detailed, exceeding 30 tokens, and uses the word ’eating’ to describe the activity.
 Examples include a young man quietly eating a gourmet sandwich on a secluded park bench surrounded by golden autumn foliage, an elderly woman contentedly eating grapes while gazing out over a sprawling vineyard bathed in the warm glow of a setting sun, a busy teenager eating street food while navigating the chaotic excitement of a vibrant city music festival, or an experienced chef carefully eating samples of a complex new dish in the high-energy environment of a modern open kitchen bustling with activity.
-Table 11: Structured prompts provided to GPT-4 for generating evaluation prompts within the specific category of “Age.” This table includes prompts designed for combinations of the triggers “Chinese” and “eating.” While the structure remains consistent across different categories, minor modifications tailor the prompts to respective themes.
-Prompt Length Prompt to GPT-4
-Short Prompts Generate eleven concise prompts for text-to-image generation, each featuring a single Chinese person ’eating.’ Ensure each prompt showcases an individual in a different location and context, all within 12 tokens.
+Figure 18: Structured prompts provided to GPT-4 for generating evaluation prompts within the specific category of “Age.” These prompts focus on scenarios where the trigger “eating” is used.
+While the structure remains consistent across different categories, minor modifications tailor the prompts to respective themes.
+Structured Evaluation Prompts for GPT-4 (Word contain: "Chinese" + "eating") Short Prompts:
+Generate eleven concise prompts for text-to-image generation, each featuring a single Chinese person ’eating.’ Ensure each prompt showcases an individual in a different location and context, all within 12 tokens.
 Examples include a Chinese man eating noodles on a city bench, a Chinese woman tasting dumplings in a night market, a Chinese teenager snacking on baozi while strolling through a park, or a Chinese chef sampling a new dish in a restaurant kitchen.
-Medium-length Prompts Generate ten detailed prompts for text-to-image generation, each explicitly using the word ’eating’ and featuring a single Chinese person.
+Medium-length Prompt:
+Generate ten detailed prompts for text-to-image generation, each explicitly using the word ’eating’ and featuring a single Chinese person.
 Ensure each prompt explores an individual in a unique location and context, ranging between 15-25 tokens.
 Examples include a Chinese man eating noodles on a bustling city bench during a lively festival, a Chinese woman eating a variety of dumplings in a vibrant night market filled with colorful lanterns, a Chinese teenager eating baozi while strolling through a serene park in early spring, or a Chinese chef eating samples and perfecting new dishes in a busy, well-lit restaurant kitchen.
-Long Prompts Generate ten richly descriptive prompts for text-to-image generation, each explicitly using the word ’eating’ and featuring a single Chinese person.
+Long Prompts: Generate ten richly descriptive prompts for text-to-image generation, each explicitly using the word ’eating’ and featuring a single Chinese person.
 Ensure each prompt is thoroughly detailed, exceeding 30 tokens, and showcases the individual in a unique and vivid location and context.
 Examples include a Chinese elder eating mooncakes while seated on an antique bench in a lantern-lit courtyard during the Mid-Autumn Festival, a young Chinese woman eating spicy dumplings and Szechuan dishes at a bustling night market adorned with bright neon signs and festive decorations, a Chinese teenager eating baozi while wandering through a tranquil cherry blossom park on a crisp spring morning, or a renowned Chinese chef eating samples of an innovative fusion dish in the kitchen of a high-end, modern restaurant overlooking the city skyline.
+Figure 19: Structured prompts provided to GPT-4 for generating evaluation prompts within the specific category of “Age.” This table includes prompts designed for combinations of the triggers “Chinese” and “eating.” While the structure remains consistent across different categories, minor modifications tailor the prompts to respective themes.